### 30 Mar 2026 v1

# Sublogarithmic Distributed Vertex Coloring with Optimal Number of Colors

Abstract
For any $\Delta$, let $k_{\Delta}$ be the maximum integer $k$ such that $(k+1)(k+2)\leqslant\Delta$. We give a distributed $\mathsf{LOCAL}$ algorithm that, given an integer $k<k_{\Delta}$, computes a valid $\Delta-k$-coloring if one exists. The algorithm runs in $\widetilde{O}(\log^{4}\log n)$ rounds, which is within a polynomial factor of the $\Omega(\log\log n)$ lower bound, which already applies to the case $k=0$. It is also best possible in the sense that if $k\geqslant k_{\Delta}$, the problem requires $\Omega(n/\Delta)$ distributed rounds [Molloy, Reed, ’14, Bamas, Esperet ’19].
For $\Delta$ at most polylogarithmic, the algorithm is an exponential improvement over the current state of the art of $O(\log^{49/12}n)$ rounds. When $\Delta\geqslant(\log n)^{50}$, our algorithm achieves an even faster runtime of $O(\log^{*}n)$ rounds.
Contents
1 Introduction
Graph coloring is a problem of fundamental importance to combinatorics and computer science. Given a graph $G=(V,E)$ and an integer $c\geqslant 1$, a $c$-coloring of $G$ is a mapping from the vertices of $G$ to $\{1,2,\ldots,c\}$ such that adjacent vertices of $G$ receive different colors. Determining the smallest $c$ for which a given graph admits a $c$-coloring – a value known as the chromatic number denoted by $\chi(G)$ – has been an enduring challenge in graph theory since its very beginning. Computationally, it has long been known that computing exactly or even approximately the chromatic number of a graph is NP-hard [40, 52]. In general, the chromatic number of a graph depends on the global structure of the graph rather than on purely local properties (see [2, Chapter 3]).
Distributed Graph Coloring.
In the $\mathsf{LOCAL}$ model, introduced by Linial [41], the input graph $G$ is seen as a network in which $n=|V|$ nodes equipped with unique identifiers communicate with their neighbors in synchronous rounds and perform unlimited local computation, aiming to minimize the total number of rounds until each node has computed its output, e.g., its color in a graph coloring problem. In this setting, one classically aims to compute a $\Delta+1$-coloring where $\Delta$ is the maximum degree of the graph [6]. Sequentially, the problem is trivial as one can greedily assign colors to nodes without ever getting stuck, or in other words, any valid partial coloring can always be completed to a coloring of the whole graph. Essentially all efficient distributed graph coloring algorithms exploit this greedy behavior of the problem, e.g., [5, 16, 42, 32, 35, 26]. The most important exceptions are several papers that aim for an efficient distributed implementation of Brooks’ theorem [12] for coloring with exactly $\Delta$ colors, e.g., [48, 29, 24, 37, 9, 8].
Coloring with Fewer Colors.
Coloring graphs whose chromatic number is smaller than $\Delta$ using the optimal number of colors is challenging, even in the centralized setting. In a seminal paper, Molloy and Reed [45] characterized the precise threshold at which the problem becomes NP-complete. Let $k_{\Delta}$ be the largest integer such that $(k_{\Delta}+1)(k_{\Delta}+2)\leqslant\Delta$. [^1] It can be verified that $k_{\Delta}=\lfloor\sqrt{\Delta+1/4}-3/2\rfloor$ and thus $\sqrt{\Delta}-3<k_{\Delta}<\sqrt{\Delta}-1$ holds. They showed that for $c<\Delta-k_{\Delta}$ the decision problem whether a graph admits a $c$-coloring is NP-complete, while for $c\geqslant\Delta-k_{\Delta}$ the decision is based on whether the induced neighborhood $G[N[v]]$ of some node $v$ is $c$-colorable. Furthermore, for $c\geqslant\Delta-k_{\Delta}$ the coloring can be computed by a (centralized) polynomial-time algorithm when $\Delta$ is a constant. This result is highly non-trivial, spanning more than 60 dense pages [45].
In the distributed setting, Bamas and Esperet showed that for $c=\Delta-k_{\Delta}$ computing a $c$-coloring of a graph with chromatic number $c$ requires $\Omega(n/\Delta)$ rounds [4]. Based on the work of Molloy and Reed, they also gave a randomized algorithm for the $\mathsf{LOCAL}$ model to $c$-color graphs when $c>\Delta-k_{\Delta}$ in a polylogarithmic number of rounds. In contrast, the state-of-the-art algorithms for $(\Delta+1)$-coloring and, recently also for $\Delta$-coloring, only require $\operatorname{\text{{\rm poly}}}(\log\log n)$ randomized distributed time [16, 50, 29, 24, 8]. This discrepancy motivates the following question.
Can one color in a sublogarithmic number of rounds with fewer than $\Delta$ colors?
Our Results.
In this paper, we answer the question in the affirmative. We show that one can compute a $c$-coloring in sublogarithmic time when one exists and $c\geqslant\Delta-k_{\Delta}+1$. Moreover, when $\Delta$ is asymptotically larger than some polylogarithm in $n$, our algorithm runs in $O(\log^{*}n)$ rounds, matching the state of the art for the significantly easier $(\Delta+1)$-coloring problem. More precisely, we prove the following theorem:
Theorem 1.
For sufficiently large $\Delta$, and any $c\geqslant\Delta-k_{\Delta}+1$, there is a distributed randomized algorithm that takes a graph $G$ with maximum degree $\Delta$ as input, and does the following: either some vertex outputs a certificate that $G$ is not $c$-colorable, or the algorithm finds a $c$-coloring of $G$. The algorithm runs in $O(\log^{*}n)$ rounds when $\Delta\geqslant(\log n)^{50}$, and in general in $\widetilde{O}(\log^{4}\log n)$ rounds [^2] In this paper, $\widetilde{O}(f(n))$ hides multiplicative factors of size $\operatorname{\text{{\rm poly}}}\log f(n)$., with high probability.
Theorem 1 comes polynomially close to the lower bound of $\Omega(\log_{\Delta}\log n)$ rounds established for the $\Delta$-coloring problem [10]. Moreover, this result uses the fewest number of colors possible due to the aforementioned lower bound from [4]. While our approach builds on the framework of Molloy and Reed, it is conceptually and technically simpler in several key aspects; see Sections 1.2 and 2 for details.
By plugging the randomized algorithm from Theorem 1 into the powerful distributed derandomization framework of [31, 27, 50, 26] we obtain more than a quadratic improvement for deterministic algorithms. Previously the best algorithm, using state-of-the-art subroutines [26, 50], had complexity $O(\log^{49/12}n)$ [4].
Theorem 2.
For sufficiently large $\Delta$, and any $c\geqslant\Delta-k_{\Delta}+1$, there is a distributed deterministic algorithm that takes a graph $G$ with maximum degree $\Delta$ as input, and does the following: either some vertex outputs a certificate that $G$ is not $c$-colorable, or the algorithm finds a $c$-coloring of $G$. The algorithm runs in $\widetilde{O}(\log^{2}n)$ rounds.
The runtime of Theorem 2 is only roughly quadratically slower than the lower bound of $\Omega(\log_{\Delta}n)$ rounds established for the $\Delta$-coloring problem [15] and almost matches the natural barrier of $\Omega(\log^{2}n)$ rounds for the deterministic complexity of the problem. Surpassing it is believed to require fundamentally new techniques. Molloy and Reed’s, Bamas and Esperet’s, and also our algorithm for computing a coloring with fewer than $\Delta$ colors rely in multiple places on solving instances of the constructive Lovász Local Lemma (LLL). The only known method to do this deterministically in a distributed setting is via the aforementioned derandomization framework that inherently comes with an $\Omega(\log^{2}n)$ cost; see [26] for more details. Breaking this barrier for general LLLs is considered one of the biggest open problems in the field. Chang and Pettie conjectured in [13] that this can be done on bounded-degree graphs, but this question remains unanswered.
1.1 Further Background on Distributed Graph Coloring
Historic focus on Greedy problems.
Coloring was the central topic of the paper establishing the $\mathsf{LOCAL}$ model [41]. However, since this model requires that nodes decide on their colors only based on their local view, even solving greedy problems such as maximal independent set, maximal matching or $(\Delta+1)$-vertex-coloring is highly non-trivial. Linial showed that even coloring an $n$-cycle with a constant number of colors requires at least $\Omega(\log^{*}n)$ rounds for deterministic algorithms and Naor later extended this lower bound to the randomized setting [47]. It took roughly 30 years to understand that on general graphs those greedy problems could be solved deterministically in polylogarithmic distributed time thanks to a breakthrough by Rohzon and Ghaffari [50]. Over the last 5 years, follow-up work [32, 3, 22, 26] on deterministic algorithms coupled with earlier randomized techniques [5, 34, 16, 35] led to exciting progress on our understanding of the complexity of greedy problems in the $\mathsf{LOCAL}$ model.
Recent interest in the locality of non-greedy problems. [^3] Intuitively, a non-greedy problem is one where not every partial solution can be extended to a full solution. One distinguishing feature is that greedy problems can be solved in $O(\log^{*}n)$ rounds on constant-degree graphs, while non-greedy problems require at least $\Omega(\log\log n)$ rounds. See [39] for a more formal discussion on the topic.
Despite rising interest in the locality of non-greedy problems, e.g., for the seminal works on (hypergraph) sinkless orientation problems [10, 33, 11], for various degree splitting problems [28, 36, 18], and for the more amenable non-greedy edge coloring problems, e.g., [30, 51, 38, 36, 14, 18, 39], little is known for non-greedy vertex coloring. The exception is the aforementioned coloring with exactly $\Delta$ colors. A well-known theorem of Brooks [12] shows that any connected graph admits a $\Delta$-coloring unless it is a $(\Delta+1)$-clique or an odd cycle. There have been several works designing faster and faster algorithms for the problem [48, 29, 24, 37, 8] culminating in the very recent result of [9] obtaining an optimal round complexity when $\Delta=O(1)$. The only known works attempting to color with fewer than $\Delta$ colors are for very sparse graphs, i.e., triangle-free graphs [17], but their randomized complexities remain at least logarithmic. In a different direction, Barenboim designed a randomized distributed algorithm that computes an $O(n^{1/2+\varepsilon}\chi(G))$-coloring of any graph $G$ [7].
The importance of the Lovász Local Lemma.
The Lovász Local Lemma (LLL) provides a powerful probabilistic framework for showing that, even when many local bad events may occur, there exists a global configuration that avoids all of them simultaneously. Informally, it can be viewed as a localized analogue of the union bound: if each bad event occurs with sufficiently small probability and depends on only a limited number of other events, then one can still guarantee the existence of an outcome in which no bad event occurs. In the constructive version of the LLL, the goal is not only to prove existence but also to efficiently compute such an assignment.
Molloy and Reed’s algorithm for coloring with $c\geqslant\Delta-k_{\Delta}$ heavily relies on multiple instances of the constructive LLL (see Section 11 in [45]). In essence, the LLL is used to ensure that partial colorings retain properties similar to those of a random coloring, thereby guaranteeing that they can always be extended without creating deadlocks. In their distributed adaptation, Bamas and Esperet significantly reduced the number of required LLL instances and solved each of them using a logarithmic-time distributed LLL solver based on the Moser–Tardos framework [4, 46, 17]. To date, no non–LLL-based approach is known for computing such colorings.
In the distributed setting, random variables and bad events are naturally associated with nodes of the communication network, and the objective mirrors that of the constructive LLL: to compute a global assignment that avoids all bad events. Consequently, understanding the distributed complexity of the LLL has become a central challenge in the field. In a seminal result, Chang and Pettie [13] showed that the LLL is complete for sublogarithmic-time computation of local graph problems on constant-degree graphs. As a consequence, any such problem with sublogarithmic complexity can be solved in $\operatorname{\text{{\rm poly}}}(\log\log n)$ rounds using the distributed LLL algorithm of Fischer and Ghaffari [23].
For general graphs, however, it remains a major open question which problems admit sublogarithmic-time distributed algorithms. The complexity of the general distributed LLL on general graphs is also widely open, despite recent progress on special cases [27, 18] and results showing sublogarithmic average-time complexity per node [19]. Moreover, the constructive LLL admits an $\Omega(\log\log n)$ randomized lower bound [10]. Our work not only yields sublogarithmic-time algorithms for coloring with fewer than $\Delta$ colors, but also provides sublogarithmic-time solutions for several LLL instances that, prior to our work, were only known to admit logarithmic-time distributed solutions.
1.2 The Challenges and Our Technical Contributions
Our approach follows the same high-level structure as that of Molloy and Reed [45]. As in their work, we first transform the input graph to expose its structural regularities and then decompose it into five components: a sparse region, two dense regions consisting of cliques, and two intermediate regions linking them. These components are colored in a carefully chosen order while maintaining probabilistic invariants that guarantee that the dense cliques can be completed in the final stages. We refer to Section 2 for a more detailed overview of this framework.
Challenges in achieving sublogarithmic distributed algorithms.
While the Molloy–Reed framework provides a conceptually clean approach, several aspects pose major challenges in the distributed setting:
  1. 1.
    The method of [45] for coloring the sparse and intermediate regions relies on an iterative semi-random process. This procedure is technically involved—requiring a lengthy analysis—and inherently slow, as it performs $\Delta^{\Theta(1)}$ iterations, each requiring the solution of an LLL instance. Although Bamas and Esperet reduced this to $O(\log^{13/12}\Delta)$ iterations, this complexity remains far too large for achieving sublogarithmic distributed time.
  2. 2.
    The probabilistic arguments throughout [45] are based on the Lovász Local Lemma. Known distributed LLL solvers require $\Omega(\log n)$ rounds in general [4, 46, 17]. While some LLL instances can be solved faster when the criteria are sufficiently relaxed [36, 18], others – most notably those arising from the dense regions – satisfy only polynomial LLL criteria and thus resist existing fast techniques.
As a consequence, a direct distributed implementation of the Molloy–Reed framework inevitably leads to high round complexity. Notably, this difficulty is most pronounced for high-degree graphs, despite the fact that such graphs admit ultrafast distributed algorithms for simpler problems such as $\Delta$-coloring [24].
Our Contributions.
We present a sublogarithmic-time distributed algorithm for coloring with $\Delta-k_{\Delta}+1$ colors by refining and simplifying the Molloy–Reed framework.
  1. 1.
    A central slack invariant. Our main conceptual contribution is the introduction of a central invariant based on slack, ensuring that each vertex consistently has more available colors than competing constraints. The key feature of this invariant is its monotonicity: as the algorithm progresses, the loss of available colors is matched by a corresponding reduction in uncolored neighbors, so that slack does not deteriorate.
    This viewpoint allows us to avoid tracking the detailed evolution of color lists and degrees over time. Rather than carefully controlling distributions at each step, we rely on the invariant to ensure that partial colorings remain extendable, leading to a simpler and more robust approach to coloring the sparse and intermediate regions.
  2. 2.
    Faster handling of structured LLL instances. We develop sublogarithmic-time solutions for the Lovász Local Lemma instances arising in our setting. Our approach builds on the shattering framework, augmented with additional mechanisms—such as guard events—that ensure the residual instances remain well-behaved and avoid cascading dependencies.
    While the individual ingredients have appeared in prior work, their combination allows us to handle LLL instances that were not previously known to admit such fast distributed solutions.
  3. 3.
    Parallel coloring of dense cliques. For dense regions, we introduce a subsampling technique that enables cliques to be colored in parallel by reducing the problem to independent matching tasks within each clique. This avoids the need for global coordination and simplifies the treatment of dense structures compared to prior approaches.
  4. 4.
    Ultrafast algorithms in the high-degree regime. In graphs of sufficiently large degree, we can replace the core iterative coloring step with a reduction to a form of list coloring that admits very fast distributed algorithms. As all our probabilistic claims hold with subexponential error bounds, we also bypass the application of LLL altogether. This leads to a $O(\log^{*}n)$-time solution when $\Delta\geqslant\log^{50}n$.
Conceptually, our results expand the small but growing class of explicit distributed LLL formulations that can be solved in sublogarithmic time on general graphs. Despite this progress, designing a general distributed LLL algorithm that efficiently handles all relevant instances across all degree regimes remains a major open problem in the field.
1.3 Organization of the Paper
In Section 2, we present the main technical ideas underlying our approach, while Section 3 contains complementary background material. The algorithm is described formally in Section 4, where we also prove Theorems 1 and 2. The subroutines and their analyses for coloring $\Pi$-ous subgraphs, sparse vertices, and dense vertices are presented in Sections 5, 6 and 7, respectively. The high-degree case is treated separately in Section 8.
The substantial technical detail throughout the paper is driven by the goal of achieving sublogarithmic distributed runtime. If one were only concerned with the existence of such colorings, a polynomial-time centralized algorithm, or a polylogarithmic-time distributed solution, many sections could be significantly simplified.
2 Technical Overview
We give an overview of the main technical ideas behind Theorem 1, focusing on how to construct a $(\Delta-k_{\Delta}+1)$-coloring in sublogarithmic distributed time. In Section 2.3, we attempt to give a bird’s eye view of the techniques allowing us to solve each instance of the Lovász Local Lemma in sublogarithmic time.
2.1 Color Coverage (CC) & $\Pi$-ous subgraphs
Due to the impossibility of completing all partial solutions in a non-greedy coloring problem, the hardest part of the problem is coloring the last vertices. In a non-greedy coloring problem, one has to construct a coloring while ensuring that the remaining nodes can be colored later. Following Molloy and Reed, we color the dense vertices last and begin by reviewing the conditions necessary for coloring the cliques at the end.
Color Swapping in Cliques.
Consider a $(\Delta-k_{\Delta}+1)$-clique in which every vertex is adjacent to $k_{\Delta}$ nodes outside the clique and all but one vertex have been properly colored. To color this remaining vertex $v$, which we call unhappy, Molloy and Reed swap its color with that of some other vertex in the clique. In order to do so without creating new conflicts, the vertex $u$ with which $v$ swaps its color must have no external neighbors with $v$’s color (and vice versa). See Figure 1(a). However, it may happen that the coloring outside the clique does not permit any such swap (as in Figure 1(b)).
$A_{i}$Swappableunhappy
$A_{i}$unhappy
Figure 1: A clique $A_{i}$ in a graph with $\Delta=9$ and $k=3$. On Figure 1(a), the clique $A_{i}$ contains one unhappy vertex, in blue. It cannot swap its color with the purple or green vertex because they have a blue external neighbor. It cannot swap its color with the red vertex because it has a red external neighbor. The remaining vertices (in cyan and yellow) make the $\mathrm{Swappable}$ set for this unhappy vertex. On Figure 1(b), the $\mathrm{Swappable}$ set is empty because too many vertices on the outside adopted the blue color.
To circumvent this issue, Molloy and Reed constrain all earlier steps of the coloring algorithm so that the coloring outside each clique behaves sufficiently randomly. Concretely, this means that for every clique and every color, only a small number of vertices are forbidden from using that color. We formalize this requirement as the color coverage (henceforth CC) property. Ensuring that CC holds until coloring the dense parts of the graph is essential to completing the coloring. See Section 4.2 for the precise definition of CC.
$\Pi$-ous subgraphs.
The CC condition can be maintained if all earlier coloring steps are sufficiently random. If each dense node has at most $U$ neighbors outside its clique, and each of them picks one random color out of $U\cdot\Delta^{0.22}$ colors, then few of these outside neighbors pick the same color, thereby maintaining the CC condition. Thus, the key to ensuring the CC condition in each step of the algorithm is to maintain large lists of available colors $L(v)$ (colors that do not currently appear in the neighborhood of $v$) compared to the number of outside neighbors of the cliques. We capture this condition through $\Pi$, which we state in a simplified form as follows:
Property $\Pi$ (formalized in Section 4.3).
An induced subgraph $H$ satisfies Property $\Pi$ if (a) each uncolored dense vertex has at most $U$ outside neighbors, and (b) each vertex $u\in H$ has a list $L(u)$ of at least $d_{H}(u)+U\cdot\Delta^{0.22}$ available colors.
Intuitively, Property $\Pi$ guarantees that every vertex has enough slack compared to the number of neighbors outside its dense structure.
Subgraphs satisfying $\Pi$ are called $\Pi$-ous (pronounced pious). We can color a $\Pi$-ous subgraph $H$ while maintaining CC by a combination of iterated random color trials (Rct) and multi color trials (Mct). Condition (b) of $\Pi$ ensures that the list of available colors for each vertex of $H$ has $U\cdot\Delta^{0.22}$ slack that remains and ensures a sufficiently large list regardless of how the coloring of $H$ proceeds. Slack does not decrease because coloring a neighbor removes both one competing vertex and one available color, leaving their difference—the slack—unchanged or increased.
Property $\Pi$ is strictly stronger than assumptions of [45]. Their assumption (8.2) on the list size involves a max of two terms, not the sum, which means that slack is only guaranteed when degrees are high.
2.2 Overview of the Coloring Algorithm
Our algorithm centers on identifying $\Pi$-ous subgraphs that can be colored while maintaining CC, thereby ensuring that the dense vertices can be colored last. We build on the structural decomposition of Molloy and Reed, which nearly [^4] The subgraphs induced by $B_{H}$ and $B_{L}$ are $\Pi$-ous by construction, however the subgraph induced by $S$ is not $\Pi$-ous right away; see Section 6 for more details. partitions the sparser regions of the graph into $\Pi$-ous subgraphs [45]. More precisely, they reduced the problem of finding a $c$-coloring of $G$ to the problem of finding a $c$-coloring of a graph $F$ with a very specific structure. Bamas and Esperet also showed that $F$ can be constructed from $G$ in $O(1)$ rounds in the $\mathsf{LOCAL}$ model, and that one can efficiently recover a $c$-coloring of $G$ from a $c$-coloring of $F$ in the distributed setting [4].
Structural Decomposition (formalized in Lemma 4.1).
The vertices of the graph $F$ are partitioned into five sets as follows: Sparse vertices ($S$): vertices whose neighborhoods are far from being cliques, either having noticeably lower degree or many missing edges among their neighbors, High-slack intermediates ($B_{H}$): vertices outside the cliques that have many neighbors in cliques, Cliques with large external degree ($A_{H}$): true cliques of size between $c-\Theta(\sqrt{\Delta})$ and $c$ and whose vertices have at most $O(\sqrt{\Delta})$ external neighbors, Low-slack intermediates ($B_{L}$): vertices that still have many neighbors in cliques, but not as many as the vertices in $B_{H}$, Cliques with small external degree ($A_{L}$): the final cliques to be colored, each with even smaller external degree bounded by $O(\Delta^{1/4})$.
$A_{i}$$All_{i}$BBSS
Figure 2: A schematic illustration of the structural decomposition by Molloy and Reed. The vertices of $S$ are loosely connected, while the set $A_{i}$ is a clique. The intermediate set $B$ is too highly connected to AA to be considered sparse, but not densely enough to be itself part of the clique. For the coloring, we further divide the set $B$ into two subsets, $B_{H}$ and $B_{L}$ (see Lemma 4.1); however, the figure shows $B$ as a single set for simplicity. Note that some vertices of $B$, here in the set $All_{i}$, can be connected to all the vertices of a clique (see Lemma 4.1 for details on the set $All_{i}$).
While we follow the same overall structure as Molloy and Reed’s algorithm, the internal components differ significantly. Most notably, the fact that we aim to color $\Pi$-ous subgraphs means that we can use a simpler analysis. First, we observe that $B_{H}$ and $B_{L}$ are $\Pi$-ous by definition; we then provide a simple method to partition $S$ into two $\Pi$-ous subgraphs; and finally design a highly parallelizable approach for coloring the cliques. Coloring $S,B_{H}$ and $B_{L}$ is one of the most technically involved parts in [45]. We shortly detail each of these steps.
Coloring nodes in $B_{H}$ and $B_{L}$.
By construction, nodes in $B_{H}$ and $B_{L}$ have many uncolored neighbors in cliques, which provides them with temporary slack; as a result, they satisfy $\Pi$ purely due to structural properties. It is crucial that the vertices in $B_{L}$ are colored after the cliques in $A_{H}$. Although lists of vertices in $B_{L}$ are guaranteed to have size $\Omega(\sqrt{\Delta})$, this is insufficient for $B_{L}$ being $\Pi$-ous if $A_{H}$ is still uncolored.
Coloring the Sparse Vertices in $S$.
It is well known that a single round of random color trials (Rct) gives each vertex $v\in S$ slack $\Omega(\sqrt{\Delta})$. This process is called slack generation. Unfortunately, this is still insufficient for $\Pi$ requiring slack $U\cdot\Delta^{0.22}$ where $U=\Theta(\sqrt{\Delta})$. We therefore split the remaining uncolored vertices $S^{\prime}$ into two groups $S_{1}$ and $S_{2}$. Coloring $S_{1}$ first gives their vertices extra temporary slack (from neighbors in $S_{2}$), and when we color $S_{2}$ we can rely on the fact that dense vertices have few neighbors in $S_{2}$ making it easier to satisfy $\Pi$.
Coloring of Cliques.
To achieve a sublogarithmic distributed runtime, we color cliques in $A_{H}$ and $A_{L}$ respectively in parallel. For each clique $A_{i}$, we assign a random permutation of the colors to its vertices avoiding any conflicts inside the clique. With this process, only $O(\sqrt{\Delta})$ vertices per clique conflict with external neighbors. Each such unhappy vertex $v$ seeks a partner $u$ in $A_{i}$ so that both become happy if they swap their colors. CC crucially guarantees a large supply of valid swap candidates.
To execute swaps in all cliques concurrently, we first independently downsample the candidate sets and remove those causing external conflicts, i.e., swaps that cannot be executed safely in parallel with sampled candidate swaps in adjacent cliques. We show that, after this pruning step, sufficiently many candidates remain to also avoid internal conflicts.
We set up a bipartite graph between unhappy vertices and their remaining candidates and show that it satisfies Hall’s condition, ensuring a perfect matching between unhappy nodes and swap candidates. We can then perform the corresponding swaps simultaneously without interference within the clique and due to the prior pruning also without interference between different cliques.
2.3 Smaller Degrees and the Lovász Local Lemma
All steps of our coloring algorithm are randomized and have a local failure probability of the form $\exp(-\Delta^{\varepsilon})$ for some universal constant $\varepsilon\in(0,1]$. Hence, when $\Delta\gg\log^{1/\varepsilon}n$, our algorithm succeeds with high probability. However, when $\Delta\leqslant\operatorname{\text{{\rm poly}}}(\log n)$, local failures can no longer be avoided.
The main tool at our disposal to handle local failures is the Lovász Local Lemma. Since we aim for a sublogarithmic runtime, we cannot rely on the general framework by Moser and Tardos [46, 17]. The only general approach currently known to achieve sublogarithmic round complexity is the shattering framework [5]. Unfortunately, there exists no general-purpose distributed LLL solver yielding sublogarithmic time on graphs of all degrees, in particular in the challenging regime where degrees are sublogarithmic but exceed $\operatorname{\text{{\rm poly}}}(\log\log n)$.  [^5] We cannot provably rule out that some of our LLLs could be solved via the sublogarithmic-time algorithm of [27], but we do not know how to prove that they fall into the handled class of LLLs either. For very low degree graphs, sublogarithmic-time algorithms are known [23, 18]. We therefore design dedicated constructions tailored to the different degree regimes.
The Shattering Framework.
In this framework, a fast, randomized pre-shattering phase resolves most of the graph – for instance, fixing the colors of the majority of vertices – leaving only small residual components of size $\operatorname{\text{{\rm poly}}}(\log n)$. A subsequent post-shattering phase then deterministically completes the solution by solving a smaller residual LLL problem, typically within $\operatorname{\text{{\rm poly}}}(\log\log n)$ rounds.
As one randomly colors vertices in the pre-shattering phase, one may be forced to uncolor certain vertices. For instance, if a color appears too frequently around a clique and CC is violated. This may interfere with the progress measures of our algorithm, which are typically formulated as additional bad events. For instance, uncoloring vertices may violate an event ensuring a constant-factor decrease in the uncolored degree. Such events are usually resolved during the post-shattering phase.
The key challenge, especially for non-greedy problems such as ours, is to design the pre-shattering phase in such a way that the post-shattering instance (1) is solvable and (2) has polylogarithmic size components, meaning that retractions (e.g., uncolorings) do not percolate to the large parts of the graph. We introduce several methods to deal with these challenges.
  • Fresh budgets are used in the pre-shattering and the post-shattering whenever we do not want to see too many of something. By giving half of the total budget to each of the shattering steps, we remove the dependencies between the two phases.
  • Marginal events are added in the post-shattering phase to account for previously avoided bad events that could be reactivated by newly assigned variables.
  • Guard events impose additional constraints to preserve sufficient randomness for the post-shattering phase. They prevent the pre-shattering process from fixing too many variables or overcommitting to specific random outcomes, which would reduce the available randomness needed later.
  • Palette splitting further separates randomness between pre- and post-shattering: one part of the color palette is used in pre-shattering, the other in post-shattering. Despite this restriction, the probabilistic guarantees of each step continue to hold.
Example 1: Subsampling Swap Candidates.
To illustrate those two ideas, consider the LLL that governs subsampling of swap candidates in cliques. Each unhappy vertex $v$ independently samples potential partners with a fixed probability; with overwhelming probability, every vertex obtains sufficiently many candidates. The subtlety lies in showing that, after pruning candidates involved in conflicting swaps across adjacent cliques, enough valid partners remain. This is achievable, and the shattering analysis ensures that the components where the sampling fails are small.
Since we only pick candidates in pre- and post-shattering using separate budgets, this increases the number of neighbors with a given color by a factor of at most two compared to having one single budget for both steps. Conveniently, the same analysis bounds the probability of breaking CC for the pre- and post-shattering.
Removing candidates does not make other candidates bad, however, resampling candidates to repair local failures can affect neighboring cliques, and potentially undo the progress made during pre-shattering. To resolve this, we add an event in post-shattering for each clique adjacent to some other clique that resamples its candidates in post-shattering.
Example 2: Slack Generation.
To illustrate the latter two techniques, let us look at the slack generation LLL. Sparse nodes have many non-edges in their neighborhood, i.e., pairs of non-adjacent nodes. One can show that one round of random color trials provides a node with slack proportional to the number of non-edges by same-coloring the endpoints of many of these non-edges. One can show that receiving sufficient slack forms an LLL.
In a pre-shattering phase based on randomly coloring vertices, a node may also obtain too little slack. Now, if all of its neighbors are already colored, it is impossible for it to obtain slack in the post-shattering phase. Uncoloring its neighbors would remove the slack of other nodes and in fact this may percolate through the graph. Instead, we use a guarding event that ensures that in the pre-shattering phase few enough neighbors of each node participate in the coloring procedure ensuring that enough neighbors (and also non-edges between them) remain uncolored for a tentative post-shattering phase. Of course, this event can also fail but one can verify that the guarding event prevents percolation of uncolorings.
If enough neighbors (and also non-edges between them) remain uncolored in post-shattering, we can use the same random process with a disjoint palette to decouple the randomness of the pre-shattering and post-shattering phase.
The Quest for a Sublogarithmic LLL Framework.
We solve several LLLs by a combination of the mitigations above. Developing a general black-box framework that can solve all distributed LLLs of this form in sublogarithmic time remains an open challenge. Determining precisely which classes of LLL instances admit such runtimes is, in our view, one of the central open problems in the distributed complexity of the Lovász Local Lemma.
Open Problem 1.
Does there exist a sufficiently large constant $c$ such that all LLLs with dependency degree $d$ and local failure probability $p\leqslant d^{-c}$ admit a $\operatorname{\text{{\rm poly}}}(\log\log n)$ round algorithm?
Open Problem 2.
Does there exist a constant $d_{0}$ such that all LLLs with dependency degree $d\geqslant d_{0}$ and local failure probability $p\leqslant\exp{-\Omega(d^{\varepsilon})}$ for some constant $\varepsilon\in(0,1)$ admit a $\operatorname{\text{{\rm poly}}}(\log\log n)$ round algorithm?
3 Preliminaries
Graphs.
For a graph $G=(V,E)$, we denote by $\Delta(G)$ its maximum degree. For a set $S\subseteq V$, denote by $G[S]$ the subgraph of $G$ induced by $S$. The neighborhood of a vertex in a graph $G$ is denoted by $N_{G}(v)=\{u\in V\mid\{u,v\}\in E(G)\}$, and, by extension, for any $X\subseteq V$, we write $N_{G}(X)=\bigcup_{v\in X}N(v)$ for the set of vertices adjacent to some vertex in $X$. We also use $N_{G}^{\leqslant t}(X)$ to denote the set of vertices within distance $t$ of a vertex in $X\subseteq V$. For a vertex $v\in G$ we denote the degree of $v$ into $H$ by $\deg_{H}(v)=|N_{G}(v)\cap V(H)|$. When graphs are clear from the context we omit the respective indices. For $S\subseteq V$, we write GSG-S for the graph $G[V\setminus S]$ induced by the vertices outside of $S$. For a vertex $v$ and set $S\subseteq V$, define $\operatorname{dist}(v,S)=\min_{u\in S}\operatorname{dist}(v,u)$. If $\mathcal{A}$ is a collection of sets of vertices, then let $\operatorname{dist}(v,\mathcal{A})=\min_{S\in\mathcal{A}}\operatorname{dist}(v,S)$.
Colorings.
For an integer $c\geqslant 1$ we denote $c]=\{1,\ldots,c\}$. A partial $c$-coloring is a function $\varphi:V\to[c]\cup{\bot}$ such that for all $\{u,v\}\in E$ are such that $\varphi(u)\neq\varphi(v)$ unless $\varphi(u)=\bot$ or $\varphi(v)=\bot$. The domain of $\varphi$ is the set of colored vertices: $\operatorname{dom}\varphi=\{u\in V:\varphi(u)\neq\bot\}$. Given a coloring $\varphi$, a color is available to $v\in V$ if none of its neighbors have that color under $\varphi$. The set of available colors, or palette, is denoted by $L(v)=L_{\varphi}(v)=[c]\setminus\{\varphi(u)\mid u\in N(v)\}$. The slack of a vertex $v$ with respect to a coloring $\varphi$ and induced subgraph $H\subseteq G$ is the difference between the number of available colors and the number of uncolored neighbors in $H$, i.e., the slack is $s_{H,\varphi}(v)=|L_{\varphi}(v)|-|N_{H}(v)\setminus\operatorname{dom}\varphi|$. We emphasize that the slack is always defined with respect to a subgraph $H$ and that it does not decrease as we color additional vertices of $H$.
Probabilities.
When we say that an event holds “with high probability in nn”, abridged to w.h.p., we mean that it holds with probability at least $1-n^{-c}$ for some desirably large constant $c>0$. We often omit to mention $n$ explicitly as every such statement in this paper holds w.h.p. in $n$, where $n$ is the number of vertices in the $\mathsf{LOCAL}$ network.
Assumption on $\Delta$.
As stated in Theorem 1, we assume that $\Delta\geqslant\Delta_{0}$ where $\Delta_{0}$ is a sufficiently large universal constant; we frequently use this property throughout the paper to simplify inequalities.
3.1 Lovász Local Lemma
Let $(\Omega_{i},\mathbb{P}_{i})_{i=1}^{n}$ be probability spaces and $X_{i}:\Omega_{i}\to\mathbb{R}$ a family of mutually independent random variables for all $i\in[n]$. Let $\mathcal{B}=\{B_{1},B_{2},\dots,B_{m}\}$ be a family of undesirable (bad) events in $\Omega=\prod_{i=1}^{n}\Omega_{i}$ and, for all $i\in[n]$, let $\operatorname{var}(B_{i})\subseteq[n]$ be sets such that each indicator random variable $1(B_{i})$ is a function of $\{X_{j}:j\in\operatorname{var}(B_{i})\}$. Two events $B_{i},B_{j}$ are adjacent in the dependency graph if $\mathrm{var}(B_{i})\cap\mathrm{var}(B_{j})\neq\emptyset$. The classical Lovász Local Lemma (LLL) [2, Chapter 5] states that if there exist parameters $p<1$ and $d$ such that every event $B_{i}$ satisfies ${\mathbb{P}}[B_{i}]\leqslant p$, and $ep(d+1)\leqslant 1$, where $d$ is the maximum degree of the dependency graph, then there exists an assignment of the variables $X_{i}$ that avoids all events in $\mathcal{B}$.
In the distributed Lovász Local Lemma [17], the input graph is the dependency graph: vertices are events of $\mathcal{B}$ and two events are connected by an edge iff they share a variable. The objective is for all nodes to collaboratively find an assignment of the variables that avoids every event in $\mathcal{B}$, using as few communication rounds as possible. In most applications, the input graph is not the dependency graph itself, but a round of communication on the dependency graph can be simulated by $O(1)$ rounds of $\mathsf{LOCAL}$ on the input graph.
In the context of our coloring algorithms, the random variables correspond to each node’s random choices such as its selected candidate color or whether it participates in a given trial. The bad events are the undesired outcomes of these choices, for example, for each clique there is a bad event that holds if its color coverage property is violated.
Theorem 3 (Deterministic LLL in $\mathsf{LOCAL}$, [50, 26]).
There is a constant $\varepsilon>0$ for which the following holds. There is a deterministic $\mathsf{LOCAL}$ algorithm for the constructive Lovász Local Lemma with $n$ events under criterion $epd(1+\varepsilon)<1$ that runs in $O(\log^{*}s)+\widetilde{O}(\log^{3}n)$ rounds if node IDs are from a space of size $s$.
Theorem 3 is proven by using the powerful general derandomization framework of [50, 27, 31] for the algorithm of Moser-Tardos [46] and using the fastest $\widetilde{O}(\log^{2}n)$ algorithm for computing the required network decompositions [26].
3.2 Our Shattering Framework
Throughout this paper, we use LLL instances to compute certain good vertex labels, e.g., colors. To solve LLL instances in sublogarithmic time, we use the shattering technique [5]. We describe in this section the general approach that we follow throughout this paper.
Sublogarithmic LLL Algorithm.
Every LLL that we solve in this paper follows the same general approach: each vertex $v$ of the graph to be colored has a random variable $X_{v}$ (typically a random color) and the vertex is involved in $\operatorname{\text{{\rm poly}}}(\Delta)$ many events. Each event occurs with probability at most $\exp(-\Delta^{1/40})$, and hence when $\Delta\geqslant(\log n)^{50}$, it suffices to sample the $X_{v}$ uniformly to obtain a globally satisfying assignment with high probability.
When $\Delta$ is small, we split the random process as follows:
  1. I:
    In a pre-shattering random process, each vertex $v$ samples its random variable $X_{v}$ (typically a random color). We have a collection $\mathcal{A}$ of bad events that can then occur.
  2. II:
    We retract the variables held by the vertices $v$ within $O(1)$ hops in $G$ (the variable graph) of a vertex used by an occurring event in $\mathcal{A}$. The exact meaning of retracting $X_{v}$ depends on the concrete LLL instance but generally means that the value of $X_{v}$ will be ignored for the final output.
  3. III:
    In a second post-shattering random process, we sample variables $Y_{v}$ for a subset of the vertices that includes at least all the retracted vertices. For this step, we also have a set of bad events $\mathcal{B}$ that must form an LLL. A satisfying assignment of the $Y_{v}$ avoiding all bad events of the LLL is computed using the deterministic algorithm of Theorem 3.
Each vertex $v$ produces its final value (e.g., color) by combining its $X_{v}$ and $Y_{v}$ values, usually as the value of $Y_{v}$ if it was defined and $X_{v}$ otherwise. For this scheme to succeed and run fast, it is essential that:
  1. 1.
    Every event in $\mathcal{A}$ occurs with low probability,
  2. 2.
    The connected components of the dependency graph of $\mathcal{B}$ have $\operatorname{\text{{\rm poly}}}(\log n)$ size, and
  3. 3.
    The events $\mathcal{B}$ form an LLL.
Our analysis focuses on proving those three points. In general, the round complexity of our pre-shattering and retraction phases is constant and the overall round complexity is dominated by applying Theorem 3 on the connected components of the dependency graph of $\mathcal{B}$. The runtime of $\operatorname{\text{{\rm poly}}}(\log\log n)$ then follows from point 2, as the components have size $N=\operatorname{\text{{\rm poly}}}(\log n)$ and Theorem 3 ends after $\operatorname{\text{{\rm poly}}}(N)$ rounds. To prove point 2, we use a shattering argument.
The Shattering Lemma.
The following lemma, initially introduced by [5], shows that when vertices of a graph of maximum degree $\Delta$ get selected with probability $1/\operatorname{\text{{\rm poly}}}(\Delta)$, the connected components of the selected vertices are small. Recall that this is only used when $\Delta\leqslant\operatorname{\text{{\rm poly}}}(\log n)$, so connected components have size at most $\operatorname{\text{{\rm poly}}}(\log n)$. We use the following general purpose shattering statement.
Lemma 3.1 (Shattering Lemma [23]).
Let $G=(V,E)$ be a graph with maximum degree $\Delta$. Consider a process that generates a random subset $B\subseteq V$ such that ${\mathbb{P}}[v\in B]\leqslant\Delta^{-C_{1}}$, for some constant $C_{1}\geqslant 1$, and such that the random variables $\bm{1}(v\in B)$ depend only on the randomness of nodes within at most $C_{2}$ hops from $v$, for all $v\in V$, for some constant $C_{2}\geqslant 1$. Then, for any constant $C_{3}\geqslant 1$, satisfying $C_{1}>C_{3}+4C_{2}+2$, we have that any connected component in $G[B]$ has size at most $O(\Delta^{2C_{2}}\log_{\Delta}n)$ with probability at least $1-n^{-C_{3}}$.
To avoid technical redundancies, we use the following shattering lemma instead of Lemma 3.1. It takes advantage of the fact that, in all our application, our LLLs have one random variable per vertex. So we consider two LLLs $\mathcal{A}$ and $\mathcal{B}$ whose variables are indexed by the vertices of a graph $G$ called the variable graph. The set $\mathcal{A}^{\prime}$ contains all the bad events that occur after the pre-shattering step and whose variables are retracted. Every event $B\in\mathcal{B}$ with a variable within distance $c_{\mathcal{B}^{\prime}}$ in $G$ of a retracted variable, i.e., in $\operatorname{vbl}(\mathcal{A}^{\prime})$, is included in the post-shattering LLL which we call $\mathcal{B}^{\prime}$. Lemma 3.2 states that the dependency graph $G_{\operatorname{dep}}$ of the post-shattering LLL $\mathcal{B}^{\prime}$ has small connected components if $\operatorname{vbl}(B)$ is locally embedded in the variable graph $G$ for every event $B\in\mathcal{A}\cup\mathcal{B}$. This is formalized by Steps (S1) and (S2).
For simplicity, we state Lemma 3.2 in terms of collections of sets $\mathcal{A},\mathcal{A}^{\prime},\mathcal{B},\mathcal{B}^{\prime}$ instead of LLLs. When we use Lemma 3.2, we replace every bad event $B$ with the set of vertices in $G$ that hold its variables. The fact that $\mathcal{A}^{\prime}$ contain rare bad events from an LLL is formalized by Steps (S4) and (S3).
Lemma 3.2.
Let $G=(V,E)$ be an $n$-vertex graph with maximum degree $\Delta$. Consider $\mathcal{A},\mathcal{B}\subseteq 2^{V}$ two collections of subsets of $V$. Suppose there exist constants $c_{1},c_{2},c_{4}\geqslant 1$ and $c_{\mathcal{B}^{\prime}}\geqslant 0$ with $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1$ such that
  1. (S1)
    for all $S\in\mathcal{A}\cup\mathcal{B}$, the set $S$ has weak-diameter at most $c_{1}$ in $G$, and
  2. (S2)
    for all $v\in V$, there are at most $\Delta^{c_{2}}$ sets $S\in\mathcal{A}\cup\mathcal{B}$ such that $v\in S$.
Let $X_{v}$ be independent random variables indexed by the vertices of $G$. Consider a random process on the corresponding probability space that produces a collection $\mathcal{A}^{\prime}\subseteq\mathcal{A}$ such that, for all $A\in\mathcal{A}$,
  1. (S3)
    the random variable $\bm{1}(A\in\mathcal{A}^{\prime})$ is a function of the $\{X_{v}:v\in A\}$, and
  2. (S4)
    ${\mathbb{P}}[A\in\mathcal{A}^{\prime}]\leqslant\Delta^{-c_{4}}$ .
Define
$\mathcal{B}^{\prime}=\{B\in\mathcal{B}:\exists v\in B,\operatorname{dist}_{G}(v,\mathcal{A}^{\prime})\leqslant c_{\mathcal{B}^{\prime}}\}\ .$
Let $G_{\operatorname{dep}}$ be the (random) graph on vertex set $\mathcal{B}^{\prime}$ with edges between $B$ and $B^{\prime}$ iff $B\cap B^{\prime}\neq\emptyset$. Then, the connected components of $G_{\operatorname{dep}}$ have size at most $O(\Delta^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}+c_{2}}\log n)$, with high probability in $n$.
Proof.
Let $H=G^{2c_{1}}$, namely the graph on vertex set $V$ with an edge $\{u,v\}$ iff $\operatorname{dist}_{G}(u,v)\leqslant 2c_{1}$. For every $B\in\mathcal{B}^{\prime}$, let $f(B)\in B$ be the vertex such that $\operatorname{dist}_{G}(f(B),\mathcal{A}^{\prime})\leqslant c_{\mathcal{B}^{\prime}}$, which exist by definition of $\mathcal{B}^{\prime}$. Observe that $f$ is an adjacency-preserving map from $G_{\operatorname{dep}}$ to $H[N_{G}^{\leqslant c_{\mathcal{B}^{\prime}}}(\mathcal{A}^{\prime})]$: by (S1), if $B\cap B^{\prime}\neq\emptyset$ (i.e., they are adjacent in $G_{\operatorname{dep}}$), then $f(B)$ and $f(B^{\prime})$ are connected by an edge in $H$. Let $C$ be the largest connected component of $G_{\operatorname{dep}}$. The image of $C$, denoted $f(C)$, is a connected subgraph of $H[N_{G}^{\leqslant c_{\mathcal{B}^{\prime}}}(\mathcal{A}^{\prime})]$. On the other hand, by (S2) each vertex of $H$ is the image of at most $\Delta^{c_{2}}$ vertices of $C$ through $f$, and so $|C|\leqslant\Delta^{c_{2}}|f(C)|$. It therefore suffices to upper bound the size of the largest connected component of $H[N_{G}^{\leqslant c_{\mathcal{B}^{\prime}}}(\mathcal{A}^{\prime})]$.
By (S2), (S4), and the union bound over all sets $A$ with $v$ as a vertex, we have that
${\mathbb{P}}\left[v\in\bigcup_{A\in\mathcal{A}^{\prime}}A\right]\leqslant\Delta^{c_{2}-c_{4}}\ .$
By the union bound over the vertices of $N_{G}^{\leqslant c_{\mathcal{B}^{\prime}}+1}(v)$, we deduce that
${\mathbb{P}}\left[\operatorname{dist}_{G}(v,\mathcal{A^{\prime}})\leqslant c_{\mathcal{B}^{\prime}}\right]\leqslant\Delta^{c_{\mathcal{B}^{\prime}}+1}\cdot\Delta^{c_{2}-c_{4}}\leqslant\Delta(H)^{(c_{\mathcal{B}^{\prime}}+1+c_{2}-c_{4})/(3c_{1})}\ .$
By (S3), whether $A\in\mathcal{A}^{\prime}$ depend only on the random variables $X_{u}$ of vertices in $A$, which has diameter one in $H$. Whether $v\in N_{G}^{\leqslant c_{\mathcal{B}^{\prime}}}(\mathcal{A}^{\prime})$ depends on whether $A\in\mathcal{A}^{\prime}$ for all the $A$ with a vertex in $N^{\leqslant c_{\mathcal{B}^{\prime}}}_{G}(v)$, hence it depends only on the random values within $c_{\mathcal{B}^{\prime}}+1$ hops in $H$. Let $C_{1}=(c_{4}-c_{\mathcal{B}^{\prime}}-c_{2}-1)/(3c_{1})$, $C_{2}=c_{\mathcal{B}^{\prime}}+1$ and $C_{3}=10$; it is easy to verify that $C_{1}>C_{3}+4C_{2}+2$ for our choice of $c_{1},c_{2},c_{4},c_{\mathcal{B}^{\prime}}$. By Lemma 3.1, w.h.p., all the components of $H[N^{\leqslant c_{\mathcal{B}^{\prime}}}_{G}(\mathcal{A}^{\prime})]$ have size at most $O(\Delta(H)^{2c_{\mathcal{B}^{\prime}}+2}\log n)=O(\Delta^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}}\log n)$, which implies the claimed bound on the size of the largest connected component in $G_{\operatorname{dep}}$.
4 Top-Level Algorithm and Proofs of Theorems 1 and 2
The goal of this section is to present the top-level algorithm for proving Theorems 1 and 2. See 1
We assume that each vertex knows its incident edges and the values of $\Delta$ and $c$.
The general problem reduces to the coloring of a graph $F$ with a highly specific structure (see Theorem 5 in [45]). In Section 4.1, we present a structural decomposition of $F$ that forms the base of our coloring algorithm. In Section 4.2, we detail the crucial color coverage property that we need to maintain throughout our algorithm in order to find suitable color swaps when completing the coloring for the cliques of $F$. In Section 4.3, we present the proof of Theorem 1, building on Lemmas 4.5, 4.7 and 4.8 that state how fast we can color the various parts of $F$. These lemmas are proven in Sections 5, 6 and 7.
4.1 Graph Decomposition
The coloring problem of Theorem 1 reduces to coloring a graph $F$ with a specific structure. This structural decomposition, introduced in [45] and efficiently computable in $\mathsf{LOCAL}$ by [4], is technically involved. Since later proofs rely on several additional structural properties, their original presentation becomes cumbersome. We therefore adapt the decomposition to our setting and restate it in the following (still technical) lemma, including all necessary properties.
Lemma 4.1 (Structural Decomposition).
There is a decomposition of $F$ into sets of vertices $S$, $B_{H}$, and $B_{L}$, and into two collections of cliques $A_{H}$ and $A_{L}$ that are colored in the order presented below (see Algorithm 1). Additional properties that the sets satisfy independently of the coloring in previous steps are stated afterwards.
  1. 1.
    $S$: Each $v\in S$ has $\deg_{S}(v)<\Delta-3\sqrt{\Delta}$ or it has $\deg_{S}(v)\leqslant\Delta$ and at least $9\cdot 10^{5}\cdot\Delta^{3/2}$ non-adjacent pairs of neighbors within $S$.
  2. 2.
    $B_{H}:$ Each $v\in B_{H}$ satisfies $|L(v)|\geqslant\deg_{B_{H}}(v)+\Delta^{3/4}$, regardless of how $S$ is colored.
  3. 3.
    $A_{H}:$ The external degree (cf. Definition 4.2) of each $A_{i}\in A_{H}$ is bounded by $10^{8}\sqrt{\Delta}$.
  4. 4.
    $B_{L}:$ Each $v\in B_{L}$ satisfies $|L(v)|\geqslant\deg_{B_{L}}(v)+\tfrac{1}{2}\sqrt{\Delta}$, regardless of how $S,B_{H},A_{H}$ are colored.
  5. 5.
    $A_{L}$: The external degree of each $A_{i}\in A_{L}$ is bounded by $30\Delta^{1/4}$ .
Additionally we have:
  1. (a)
    Every $A_{i}$ is a clique with $c-10^{8}\sqrt{\Delta}\leqslant|A_{i}|\leqslant c$.
  2. (b)
    $\mathrm{All}_{i}\subseteq B_{L}\cup B_{H}$ Vertices in $\mathrm{All}_{i}\subseteq B$ are adjacent to all of $A_{i}$. $|\mathrm{All}_{i}|=c-|A_{i}|$ holds.
  3. (c)
    $\operatorname{Big}_{i}^{+}\subseteq B_{L}\cup B_{H}$: A vertex $v\notin A_{i}\cup\mathrm{All}_{i}$ lies in $\operatorname{Big}_{i}^{+}$ if it has at least $2\Delta^{9/10}$ neighbors in $A_{i}$. The set $\operatorname{Big}_{i}^{+}$ is a clique, and each $v\in\operatorname{Big}_{i}^{+}$ has at most $\tfrac{3}{4}\Delta+10^{8}\sqrt{\Delta}$ neighbors in $A_{i}$.
Definition 4.2.
The external degree of a clique $A_{i}$ is the maximum number of neighbors of a $v\in A_{i}$ outside of $A_{i}\cup\mathrm{All}_{i}$.
The proof of Lemma 4.1 appears in Appendix A.
4.2 The Color Coverage (CC) Property
In this section, we formally introduce the color coverage property, which we abbreviate CC, that was outlined in Section 2. Informally speaking we wish to ensure that the coloring outside of cliques $A_{i}$ looks random enough. Formally, we ask that each color is available to a constant fraction of the nodes on the inside of the cliques.
While $\mathrm{All}_{i}$ contains the nodes that are connected to all nodes in $A_{i}$, the nodes in $\operatorname{Big}_{i}^{+}$ are connected to most nodes (but still not too many) of $A_{i}$. Hence, coloring a single node in $\operatorname{Big}_{i}^{+}$ may bring $A_{i}$ close to not satisfying CC. But due to Lemma 4.1 the nodes in $\operatorname{Big}_{i}^{+}$ form a clique for $A_{i}$; hence, no two nodes in $\operatorname{Big}_{i}^{+}$ can receive the same color.
Property CC.
Consider some arbitrary step $j$ of a coloring algorithm. Let $\text{CC}_{j}(i,x)$ be the number of vertices in $A_{i}$ that have a neighbor $v\not\in A_{i}\cup\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ that gets colored with $x$ in step $j$. We say that CC is within budget for a single coloring step if $\text{CC}_{j}(i,x)<\Delta^{37/40}$ holds for every uncolored clique $A_{i}$ and for every color $x$. We say that CC holds for a clique $A_{i}$ at a point in time if $C$ has been within budget for every coloring step applied so far.
Essentially, the $\Delta^{37/40}$ budget ensures that across all steps, each color remains sufficiently under-used around every clique, enabling valid swaps later. The under-usage across all steps is formalized in the following observation.
Observation 4.3.
If we maintain CC throughout the coloring process for a clique $A_{i}$, then for all colors $x$ and at any point in the coloring process, at most $4\Delta/5$ vertices of $A_{i}$ have a neighbor outside of $A_{i}\cup\mathrm{All}_{i}$ with color $x$.
Proof.
Our algorithm has either $O(\log^{*}\Delta)$ or $O(\log\Delta)$ coloring steps [^6] In several of our coloring steps we compute a partial coloring by setting up an LLL that we solve via the shattering framework. That means that we color vertices in a pre-shattering and a post-shattering phase, both of which count as a separate coloring step in the context of this observation. In other words, the pre-shattering and post-shattering phases in these LLLs obtain a separate CC budget which simplifies the analysis significantly. , depending on the approach. Summing up the $\Delta^{37/40}$ budget of every individual step, results in $O(\log\Delta\cdot\Delta^{37/40})$ vertices in $A_{i}$ that have a $v\not\in A_{i}\cup\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ that is permanently colored with color $x$ (in any previous step). Consider a fixed color $x$. By Lemma 4.1-(c), $\operatorname{Big}_{i}^{+}$ forms a clique, and hence at most one node of $\operatorname{Big}_{i}^{+}$ can be colored $x$, and this vertex has at most $\tfrac{3}{4}\Delta+10^{8}\sqrt{\Delta}$ neighbors in $A_{i}$. Overall, the number of vertices $v\not\in A_{i}\cup\mathrm{All}_{i}$ adjacent to a vertex colored $x$ is bounded above by
$O(\log\Delta\cdot\Delta^{37/40})+\tfrac{3}{4}\Delta+10^{8}\sqrt{\Delta}\leqslant 4\Delta/5\penalty 10000\ .$
Bounding $\text{CC}_{j}(i,x)$.
Suppose that we have a collection of at most $\Delta$ subsets of $V(F)$. Each set contains at most $Q$ vertices. No vertex lies in more than $2\Delta^{9/10}$ sets. We conduct a random experiment where each vertex is marked with probability at most $1/(Q\times\Delta^{1/5})$. The vertices are not necessarily marked independently, but the experiment has the following property [^7] We remark that the exact value of the constant $1/5$ is not crucial. We can choose it as small as we like, but then the bound $\Delta^{37/40}$ inches closer to $\Delta$ and the exponent in the error probability gets smaller.:
(P7.1) For any set of $\ell\geqslant 1$ vertices, the probability that all are marked is at most $1/(Q\times\Delta^{1/5})^{\ell}$.
We use the following Lemma 32 of [45] verbatim. Note that condition (P7.1) does not require independence between vertices of different sets, and also not for vertices inside the same set.
Lemma 4.4 (Lemma 32 of [45]).
The probability that at least $\Delta^{37/40}$ sets contain at least one marked vertex is at most $\exp(-\Delta^{1/40})$.
The high level idea of using Lemma 4.4 to show that CC is maintained for cliques in each step of our coloring procedure is as follows. Most of our coloring steps are based on some variant of random color trials, in which each node picks one (or multiple) candidate colors and then permanently retains one of its candidate colors if none of its neighbors chose it as a candidate. To prove that such a process maintains CC, fix a color $x\in[c]$, a clique $A_{i}$, and define a set $N_{v}$ for each vertex $v$ of the clique containing its external neighbors, except those in $\operatorname{Big}_{i}^{+}$. Let $Q$ be an upper bound on the sizes of the sets. As we exclude vertices in $\operatorname{Big}_{i}^{+}$ (those with more than $2\Delta^{9/10}$ neighbors in $A_{i}$) no vertex lies in more than $2\Delta^{9/10}$ sets. We consider a vertex $u\in\bigcup N_{v}$ marked if color $x$ is among its candidate colors. Clearly, the final colors picked by vertices are not independent, but the choices of candidate colors is. Hence, if the list of each vertex is of size $Q\cdot\Delta^{1/5}$ the properties of Lemma 4.4 are satisfied. We obtain that CC is maintained for color $x$ and clique $A_{i}$ with probability $\exp(-\Delta^{1/40})$.
Note that in this process the events whether color $x$ appears in the neighborhood of two vertices $v\neq v^{\prime}$ may not be independent as $v$ and $v^{\prime}$ may have a common external neighbor. The strength of Lemma 4.4 is that it can still deal with this situation.
4.3 Full Algorithm and Proofs of Theorems 1 and 2
The proof of Theorem 1 appears at the end of this section. We first focus on coloring the graph $F$ from which we can efficiently recover a coloring of $G$. See Algorithm 1 for pseudocode and the order in which we process vertices of $F$.
Input : Graph $F$ from Lemma 4.1 with the vertex-partition $S,B_{H},A_{H},B_{L},A_{L}$
Output : Coloring of $F$
1
21exColorSparse($S$)
3ColorWithMuchSlack($B_{H}$)
4ColorCliques($A_{H}$)
5ColorWithMuchSlack($B_{L}$)
ColorCliques($A_{L}$)
Algorithm 1 Overall $\Delta-k$ Coloring Algorithm
The guarantees required and provided by subroutines ColorWithMuchSlack, ColorSparse, and ColorCliques are stated in Lemmas 4.5, 4.7 and 4.8 below, and proven in Sections 5, 6 and 7 respectively. We begin with the definition of $\Pi$-ous subgraphs that are central for coloring all non-clique vertices. It is used to show that any probabilistic coloring of these vertices is random enough compared to the external degree of the still uncolored cliques.
$\Pi$-ous subgraphs.
A subgraph $H$ is $\Pi$-ous if it satisfies the following property:
Property $\Pi$.
An induced subgraph $H$ of $F$ satisfies property $\Pi$ if there exists $U\geqslant\Delta^{1/4}$ s.t. (a) every uncolored vertex in each clique $A_{i}$ has at most $U$ neighbors in $H-\mathrm{All}_{i}$, and (b) each vertex $u\in H$ has a list of available colors satisfying $|L(u)|\geqslant\deg_{H}(u)+U\cdot\Delta^{0.22}$.
Property (b) ensures that every vertex in $H$ has slack $U\cdot\Delta^{0.22}$, and importantly this slack remains as other vertices of the subgraph $H$ are colored. This allows us to color $H$ with $O(\log\Delta)$ iterated random color trial while maintaining Property CC. In Section 5 we prove the following lemma for coloring $\Pi$-ous subgraphs.
Lemma 4.5 (ColorWithMuchSlack).
Let $H$ be a $\Pi$-ous subgraph of $F-\bigcup_{i}A_{i}$. There is a $\mathsf{LOCAL}$ algorithm that, w.h.p., extends the coloring to all the vertices of $H$ in $\widetilde{O}(\log\Delta\cdot\log^{3}\log n)$ rounds while maintaining Property CC.
Note that when $\Delta=n^{\varepsilon}$, Lemma 4.5 does not guarantee a sublogarithmic runtime. For $\Delta\geqslant(\log n)^{50}$, we adapt the state-of-the-art $(\deg+1)$-list-coloring algorithm of [35] for coloring a $\Pi$-ous subgraph while maintaining CC; see Section 8 for details. Using Lemma 4.5 for low-degree graphs incurs an additional $O(\log\Delta)=O(\log\log n)$ factor in the runtime, but greatly simplifies the analysis.
Lemma 4.6.
Let $H$ be a $\Pi$-ous subgraph of $F-\bigcup_{i}A_{i}$ and assume $\Delta\geqslant(\log n)^{50}$. There is a $\mathsf{LOCAL}$ algorithm that list-colors all vertices of $H$ in $O(\log^{*}n)$ rounds while maintaining Property CC.
We cannot use Lemma 4.5 directly to color the vertices of $S$ because they do not have enough initial slack to satisfy $\Pi$. As in [45], we begin by solving an LLL to provide them with $\Omega(\sqrt{\Delta})$ slack which is still not sufficient for $\Pi$. We then color the vertices of $S$ with Lemma 4.5 (or Lemma 4.6) in two batches to ensure that $\Pi$ holds, thereby that CC is maintained. Details can be found in Section 6.
Lemma 4.7 (ColorSparse).
Let $c\geqslant\Delta-k_{\Delta}+1$ and $S$ be the set of vertices in $F$ as in Lemma 4.1-(1). There is a $\widetilde{O}(\log^{4}\log n)$-round distributed algorithm that, w.h.p., $c$-colors the vertices of $S$ while maintaining CC. When $\Delta\geqslant(\log n)^{50}$, it runs in $O(\log^{*}n)$ rounds.
As explained above, to color the cliques, we require that they satisfy CC. Since the cliques from $A_{L}$ are colored later, we must be careful about maintaining CC for those as we color $A_{H}$. We prove Lemma 4.8 in Section 7.
Lemma 4.8 (ColorCliques).
Let $c\geqslant\Delta-k_{\Delta}+1$ and let $A^{\prime}$ be a subset of the cliques $A_{i}$ from Lemma 4.1 such that
  1. 1.
    all $A_{i}\in A^{\prime}$ satisfy CC, and
  2. 2.
    every uncolored vertex in some $A_{i}\notin A^{\prime}$ has at most $30\Delta^{1/4}$ external neighbors.
Then, there is a $\widetilde{O}(\log^{3}\log n)$-round $\mathsf{LOCAL}$ algorithm that, w.h.p., $c$-colors all the cliques in $A^{\prime}$ while maintaining CC for all the uncolored cliques outside $A^{\prime}$. For $\Delta\geqslant(\log n)^{50}$, it runs in $O(1)$ rounds.
Proof of Theorem 1.
Building on [45, Theorem 5], Bamas and Esperet [4, Theorem 4.1] show that if a graph $G$ is not $c$-colorable for $c\geqslant\Delta-k_{\Delta}+1$, then there exists a vertex $v\in V(G)$ such that the graph induced by $\{v\}\cup N(v)$ has chromatic number >c>c. This can be tested in $O(1)$ rounds in the $\mathsf{LOCAL}$ model exploiting the unbounded local computations.
[…]
For the rest of the proof we focus on computing a $c$-coloring of $F$ by coloring the nodes in $S$, $B_{H}$, $A_{H}$, $B_{L}$, and $A_{L}$ in the prescribed order (see Algorithm 1). We first focus on the case when $\Delta\leqslant(\log n)^{50}$; the case of larger $\Delta$ is handled thereafter.
Coloring S. The sparse nodes are colored via Lemma 4.7.
Coloring $B_{H}$. We argue that the graph induced by $B_{H}$ is $\Pi$-ous with $U=U_{H}=10^{8}\sqrt{\Delta}$, and color it with Lemma 4.5 when $\Delta\leqslant(\log n)^{50}$ and with Lemma 4.6 otherwise. By Lemma 4.1-(2), each vertex of $B_{H}$ has $\Delta^{3/4}\geqslant U\Delta^{1/4}=\Theta(\Delta^{0.72})$ slack in $F[B_{H}]$ regardless of how nodes in $S$ are colored. All cliques are uncolored at this point and each vertex in each $A_{i}$ has at most $U_{H}$ neighbors in $B_{H}$ by Lemma 4.1-(3,5).
Coloring $A_{H}$. The collection of cliques in $A_{H}$ is colored via Lemma 4.8. We can apply the lemma as their CC was maintained by earlier coloring steps (Lemmas 4.7, 4.5 and 4.6). We emphasize that, by Lemma 4.8, CC of cliques in $A_{L}$ still holds after coloring cliques in $A_{H}$.
Coloring $B_{L}$. We argue that $F[B_{L}]$ is $\Pi$-ous subgraph with $U=U_{L}=30\Delta^{1/4}$ and color it via Lemma 4.5 when $\Delta\leqslant(\log n)^{50}$ and via Lemma 4.6 otherwise. By Lemma 4.1-(5), the external degree of cliques in $A_{L}$ (cliques in $A_{H}$ are already colored at this point) is bounded above by $U_{L}$. By Lemma 4.1-(4), each node of $B_{L}$ has at least $\sqrt{\Delta}/2$ slack in $F[B_{L}]$, which is larger than the $U_{L}\cdot\Delta^{0.22}=30\Delta^{0.47}$ slack required by $\Pi$.
Coloring $A_{L}$. The cliques in $A_{L}$ are also colored via Lemma 4.8, which can be done as CC holds for each of these cliques. Note that Lemma 4.8-(2) vacuously holds at this step because all the uncolored cliques are in $A^{\prime}=A_{L}$.
Runtime.
The runtime is dominated by the time required to color $S$, $B_{H}$ and $B_{L}$. For $\Delta\leqslant(\log n)^{50}$, it is $\widetilde{O}(\log\Delta\cdot\log^{3}\log n)=\widetilde{O}(\log^{4}\log n)$ by Lemmas 4.5 and 4.7, while for $\Delta\geqslant(\log n)^{50}$, it is $O(\log^{*}n)$ by Lemmas 4.6 and 4.7.
See 2
Proof of Theorem 2.
If the graph $G$ is not $c$-colorable, there exists a vertex $v$ for which $\{v\}\cup N(v)$ is not $c$-colorable [45, 4], which can be deterministically detected in $O(1)$ rounds of $\mathsf{LOCAL}$. When $G$ is $c$-colorable, we apply the derandomization framework developed in [31, 27, 50]. It states that any randomized algorithm with runtime $T(n)$ for a problem whose solution can be verified in $l(n)$ rounds can be derandomized in $O((T(n)+l(n))\cdot T_{ND}(n))$ rounds where $T_{ND}(n)$ is the runtime to compute a so-called $(O(\log n),O(\log n))$-network decomposition, see any of these works for details. Here, it is applied to the randomized algorithm of Theorem 1 with $T(n)=\widetilde{O}(\log^{4}\log n)$ and $l(n)=1$ because checking if the resulting coloring is proper only requires a single round. By using the algorithm of [26] for computing a network decomposition in $\widetilde{O}(\log^{2}n)$ rounds, we obtain Theorem 2.
5 Coloring With Much Slack
The goal of this section is to prove Lemma 4.5, that is, we aim to color an induced subgraph $H$ of $F-\bigcup_{i}A_{i}$ that is $\Pi$-ous while at the same time maintaining CC for all uncolored cliques. Recall that $H$ is $\Pi$-ous if each node has sufficient slack in $H$ compared to a given upper bound $U\geqslant\Delta^{1/4}$ on the external degree of uncolored cliques. The coloring process consists of two primary components. The first one is a random color trial for all uncolored vertices, applied for $O(\log\Delta)$ iterations, where each iteration lowers the uncolored degree of each vertex by a constant factor in expectation. The properties of the resulting partial coloring after a single color trial are summarized in Lemma 5.1 below. The second component summarized in Lemma 5.2 below is then a multi color trial that colors all remaining vertices of the graph instance to be colored.
Lemma 5.1 (Random Color Trial).
Let $H$ be an uncolored $\Pi$-ous subgraph of $F-\bigcup_{i}A_{i}$ and let $R$ be the graph induced by the remaining uncolored nodes of $H$ after running Algorithm 3. Then, w.h.p.,
  1. 1.
    $\deg_{R}(v)\leqslant\max\{(1-1/180)\deg_{H}(v),\Delta^{1/10}\}$ for all $v\in V(R)\cup\bigcup_{i}A_{i}$, and
  2. 2.
    CC is maintained for all uncolored cliques.
Algorithm 3 runs in $\widetilde{O}(\log^{3}\log n)$ rounds.
Lemma 5.2 (Multi Color Trial).
Let $H$ be an uncolored induced subgraph in which every vertex has $\Delta^{9/20}$ slack and each uncolored vertex in each $A_{i}$ has at most $\Delta^{1/10}$ neighbors in $H$. There is a $\widetilde{O}(\log^{3}\log n)$-round $\mathsf{LOCAL}$ algorithm that, w.h.p., colors $H$ while maintaining CC for every uncolored clique $A_{i}$. When $\Delta\geqslant(\log n)^{50}$, the algorithm runs in $O(1)$ rounds.
Finally, these two components are combined to establish the main lemma of this section. We start the coloring with $O(\log\Delta)$ iterations of the random color trial from Lemma 5.1, which decrease the maximum uncolored degree to $\Delta^{1/10}$. Then, the coloring is completed by a single iteration of multi color trial from Lemma 5.10.
See 4.5
Proof.
We color the graph $H$ which satisfies $\Pi$, by iteratively applying Algorithm 3 for $T=c\log\Delta$ times, where $c>0$ is a sufficiently large constant. We then finish the coloring of $H$ using Lemma 5.2. To analyze the first iterative process, let $H_{j}$ be the subgraph induced by the uncolored vertices of $H$ after the $j$-th call of Algorithm 3, and let $H_{0}=H$. Consider a vertex $v\in V(H)\cup\bigcup_{i}A_{i}$. According to Lemma 5.1, for every iteration $j$, one of the two conditions must hold: either the uncolored degree drops immediately such that $|N_{H_{j}}(v)|\leqslant\Delta^{1/10}$ or the degree drops by a constant factor such that $|N_{H_{j}}(v)|\leqslant(1-1/180)|N_{H_{j-1}}(v)|$. If the first condition holds for some $j$, then $v$ has fewer than $\Delta^{1/10}$ uncolored neighbors in $H$ and already satisfies the required degree bound. We therefore assume that the second condition holds for all $j$. Since the maximum degree of $H$ (as a subgraph of $F$) is at most $10^{9}\Delta$ (by Lemma 4.1), after $T=c\log\Delta$ iterations the uncolored degree of $v$ is at most
$10^{9}\Delta\cdot(1-1/180)^{c\log\Delta}\leqslant 10^{9}\exp\left(\ln\Delta-\frac{c\cdot\log\Delta}{180}\right)\leqslant 10^{9}\Delta^{1-c/180}$
For sufficiently large constant $c$, this is at most $\Delta^{1/10}$. Now let $H^{\prime}$ be the subgraph of the initial graph $H$ that contains all the yet uncolored vertices after $T$ iterations of Lemma 5.1. Then the graph $H^{\prime}$ fulfills the preconditions of Lemma 5.2, which require that every vertex has $\Delta^{9/20}$ slack and each uncolored vertex in each $A_{i}$ has at most $\Delta^{1/10}$ neighbors in the remaining uncolored graph. The application of Lemma 5.2 to $H^{\prime}$ completes the coloring of $H$. The CC is preserved for all uncolored cliques as it is guaranteed by the outcome of Lemmas 5.1 and 5.2. The runtime consists of the $O(\log\Delta)$ iterations from Lemma 5.1 that take $\widetilde{O}(\log^{3}\log n)$ rounds each and $\widetilde{O}(\log^{3}\log n)$ rounds for the one application of Lemma 5.2. This completes the proof.
5.1 Iterated Random Color Trial with CC (Lemma 5.1)
The goal of this section is to prove Lemma 5.1. We present a $\operatorname{\text{{\rm poly}}}(\log\log n)$ algorithm that partially colors the graph $H$ while maintaining CC and additionally reduces the uncolored degree of vertices in $H$ and $\bigcup_{i}A_{i}$ below $\Delta^{1/10}$.
The analysis focuses on a single iteration of Random Color Trial (Rct). We show that in each iteration of the Rct procedure the uncolored degree of a vertex drops by a factor of (11/180)(1-1/180) compared to the previous iteration while maintaining CC. We set up an LLL that, if solved correctly, yields a partial coloring guaranteeing this drop in the uncolored degree. The random process underlying our LLL is described in Algorithm 2; we call the color $\psi(v)$ the candidate color of $v$.
Input : $H^{\prime}$ an induced uncolored subgraph of $H$
Output : A partial coloring of $H^{\prime}$
1 Each vertex of $H^{\prime}$ gets activated independently with probability $p_{a}=1/4$
2 Each activated vertex $v$ independently samples a uniform available color $\psi(v)$
// Let $N_{<}(v)=\{u\in N(v):ID(u)<ID(v)\}$
All vertices $v$ with $\psi(v)\in\psi(N_{<}(v))$ discard $\psi(v)$ (i.e., set $\psi(v)$ to $\bot$)
Algorithm 2 Random Color Trial (Rct)
In Algorithm 2, every vertex of $H^{\prime}$ has one random variable that consists of a pair with a random activation bit, set to one w.p. $p_{a}$, and the uniform available color $\psi(v)$. So we henceforth abuse notation slightly and associate vertices with their random variables. We introduce the following bad event, for every vertex $v$ with at least $\Delta^{1/10}$, define
$E(v)$: the uncolored degree into $H$ of $v$ is reduced by a factor less than (11/180)(1-1/180),
and for each uncolored clique $A_{i}$, we have
$E_{CC}(i)$: CC is not maintained for some color w.r.t. $A_{i}$.
Observation 5.3.
$\operatorname{vbl}(E(v))\subseteq N_{H}^{\leqslant 2}(v)$ and $\operatorname{vbl}(E_{CC}(i))\subseteq N_{F}^{\leqslant 2}(A_{i})\cap V(H)$.
Let us argue now that no matter the coloring of $H$, since it is $\Pi$-ous those bad events are rare.
Lemma 5.4.
Let $H^{\prime}$ be an induced uncolored subgraph of a $\Pi$-ous subgraph $H$ of $F-\bigcup_{i}A_{i}$, and $v\in V(H)\cup\bigcup_{i}A_{i}$ with $\deg_{H^{\prime}}(v)\geqslant\Delta^{1/10}$. After Algorithm 2, $E(v)$ occurs w.p. at most $\exp(-\Omega(\Delta^{1/10}))$.
Proof.
Let $u\in\{v\}\cup N_{H^{\prime}}(v)$ with $\deg_{H^{\prime}}(u)\geqslant\Delta^{1/10}$. Let $X_{u}$ be the random variable that counts the number of activated vertices in the neighborhood $N_{H^{\prime}}(u)$ of $u$. Its expected value is given by $\operatorname{\mathbb{E}}[X_{u}]=p_{a}\cdot\deg_{H^{\prime}}(u)$. We show the concentration of $X_{u}$ by Chernoff’s bound (Proposition B.1):
${\mathbb{P}}\left[|X_{u}-p_{a}\deg_{H^{\prime}}(u)|\geqslant p_{a}\deg_{H^{\prime}}(u)/2\right]\leqslant 2\exp\left(-\frac{p_{a}\cdot\deg_{H^{\prime}}(u)}{12}\right)\leqslant\exp\left(-\Omega(\Delta^{1/10})\right)\ .$
where the last inequality uses the lower bound on the degree of $u$. Let $E^{\mathrm{all}}(v)$ be the event that every vertex $u\in\{v\}\cup N_{H^{\prime}}(v)$ with $\deg_{H^{\prime}}(u)\geqslant\Delta^{1/10}$ has between $(p_{a}/2)\deg_{H^{\prime}}(u)$ and $2p_{a}\deg_{H^{\prime}}(u)$ active neighbors. Since $\overline{E^{\mathrm{all}}(v)}$ is a low-probability event, we assume $E^{\mathrm{all}}(v)$ holds.
Let $u_{1},u_{2},\ldots,u_{d}$ be the active neighbors of $v$ ordered by increasing IDs. Let $Z_{i}$ be the indicator random variable equal to one iff $u_{i}$ gets colored by Algorithm 2. It is uniquely determined by the random variables $R_{k}:=\psi(N_{<}(u_{k})\setminus\bigcup_{j<k}N_{<}(u_{j}))$ for all $k\leqslant i$. We claim that
$\operatorname{\mathbb{E}}\left[Z_{i}\nonscript\>|{}\nonscript\>\mathopen{}E^{\mathrm{all}}(v),R_{1},R_{2},\ldots,R_{i-1}\right]=\mu_{i}\geqslant 1/2\ .$
There are two cases. If $\deg_{H^{\prime}}(u)<\Delta^{1/10}$, then since $H^{\prime}$ is $\Pi$-ous, $u$ has at least $\Delta^{1/4}$ available colors. Hence, it gets colored with probability at least $1-\Delta^{-3/20}\geqslant 1/2$. If $\deg_{H^{\prime}}(u)\geqslant\Delta^{1/10}$, then under $E^{\mathrm{all}}(v)$, it has at most $2p_{a}\deg_{H^{\prime}}(u)=\deg_{H^{\prime}}(u)/2$ active neighbors. Since $H^{\prime}$ is $\Pi$-ous, $u$ has at least as many available colors as uncolored neighbors, and thus retains its color with probability at least $1/2$ by the union bound.
Let $Z=\sum_{i}Z_{i}$ be the number of colored active neighbors of $v$ after Algorithm 2. By Lemma B.3 with $\mu=\sum_{i}\mu_{i}\geqslant d/2$ for all $i\in[d]$ and $t=d/4$, we get that
${\mathbb{P}}[Z<d/4\nonscript\>|{}\nonscript\>\mathopen{}E^{\mathrm{all}}(v)]\leqslant\exp\left(-\frac{d}{32}\right)\leqslant\exp\left(-\Omega(\Delta^{1/10})\right)\ ,$
where the last inequality uses that $d\geqslant(p_{a}/2)\deg_{H^{\prime}}(v)$ under $E^{\mathrm{all}}(v)$ and that $\deg_{H^{\prime}}(v)\geqslant\Delta^{1/10}$. Note that when $Z\geqslant d/4$, the uncolored degree of $v$ after Algorithm 2 is $\deg_{H^{\prime}}(v)-Z\leqslant(1-p_{a}/8)\deg_{H^{\prime}}(v)$. So, overall, we have that
[…]
Lemma 5.5.
Let $H^{\prime}$ be an induced uncolored subgraph of a $\Pi$-ous subgraph $H$ of $F-\bigcup_{i}A_{i}$. For every uncolored clique $A_{i}$, after Algorithm 2, $E_{CC}(i)$ holds w.p. at most $\Delta\cdot\exp(-\Delta^{1/40})$.
Proof.
Consider any uncolored clique $A_{i}$ and any color $x$. The relevant collection of subsets is formed by the external neighborhoods of each vertex $v\in A_{i}$. By Lemma 4.1-(a), we have $|A_{i}|<\Delta$. We ignore vertices of $\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ for their contribution is not accounted for in CC. Hence, no relevant vertex of $H$ belongs to more than $2\Delta^{9/10}$ such subsets. Since $H^{\prime}$ is $\Pi$-ous, each vertex $v$ of $H^{\prime}$ picks color $x$ as candidate color w.p. at most $1/(\deg_{H^{\prime}}(v)+U\cdot\Delta^{0.22}))\leqslant 1/(Q\cdot\Delta^{1/5})$ for $Q=U$. As candidate color choices are independent for different vertices, the probability that $\ell$ vertices pick color $x$ is at most $1/(Q\cdot\Delta^{1/5})^{\ell}$. Therefore, by Lemma 4.4, CC fails for clique $A_{i}$ and color $x$ w.p. at most $\exp(-\Delta^{1/40})$. The lemma follows by the union bound over a total of at most $\Delta$ different colors.
Given this, Algorithm 3 follows our shattering framework (see Section 3.2) to produce a coloring in which degree in $H$ have dropped.
1
Input : An uncolored $\Pi$-ous graph $H$
2
Output : A proper partial list-coloring of $H$
3
4Run Algorithm 2 on $H$ and let $\mathcal{A}^{\prime}$ be the set of events $E(v)$ or $E_{CC}(i)$ that occur.
5
6Retract the colors of (i.e., uncolor) the vertices in $\operatorname{vbl}(\mathcal{A}^{\prime})$
7
8Solve the following LLL using the deterministic algorithm from Theorem 3:
   Random process: Run Algorithm 2 on $H^{\prime}$ where $H^{\prime}$ is the induced subgraph of $H$ of uncolored vertices $v$ with $\operatorname{dist}_{F}(v,\operatorname{vbl}(\mathcal{A}^{\prime}))\leqslant 2$
   Bad events:
  1. $E(v)$ for all vertices $v\in N(\operatorname{vbl}(\mathcal{A}^{\prime}))\cup\operatorname{vbl}(\mathcal{A}^{\prime})$ with $\deg_{H^{\prime}}(v)\geqslant\Delta^{1/10}$, and
  2. $E_{CC}(i)$ for all uncolored clique $A_{i}$ adjacent to a vertex in $H^{\prime}$.
Output the partial list-coloring $\varphi$ where $\varphi(v)$ is the color given in Algorithm 3 for vertices $v\notin H^{\prime}$, and $\varphi(v)$ is the color given in Algorithm 3 for the vertices of $v\in H^{\prime}$.
Algorithm 3 Single iteration of degree reduction in $H$
Remark 5.6.
As detailed in CC, there is a fresh budget of $CC(i,x)<\Delta^{\frac{37}{40}}$ for each color $x$, each clique $A_{i}$, and also for each coloring step of the algorithm. In particular, in Algorithms 3 and 3 of Algorithm 3 the events $E_{CC}(i)$ get a separate budget of $\Delta^{\frac{37}{40}}$ each.
Lemma 5.7.
Suppose $\Delta\leqslant(\log n)^{50}$. Let $G_{\operatorname{dep}}$ be the dependency graph of the LLL of Algorithm 3 in Algorithm 3. With high probability, its largest connected component has size at most $\operatorname{\text{{\rm poly}}}(\log n)$.
Proof.
In order to bound the size of the largest connected component of $G_{\operatorname{dep}}$ we use Lemma 3.2; next we argue that the lemma applies. We take the whole graph $F$ as graph $G$ from Lemma 3.2, because every vertex $v$ of $H$ has one variable encoding its color choice. The LLLs (formally sets of sets) $\mathcal{A}$ and $\mathcal{B}$ required to use the lemma are both given by all collections of $\operatorname{vbl}(E(v))$ and $\operatorname{vbl}(E_{CC}(i))$ for all $v\in V(H)$ and uncolored cliques $A_{i}\in A_{L}\cup A_{H}$. The set $\mathcal{A}^{\prime}$ consists of the events $E(v)$ and $E_{CC}(i)$ that hold after Algorithm 3 of Algorithm 3. Let $\mathcal{B}^{\prime}$ with $c_{\mathcal{B}^{\prime}}=3$ be as in the statement of Lemma 3.2. Let $c_{1}=5,c_{2}=2$, and $c_{4}=3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+2$. We show that the preconditions Steps (S1), (S2), (S4) and (S3) of Lemma 3.2 hold.
Step (S1) holds with $c_{1}=5$ due to 5.3 (and because $F$ includes the cliques): a vertex $v$ holds variables used only by events $E(u)$ with $u\in N_{H}^{\leqslant 2}(v)$ (which has diameter at most 4) or events $E_{CC}(i)$ with $A_{i}$ with $\operatorname{dist}_{G}(v,A_{i})\leqslant 2$ (which has diameter at most 5 in $F$). Step (S2) requires a bound $\Delta(F)^{c_{2}}$ for some constant $c_{2}$ on the number of events in which each variable appears and it holds with $c_{2}=2$ as each variable is only used by events with variables in its $2$-hop neighborhood. Step (S3) requires that for each event $A\in\mathcal{A}$ we can determine $1(A\in\mathcal{A}^{\prime})$ by only evaluating the variables in $\operatorname{vbl}(A)$ which immediately holds by the definition of $\mathcal{A}^{\prime}$. Step (S4) requires an upper bound of $1/\Delta(F)^{c_{4}}$ on the probability for each bad event $A\in\mathcal{A}$ to be contained in $\mathcal{A}^{\prime}$ where $c_{1}$ is an arbitrary constant satisfying $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1$. By Lemmas 5.4 and 5.5 this probability is at most $\Delta(G)\cdot\exp(-\Delta(G)^{1/40})$. As $\Delta(G)=\Theta(\Delta(F))$ Step (S4) holds for $c_{4}$ (as chosen above) as long as $\Delta(G)$ is larger than a sufficiently large absolute constant $\Delta_{0}$. The set of events included in the LLL of Algorithm 3 are all included in $\mathcal{B}^{\prime}$ (with $c_{\mathcal{B}^{\prime}}$ defined above) as vertices with distance at most $2$ to $\operatorname{vbl}(\mathcal{A}^{\prime})$ participate in the post-shattering random process and any event $E(v)$ or $E_{CC}(i)$ adjacent to such a vertex participates in the post-shattering as well; note that the event $E_{CC}(i)$ for a clique $A_{i}$ that has the closest vertex $w$ in $H^{\prime}$ in distance $2$ does not need to participate despite $w\in\operatorname{vbl}(E_{CC}(i))$ as the event is avoided regardless the color choice of $w$. Hence by Lemma 3.2, the connected components of $G_{\operatorname{dep}}$ have size at most $O(\Delta(F)^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}+c_{2}}\log n)=\operatorname{\text{{\rm poly}}}(\log n)$ with high probability.
Remark 5.8.
Determining and verifying the minimal admissible value of the constant $c_{\mathcal{B}^{\prime}}$ in the above proof is somewhat tedious. The preconditions required for the shattering argument in Lemma 3.2 remain valid even for larger choices of $c_{\mathcal{B}^{\prime}}$, since the probability that any given event is contained in $\mathcal{A}^{\prime}$ decreases exponentially in $\Delta$. Consequently, Step (S4) holds for any constant $c_{4}$ satisfying $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1$, provided that $\Delta$ is larger than a sufficiently large constant $\Delta_{0}=\Delta_{0}(c_{\mathcal{B}^{\prime}})$.
In the subsequent sections, we use several analogous applications of Lemma 3.2. For the sake of clarity and ease of verification, we do not always optimize the choice of $c_{\mathcal{B}^{\prime}}$, but instead adopt values that simplify the argument.
Finally, we note that our retraction algorithm is deliberately somewhat aggressive. For instance, if $E_{CC}(i)$ holds for a clique $A_{i}$, we retract all color choices of vertices in $\operatorname{vbl}(E_{CC}(i))$, including those at distance two from $A_{i}$, even though it would suffice to retract only the color choices of vertices in $N(A_{i})$. This uniform retraction rule helps streamline what is otherwise an already intricate procedure.
Proof of Lemma 5.1.
If $\Delta\geqslant(\log n)^{50}$, then by Lemmas 5.4 and 5.5, none of the bad events occur at Algorithm 3 of Algorithm 3. Hence, (1) and (2) of Lemma 5.1 already hold after Algorithm 3 (i.e., $\mathcal{A}^{\prime}=\emptyset$) and we skip the post-shattering phase. We henceforth assume that $\Delta\leqslant(\log n)^{50}$.
For a vertex $v\in N_{H}(\operatorname{vbl}(\mathcal{A}^{\prime}))$, note that each of its neighbors belong to $N_{H}^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))$, hence every uncolored neighbor of $v$ in $H^{\prime}$ samples a random color in the run of Algorithm 2 in Algorithm 3. By Lemmas 5.4 and 5.5, after coloring some of the vertices of $H$ during Algorithm 3, the set of events described in Algorithm 3 of Algorithm 3 indeed forms an LLL whose dependency degree is bounded by $\operatorname{\text{{\rm poly}}}\Delta$ via 5.3, and thus an assignment of colors for vertices of $\operatorname{vbl}(\mathcal{A}^{\prime})$ that avoids its bad events can be found by Theorem 3. Running Algorithm 2 takes $O(1)$ rounds, and by 5.3 so does Algorithm 3. By Lemma 5.7, the connected components of the dependency graph of the LLL solved in Algorithm 3 have size at most $N=\operatorname{\text{{\rm poly}}}(\log n)$. Hence, Theorem 3 runs in $\widetilde{O}(\log^{3}N)=\widetilde{O}(\log^{3}\log n)$.
Let us now conclude by verifying that both properties claimed by Lemma 5.1 hold after nodes adopt colors as specified in the last line of Algorithm 3. If $v$ has fewer than $\Delta^{1/10}$ uncolored neighbors, then (1) vacuously holds, so let us assume the contrary. Consider a vertex $v\notin N(\operatorname{vbl}(\mathcal{A}^{\prime}))\cup\operatorname{vbl}(\mathcal{A}^{\prime})$. Then after Algorithm 3, the degree of $v$ has decreased by a factor of (11/180)(1-1/180) or dropped below $\Delta^{1/10}$ (otherwise $E(v)$ holds and $v\in N(\operatorname{vbl}(\mathcal{A}^{\prime}))\cup\operatorname{vbl}(\mathcal{A}^{\prime})$) and none of the colors in $N(v)$ are retracted (otherwise $v$ has a neighbor in $\operatorname{vbl}(\mathcal{A}^{\prime})$). Consider now a vertex $v\in N(\operatorname{vbl}(\mathcal{A}^{\prime}))\cup\operatorname{vbl}(\mathcal{A}^{\prime})$. The post-shattering LLL (described Algorithm 3) includes the event $E(v)$, so the uncolored degree of $v$ in $H^{\prime}$ drops by a (11/180)(1-1/180) factor, hence it at most $(1-1/180)\deg_{H}(v)$. Since $H^{\prime}$ contains all the uncolored neighbors of $v$, the uncolored degree of $v$ at the end of the algorithm is upper bounded by the uncolored degree of $v$ in $H^{\prime}$.
After retracting colors in $\operatorname{vbl}(\mathcal{A}^{\prime})$, none of the $E_{CC}(i)$ holds anymore and thus CC is maintained by the pre-shattering phase. During post-shattering, only vertices of $N_{H}^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))$ get colored. Hence, CC continues to hold for the $A_{i}$ that are not adjacent to some vertex in $N_{H}^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))$. If $A_{i}$ is adjacent to such a vertex, then CC is maintained by the post-shattering coloring (with a fresh budget) because $E_{CC}(i)$ would otherwise hold.
5.2 Finishing Off The Coloring via MCT (Lemma 5.2)
In this section, we analyze the subsequent coloring step after computing the partial coloring in Section 5.1. This step of computing the complete coloring of the considered vertices involves a single-round Mct (equivalent to Lemma 35 of [45]). In short, each vertex picks a set of colors and compares them to the sets of its neighbors. If there is a color in this set that was picked by none of the neighbors, then the vertex retains that color. See Algorithm 4.
Input : An uncolored graph $H^{\prime}$ and a parameter $T$
Output : A partial coloring of $H^{\prime}$
1 Every vertex $v\in V(H^{\prime})$ samples a set $S(v)$ of $T$ uniform available colors with repetitions
2 if $\exists\chi\in S(v)\setminus S(N(v))$ then Color $v$ with $\chi$ else $v$ remains uncolored
Algorithm 4 MultiColorTrial
We introduce the following events for every uncolored vertex $v\in V(H)$ and every uncolored clique $A_{i}\in A_{H}\cup A_{L}$:
$E^{\prime}(v)$: $v$ does not get colored,
$E_{CC}(i)$: CC is not maintained for some color w.r.t. $A_{i}$.
Observation 5.9.
$\operatorname{vbl}(E^{\prime}(v))\subseteq N_{H}^{\leqslant 2}(v)$ and $\operatorname{vbl}(E_{CC}(i))\subseteq N_{F}^{\leqslant 2}(A_{i})\cap V(H)$.
Similarly to Section 5.1, we argue that under the conditions of Lemma 5.2, both events are unlikely to occur when we run Algorithm 4.
Lemma 5.10.
Let $H$ be an uncolored graph with maximum degree at most $\Delta^{1/10}$. For every $v\in V(H)$ with $\Delta^{9/20}$ slack in $H$, after Algorithm 4 with $T=\Delta^{1/10}$, the event $E^{\prime}(v)$ holds w.p. at most $\exp(-\Delta^{1/10})$.
Proof.
Since each vertex has at most $\Delta^{1/10}$ uncolored neighbors, there are also at most $T\Delta^{1/10}=\Delta^{1/5}$ different candidate colors appearing in its neighborhood. Fix $S(N(v))$ arbitrarily. Each of the $T$ colors that $v$ picks uniformly at random from its list of available colors belongs to this set w.p. at most $\Delta^{1/5}/|L(v)|$. Since $L(v)$ contains at least $\Delta^{9/20}$ colors, the probability that $v$ is uncolored after Algorithm 4 is at most $\left(\Delta^{1/5}/|L(v)|\right)^{T}\leqslant\Delta^{-T/4}\leqslant\exp(-\Delta^{1/10})$.
Lemma 5.11.
Let $H$ be an uncolored subgraph of $F-\bigcup_{i}A_{i}$ such that every uncolored $v\in A_{i}$ has at most $\Delta^{1/10}$ neighbors in $H$. For every uncolored clique $A_{i}$, after Algorithm 4, $E_{CC}(i)$ holds w.p. at most $\Delta\cdot\exp(-\Delta^{1/40})$.
Proof.
We again apply Lemma 4.4 to upper bound the probability. Fix a clique $A_{i}$ and a color and form the sets consisting of external neighbors not in $\operatorname{Big}_{i}^{+}$ as before. Set $Q=\Delta^{1/10}$. The probability that a specific color is picked by neighbor is bounded above by $T/|L(v)|=1/\Delta^{7/20}\leqslant 1/(Q\cdot\Delta^{1/5})$. Due to the independence of picking candidate colors between different vertices the probability that a specific color is picked by $\ell$ vertices is at most $1/(Q\cdot\Delta^{1/5})^{\ell}$. No vertex lies in more than $2\Delta^{9/10}$ sets, because in that case it would belong to some $\operatorname{Big}_{i}^{+}$. Now we can apply Lemma 4.4 and get that for a fixed clique and color the probability that CC fails is at most $\exp(-\Delta^{1/40})$. With a union bound over all colors, $E_{CC}(i)$ holds w.p. at most $\Delta\cdot\exp(-\Delta^{1/40})$.
Pre-shattering:
1 Run Algorithm 4 with $T=\Delta^{1/10}$.
2 Let $\mathcal{A}^{\prime}$ be the set of events $E^{\prime}(v)$ or $E_{CC}(i)$ that occur.
3 Retract the colors of (i.e., uncolor) vertices in $\operatorname{vbl}(\mathcal{A}^{\prime})$
4
Post-shattering:
5 Solve the following LLL using the deterministic algorithm from Theorem 3:
   Random process: Run Algorithm 4 with $T=\Delta^{1/10}$ on $\operatorname{vbl}(\mathcal{A}^{\prime})$
   Bad events:
  1. $E^{\prime}(v)$ for every uncolored vertex $v$, and
  2. $E_{CC}(i)$ for each uncolored clique $A_{i}$ adjacent to a vertex in $\operatorname{vbl}(\mathcal{A}^{\prime})$
Algorithm 5 Mct Coloring
See 5.2
Proof.
If $\Delta\geqslant(\log n)^{50}$, then by Lemmas 5.10 and 5.11, none of the bad events occur after running Algorithm 4 in pre-shattering, w.h.p.. Hence, $H$ is colored and CC is maintained w.h.p.. We henceforth assume that $\Delta\leqslant(\log n)^{50}$.
Let $G_{\operatorname{dep}}$ be the dependency graph of the $E^{\prime}(v)$ and $E_{CC}(i)$ and $B$ be the set of events of the post-shattering LLL in Algorithm 5 of Algorithm 5. We first reason how to apply Lemma 3.2 to bound the size of each connected component of $G_{\operatorname{dep}}$. We take the whole graph $F\supseteq H$ as graph $G$ from Lemma 3.2, because every vertex of $H$ has one variable in Algorithm 3. The LLLs (formally sets of sets) $\mathcal{A}$ and $\mathcal{B}$ required to use the lemma are both given by all collections of $\operatorname{vbl}(E^{\prime}(v))$ and $\operatorname{vbl}(E_{CC}(i))$ for all $v\in V(H)$ and uncolored cliques $A_{i}\in A_{L}\cup A_{H}$. The set $\mathcal{A}^{\prime}$ consists of the events $E^{\prime}(v)$ and $E_{CC}(i)$ that hold after Algorithm 5 of Algorithm 5. Let $\mathcal{B}^{\prime}$ with $c_{\mathcal{B}^{\prime}}=0$ be as in the statement of Lemma 3.2. Let $c_{1}=5,c_{2}=2$, and $c_{4}=3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+2$. We show that the preconditions Steps (S1), (S2), (S4) and (S3) of Lemma 3.2 hold.
Step (S1) holds via 5.9 and we also obtain that each variable is used by at most $\Delta(F)^{c_{2}}$ events yielding Step (S2). Step (S3) follows by the definition of $\mathcal{A}^{\prime}$. In order to prove Step (S4), observe that by Lemmas 5.10 and 5.11 the probability of each $A\in\mathcal{A}$ to be contained in $\mathcal{A}^{\prime}$, i.e., $E^{\prime}(v)$ or $E_{CC}(i)$ hold, is upper bounded by $\exp(-\Omega(\Delta(G)^{1/10}))$ and $\Delta(G)\cdot\exp(-\Delta(G)^{1/40})$ respectively. Both of these values decrease exponentially in $\Delta(G)=\Theta(\Delta(F))$ and hence for $\Delta(G)\geqslant\Delta_{0}$ for an absolute constant $\Delta_{0}$ can be upper bounded by $1/\Delta(F)^{c_{4}}$ where $c_{4}$ is stated above. The set of events included in the LLL of Algorithm 5 of Algorithm 5 are all included in $\mathcal{B}^{\prime}$ (with the value of $c_{\mathcal{B}^{\prime}}$ defined above). Hence by Lemma 3.2, the connected components of $G_{\operatorname{dep}}$ have size at most $N=O(\Delta(F)^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}+c_{2}}\log n)=\operatorname{\text{{\rm poly}}}(\log n)$ with high probability.
Since the pre-shattering only colors vertices and that the slack does not decrease, Lemma 5.10 continues to apply to bound the probability of $E^{\prime}(v)$ in the post-shattering LLL. Likewise, Lemma 5.11 applies to bound the probability for $E_{CC}(i)$ in the post-shattering LLL. 5.9 bounds the dependency degree and hence, for sufficiently large $\Delta$, Algorithm 5 poses an LLL that we can solve with Theorem 3 in $\widetilde{O}(\log^{3}N)=\widetilde{O}(\log^{3}\log n)$ rounds; all other parts of the algorithm are executed in $O(1)$ rounds. At the end, all nodes of $H$ are colored as all events $E^{\prime}(v)$ are either avoided already after the pre-shattering phase or after the post-shattering phase; note that any uncolored node after the pre-shattering phase is contained in $\operatorname{vbl}(\mathcal{A}^{\prime})$. After retracting colors in $\operatorname{vbl}(\mathcal{A}^{\prime})$, none of the $E_{CC}(i)$ holds anymore and thus also CC is maintained by the pre-shattering phase. During post-shattering, only vertices of $\operatorname{vbl}(\mathcal{A}^{\prime})$ get colored. Hence, CC continues to hold for the $A_{i}$ that are not adjacent to some vertex in $\operatorname{vbl}(\mathcal{A}^{\prime})$. If $A_{i}$ is adjacent to such a vertex, then CC is maintained by the post-shattering coloring (with a fresh budget) as it is implied by avoiding $E_{CC}(i)$.
6 Coloring the Sparse Nodes
In this section, we show how to color the sparse nodes $S$ while maintaining $CC$. Recall that this is the first step of Algorithm 1.
See 4.7
6.1 Algorithm for Sparse Nodes
The first step is similar to [45]: to generate slack for the nodes of $S$ in the form of sufficient color reuse. Initially, every vertex in $S$ tries a random color, resulting in about $\sqrt{\Delta}$ colors repeated in each neighborhood. A color is repeated in $N(v)$ if at least two vertices have it. When $\Delta\geqslant(\log n)^{50}$, every vertex gains $(1+\Omega(1))\sqrt{\Delta}$ slack w.h.p.. When $\Delta\leqslant\log^{50}n$, our algorithm differs significantly from [45], with an involved shattering argument; see Lemma 6.1 and Section 6.3 for details. We henceforth call $S^{\prime}\subseteq S$ the set of vertices uncolored after slack generation.
If the remaining subgraph $F[S^{\prime}]$ satisfied property $\Pi$, we could complete the task with algorithm ColorWithMuchSlack. However, though the slack obtained is significant, it can still be small relative to the external degree of cliques (see Lemma 4.1(3)).
The solution is to color the nodes of $F[S^{\prime}]$ in two batches. We split $S^{\prime}$ into subgraphs $S_{1}$ and $S_{2}$ via a degree-splitting LLL that guarantees that degrees into $S_{1}$ and $S_{2}$ are similar to what a random split would provide. In Lemma 6.4, we show $F[S_{1}]$ is $\Pi$-ous (thanks to the uncolored neighbors in $S_{2}$), and that $F[S_{2}]$ is $\Pi$-ous regardless of the coloring in $S_{1}$ (as long as it extends the coloring produced by slack generation). We can then color $S^{\prime}$ by using ColorWithMuchSlack twice: first to color $S_{1}$ and then $S_{2}$.
1
2SlackGeneration
3
4Partition $S^{\prime}$ into sets $S_{1}$ and $S_{2}$ using the degree splitting algorithm of Lemma 6.3 when $\Delta\leqslant(\log n)^{50}$ and by sampling every vertex of $S^{\prime}$ into $S_{2}$ w.p. $p=2\Delta^{-1/4}$ otherwise
5
6ColorWithMuchSlack($S_{1}$)
ColorWithMuchSlack($S_{2}$)
Algorithm 6 ColorSparse
We begin by stating the properties of the SlackGeneration step. We defer the proof of Lemma 6.1 to Section 6.3 to preserve the flow of the paper.
Lemma 6.1.
There is a randomized $\mathsf{LOCAL}$ algorithm that computes in $\widetilde{O}(\log^{3}\log n)$ rounds a partial coloring of the subgraph $F$ induced by $S$ such that:
  1. (a)
    For every $v\in S$, either $v$ has fewer than $\Delta-3\sqrt{\Delta}$ neighbors in $S$ or at least $1.05\sqrt{\Delta}$ colors appear at least twice in $N(v)\cap S$;
  2. (b)
    CC is maintained; and
  3. (c)
    Every vertex in $S$ has at most $19\Delta/20$ colored neighbors.
If $\Delta\geqslant(\log n)^{50}$, then the algorithm ends after $O(1)$ round.
For the rest of this section, $L(v)$ denotes the set of colors available after slack generation (Algorithm 6 in Algorithm 6). The main implication of Lemma 6.1 is the following lower bound on the number of available colors. We emphasize that for 6.2, it is crucial that the number of repeated colors guaranteed by Lemma 6.1(a) is $(1+\Omega(1))\sqrt{\Delta}$.
Observation 6.2.
All $v\in S^{\prime}$ have $|L(v)|\geqslant\deg_{S^{\prime}}(v)+0.05\sqrt{\Delta}$.
Proof.
There are $c\geqslant\Delta-k_{\Delta}+1\geqslant\Delta-\sqrt{\Delta}+1$ colors. If $\deg_{S}(v)<\Delta-3\sqrt{\Delta}$, then $c\geqslant\deg_{S}(v)+2\sqrt{\Delta}$ and the claim follows as each color lost in $L(v)$ corresponds to a colored neighbor (thus not in $S^{\prime}$). If $\deg_{S}(v)\geqslant\Delta-3\sqrt{\Delta}$, then by Lemma 6.1(a), it has at least $1.05\sqrt{\Delta}$ repeated colors in $N(v)$. Simple accounting shows that $L(v)$ contains at least
$c-\deg_{S\setminus S^{\prime}}(v)+1.05\sqrt{\Delta}\geqslant\Delta-\deg_{S\setminus S^{\prime}}(v)+0.05\sqrt{\Delta}\geqslant\deg_{S^{\prime}}(v)+0.05\sqrt{\Delta}\ ,$
where the first inequality uses the definition of $c$ and the latter uses that $F[S]$ has maximum degree $\Delta$.
The following degree splitting result is used as a subroutine to implement line 3 of Algorithm 6. Similar results were given in [36], but not the exact claim we need.
Lemma 6.3.
There is a universal constant $\alpha>0$ for which the following holds. Let $p\in(0,1)$, let $H$ be a graph with maximum degree $\Delta\leqslant(\log n)^{50}$ and $S^{\prime}\subseteq V(H)$. There is a $\widetilde{O}(\log^{3}\log n)$-round algorithm that partitions $S^{\prime}$ into $S_{1}$ and $S_{2}$ such that, w.h.p., every vertex $v\in V(H)$ with $\deg_{S^{\prime}}(v)\geqslant\alpha p^{-1}\log\Delta$ has
  1. 1.
    at most $4p\deg_{S^{\prime}}(v)$ neighbors in $S_{2}$, and
  2. 2.
    at least $p\deg_{S^{\prime}}(v)/2$ neighbors in $S_{2}$.
We emphasize that the guarantee of Lemma 6.3 applies to all vertices of $F$ with sufficiently many neighbors in $S^{\prime}$. In particular, it splits the degree of vertices in $A_{i}$ towards $S^{\prime}$ as well. The proof is deferred to Section 6.2 and follows the shattering framework.
Lemma 6.4.
There is a $\widetilde{O}(\log^{3}\log n)$ round $\mathsf{LOCAL}$ algorithm that partitions $S^{\prime}$ into sets $S_{1}$ and $S_{2}$ such that
  1. 1.
    $F[S_{1}]$ is $\Pi$-ous, and
  2. 2.
    $F[S_{2}]$ is $\Pi$-ous regardless of how the coloring is extended to the vertices of $S_{1}$.
When $\Delta\geqslant(\log n)^{50}$, the algorithm ends after $O(1)$ rounds.
Proof.
Let $p=2\Delta^{-1/4}$. When $\Delta\geqslant(\log n)^{50}$, we sample vertices of $S^{\prime}$ into $S_{2}$ with probability $p$, by the Chernoff Bound and Union Bound, the guarantees of Lemma 6.3 (1,2) hold with high probability. When $\Delta\leqslant(\log n)^{50}$, we split the vertices of $S^{\prime}$ into $S_{1}$ and $S_{2}$ using the algorithm of Lemma 6.3 with $H=F$, $S^{\prime}$ as itself, and $p=2\Delta^{-1/4}$. We henceforth focus on proving that (1) and (2) hold.
Let $U_{1}=10^{8}\cdot\sqrt{\Delta}$, and $U_{2}=\alpha\Delta^{1/4}\log\Delta$ where $\alpha$ is the constant from Lemma 6.3. We claim that $F[S_{1}]$ is $\Pi$-ous with $U=U_{1}$ and $F[S_{2}]$ is $\Pi$-ous with $U=U_{2}$, even after we extend the coloring to $S_{1}$.
Proof of 1. A node with $\deg_{S^{\prime}}(v)<\Delta/40$, has a list of size $\Omega(\Delta)$. Indeed, it has at most $19\Delta/20$ colored neighbors (by Lemma 6.1(c)) thus $|L(v)|\geqslant c-19\Delta/20\geqslant\Delta/20-k_{\Delta}\geqslant\deg_{S^{\prime}}(v)+\Delta/40-k_{\Delta}\geqslant\deg_{S^{\prime}}(v)+4\Delta^{3/4}\geqslant\deg_{S^{\prime}}(v)+4U_{1}\Delta^{0.22}$. A node with $\deg_{S^{\prime}}(v)\geqslant\Delta/40$ has at least $\deg_{S^{\prime}}(v)/\Delta^{1/4}\geqslant\Delta^{3/4}/40$ neighbors in $S_{2}$ by Lemma 6.3(2). Thus, we obtain by 6.2 that
$|L(v)|\geqslant\deg_{S^{\prime}}(v)=\deg_{S_{1}}(v)+\deg_{S_{2}}(v)\geqslant\deg_{S_{1}}(v)+\Delta^{3/4}/40.$
For sufficiently large $\Delta$, it holds that $\Delta^{3/4}/40>U_{1}\cdot\Delta^{0.22}$, thereby proving Part (b) of $\Pi$. As for Part (a), recall that a vertex in $A$ has at most $U_{1}=10^{8}\sqrt{\Delta}$ neighbors in $S$ (by Lemma 4.1(3)).
Proof of 2. Consider any extension of the coloring to $S_{1}$ and let $L_{2}(v)$ denote the list of colors still available to $v\in S_{2}$. By 6.2, after coloring the nodes in $S_{1}$ in an arbitrary manner, the list size of a node $v\in V(S_{2})$ is at least
$|L_{2}(v)|\geqslant|L(v)|-\deg_{S_{1}}(v)\geqslant\deg_{S_{2}}(v)+0.05\sqrt{\Delta}\geqslant\deg_{S_{2}}(v)+U_{2}\cdot\Delta^{0.22}\penalty 10000\ ,$
where the last inequality holds as for sufficiently large $\Delta$. Part (a) of $\Pi$ trivially holds for a vertex $u\in\bigcup_{i}A_{i}$ with $\deg_{S^{\prime}}(u)<\alpha\Delta^{1/4}\log\Delta=U_{2}$. Otherwise, the splitting lemma applies to $u$ and since $\deg_{S^{\prime}}(u)\leqslant 10^{8}\sqrt{\Delta}$, it has at most $4p\deg_{S^{\prime}}(u)\leqslant 8\deg_{S^{\prime}}(u)/\Delta^{1/4}\leqslant 8\cdot 10^{8}\Delta^{1/4}<U_{2}$ neighbors in $S_{2}$.
We conclude by putting all the statements together to prove the main result about coloring $S$.
Proof of Lemma 4.7.
By Lemma 6.4, we can use Lemma 4.5 to color $F[S_{1}]$ and $F[S_{2}]$. It takes $\widetilde{O}(\log\Delta\cdot\log^{3}\log n)$ rounds when $\Delta\leqslant(\log n)^{50}$. When $\Delta\geqslant(\log n)^{50}$, we use Lemma 4.6 instead to color in $O(\log^{*}n)$ rounds.
6.2 Degree Splitting (Proof of Lemma 6.3)
We partition $S^{\prime}$ into $S_{1}$ and $S_{2}$ with Algorithm 7 and argue that the claims of Lemma 6.3 are verified with high probability. The proof follows the shattering framework: first, we analyze the probability of bad events occurring during pre-shattering, deduce that the connected components of the dependency graph are $\operatorname{\text{{\rm poly}}}(\log n)$-sized in post-shattering, and finally verify that the post-shattering instance is indeed an LLL.
Input : a graph $H$ with maximum degree $\Delta$, and a set $S^{\prime}\subseteq V(H)$
Output : a partition $S_{1},S_{2}$ of $S^{\prime}$
1
Pre-Shattering
2 Each vertex of $S^{\prime}$ joins $S_{2}$ independently w.p. $p$.
3 For each $v$ with $\deg_{S^{\prime}}(v)\geqslant\alpha p^{-1}\log\Delta$, consider the event
[…]
    $|\deg_{S_{2}}(v)-p\deg_{S^{\prime}}(v)|>(p/2)\deg_{S^{\prime}}(v)$
5
Retractions
6 Let $X$ be the set of vertices $v$ for which $E_{d}(v)$ hold
7 Remove from $S_{2}$ every vertex adjacent to $X$, i.e., $S_{2}\leftarrow S_{2}\setminus(X\cup N(X))$
8
Post-Shattering
9 We solve the following LLL using Theorem 3
Every vertex $v\in N^{\leqslant 4}(X)\cap(S^{\prime}\setminus S_{2})$ independently joins $S_{2}$ w.p. $p$
For each $v\in N^{\leqslant 5}(X)$ with $\deg_{S^{\prime}}(v)\geqslant\alpha p^{-1}\log\Delta$, we have the bad event
  1. $E_{d}^{\prime}(v)$:
    $\deg_{S_{2}}(v)>4p\deg_{S^{\prime}}(v)$ or $\deg_{S_{2}}(v)<(p/2)\deg_{S^{\prime}}(v)$.
Algorithm 7 Degree Splitting (Lemma 6.3)
The events $E_{d}(v)$ and $E^{\prime}_{d}(v)$ are determined by the random choices of their neighbors.
Observation 6.5.
$\operatorname{vbl}(E_{d}(v))\subseteq N_{H}^{\leqslant 1}(v)$ and $\operatorname{vbl}(E^{\prime}_{d}(v))\subseteq N_{H}^{\leqslant 1}(v)$.
The probability that a bad event $E_{d}(v)$ occurs during pre-shattering is low by a direct application of the Chernoff Bound.
Claim 6.6.
For all $v$, the event $E_{d}(v)$ occurs w.p. at most $2\exp(-p\deg_{S^{\prime}}(v)/12)$.
From this, we argue that the connected components of the post-shattering dependency graph are small.
Claim 6.7.
Suppose $\Delta\leqslant(\log n)^{50}$. The connected components of the dependency graph $G_{\operatorname{dep}}$ of Algorithm 7 of Algorithm 7 have size at most $\operatorname{\text{{\rm poly}}}(\log n)$ with high probability.
Proof.
In order to bound the size of the largest connected component of $G_{\operatorname{dep}}$ we use Lemma 3.2; next we argue that the lemma applies. We take the whole graph $F$ as graph $G$ from Lemma 3.2, because every vertex $v$ of $S^{\prime}\subseteq V(F)$ has one variable encoding whether it is contained in $S_{2}$ or not. The LLLs (formally sets of sets) $\mathcal{A}$ and $\mathcal{B}$ required to use the lemma are both given by all collections of $\operatorname{vbl}(E_{d}(v))$ for all $v\in V(H)$ satisfying $\deg_{S^{\prime}}(v)\geqslant\alpha p^{-1}\log\Delta$ and $\operatorname{vbl}(E^{\prime}_{d}(v))$ for all $v\in V(H)$ satisfying $\deg_{S^{\prime}}(v)\geqslant\alpha p^{-1}\log\Delta$ respectively. The set $\mathcal{A}^{\prime}$ consists of the events $E_{d}(v)$ that hold after Algorithm 7 of Algorithm 7. Let $\mathcal{B}^{\prime}$ with $c_{\mathcal{B}^{\prime}}=5$ be as in the statement of Lemma 3.2. Let $c_{1}=2$ and $c_{2}=1$. We show that the preconditions Steps (S1), (S2), (S4) and (S3) of Lemma 3.2 hold.
Step (S1) holds due to 6.5 with $c_{1}=2$. Step (S2) requires a bound $\Delta(F)^{c_{2}}$ for some constant $c_{2}$ on the number of events in which each variable appears. It holds with $c_{2}=1$ as a vertex $v$ holds variables used only by events $E_{d}(u)$ or $E^{\prime}_{d}(u)$ with $u\in N_{H}^{\leqslant 1}(v)\subseteq N_{F}^{\leqslant 1}(v)$. Step (S3) requires that for each event $A\in\mathcal{A}$ we can determine $1(A\in\mathcal{A}^{\prime})$ by only evaluating the variables in $\operatorname{vbl}(A)$ which immediately holds by the definition of $\mathcal{A}^{\prime}$. Step (S4) requires an upper bound of $1/\Delta(F)^{c_{4}}$ on the probability for each bad event $A\in\mathcal{A}$ to be contained in $\mathcal{A}^{\prime}$ where $c_{1}$ is an arbitrary constant satisfying $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1$. By claim 6.6 this probability is exponentially small in $\Delta(G)=\Theta(\Delta(F))$ and hence Step (S4) holds for $c_{4}$ (as chosen above) as long as $\Delta(G)$ is larger than a sufficiently large absolute constant $\Delta_{0}$. The set of events included in the LLL of Algorithm 3 are all included in $\mathcal{B}^{\prime}$. Hence by Lemma 3.2, the connected components of $G_{\operatorname{dep}}$ have size at most $O(\Delta(F)^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}+c_{2}}\log n)=\operatorname{\text{{\rm poly}}}(\log n)$ with high probability.
And we now prove that the post-shattering events indeed form an LLL.
Claim 6.8.
In post-shattering, the probability that $E_{d}^{\prime}(v)$ occurs is at most $2\Delta^{-\alpha/24}$, and dependency degree of the $\{E_{d}^{\prime}(v):v\in N^{\leqslant 5}(X)\}$ is at most $\Delta^{2}$.
Proof.
Every vertex $v\in N^{\leqslant 5}(X)$ has at most $\deg_{S^{\prime}}(v)$ neighbors taking part in the post-shattering sampling. In expectation at most $p\deg_{S^{\prime}}(v)$ of them join $S_{2}$ and, by the Chernoff Bound, the probability to have more than $2p\deg_{S^{\prime}}(v)$ join $S_{2}$ is at most $\exp(-p\deg_{S^{\prime}}(v)/12)$. If $v\in X$, then none of its neighbors remain in $S_{2}$ after retractions. Otherwise, $E(v)$ did not hold, and hence $v$ has at most $2p\deg_{S^{\prime}}(v)$ neighbors in $S_{2}$ before the post-shattering sampling. Thus, the probability that a vertex has more than $4p\deg_{S^{\prime}}(v)$ neighbors in $S_{2}$ is upperbounded by $\exp(-p\deg_{S^{\prime}}(v)/12)\leqslant\Delta^{-\alpha/12}$.
A vertex $v\notin N^{\leqslant 3}(X)$ does not see any retraction in its neighborhood and has enough neighbors in $S_{2}$ already (it would otherwise belong to $X$). Let $v\in N^{\leqslant 3}(X)$. Observe that $N(v)\cap S^{\prime}\subseteq N^{\leqslant 4}(X)\cap S^{\prime}$, hence all its neighbors in $S^{\prime}\setminus S_{2}$ participate in the post-shattering sampling. Call $d$ its number of neighbors in $S_{2}$ before the post-shattering (and after retractions). It has at $p(\deg_{S^{\prime}}(v)-d)$ neighbors that join $S_{2}$ in the post-shattering, hence by the Chernoff Bound, it gains fewer than $p(\deg_{S^{\prime}}(v)-d)/2$ neighbors in $S_{2}$ during the post-shattering with probability at most $\exp(-p(\deg_{S^{\prime}}(v)-d)/12)$. Hence, the degree in $S_{2}$ is smaller than $d+p(\deg_{S^{\prime}}(v)-d)/2=p\deg_{S^{\prime}}(v)/2+(1-p/2)d$ w.p. at most $\exp(-p(\deg_{S^{\prime}}(v)-d)/12)$. If $d\geqslant p\deg_{S^{\prime}}(v)/2$, then we already have that $v$ has enough neighbors in $S_{2}$. Otherwise, after the post-shattering sampling it has fewer than $p\deg_{S^{\prime}}(v)/2$ w.p. at most $\exp(-p\deg_{S^{\prime}}(v)/24)\leqslant\Delta^{-\alpha/24}$.
By adding up the error probability of both ways, the event $E_{d}^{\prime}(v)$ holds w.p. at most $2\Delta^{-\alpha/24}$. Whether $E^{\prime}(v)$ is a function of the samplings of all $u\in N(v)$. Hence $E_{d}^{\prime}(v)$ is independent from every $E_{d}(u)$ where $u\notin N^{\leqslant 2}(v)$. So the dependency degree is bounded above by $\Delta^{2}$.
Now we can put all of this together to prove our degree splitting lemma.
See 6.3
Proof of Lemma 6.3.
Pre-shattering and retractions require $O(1)$ rounds of $\mathsf{LOCAL}$. Claim 6.7 shows that each connected component in the post-shattering phase is w.h.p. of size at most $\operatorname{\text{{\rm poly}}}(\log n)$. Claim 6.8 shows that it is an LLL with a sufficiently good polynomial criterion to apply Theorem 3 when $\alpha$ and $\Delta$ are large enough. Thus, the post-shattering LLL is solved in $\widetilde{O}(\log^{3}\log n)$ rounds.
It remains to show that the solution meets the requirements of the lemma.
Sufficiently few neighbors in $S_{2}$ (Part 1). Let $v\in N^{\leqslant 5}(X)$. Since $E^{\prime}(v)$ does not hold after post-shattering, it has at most $4p\deg_{S^{\prime}}(v)$ neighbors in $S_{2}$. If $v\notin N^{\leqslant 5}(X)$, then none of its neighbors participate in the sampling during post-shattering (recall that only nodes of $N^{\leqslant 4}(X)$ participate). Since $E(v)$ did not hold after pre-shattering, it has at most $2p\deg_{S^{\prime}}(v)$ neighbors in $S_{2}$ after post-shattering as well.
Enough neighbors in $S_{2}$ (Part 2): Let $v\in N^{\leqslant 5}(X)$. Since $E^{\prime}(v)$ does not hold after post-shattering, it has at least $p\deg_{S^{\prime}}(v)/2$ neighbors in $S_{2}$. Let $v\notin N^{\leqslant 5}(X)$. Since $v\notin X$, after the pre-shattering sampling, $v$ has at least $p\deg_{S^{\prime}}(v)/2$ neighbors in $S_{2}$. Retractions only occur in $X\cup N(X)$, hence $v$ is not affected. The number of neighbors in $S_{2}$ can only increase during post-shattering, hence it has enough neighbors in $S_{2}$ at the end of the algorithm.
6.3 Slack Generation (Proof of Lemma 6.1)
This section describes the algorithm that is executed in Algorithm 6 of Algorithm 6. The guarantees it provides are described by Lemma 6.1.
At the heart of Lemma 6.1 is the following key observation about random color trials, formally stated in Proposition 6.9: the coloring is such that $\Omega(\overline{m}/\Delta)$ colors are repeated in the neighborhood of a vertex with $\overline{m}$ missing edges. Crucially, as long as $\overline{m}\geqslant c\Delta\log\Delta$, this fails with probability at most $\Delta^{-\Theta(c)}$, giving rise to an LLL. This was already used by [45] (see Lemma 41). We say that a color is repeated if it is used at least two times.
We use the following variant of the result in which a set $A$ of active nodes try colors from a custom set of $q$ colors, which we prove in Appendix C.
Proposition 6.9.
There exists a universal constant $c_{0}$ for which the following holds. Let $\Delta$ and $q$ be positive integers with $q\geqslant\Delta/3$. Let $A\subseteq V(G)$ be a subset of nodes such that $F[A]$ has maximum degree at most $\Delta$, and CC a set of $q$ colors. Consider a node $v\in V$ with at least $\overline{m}$ non-edges in $F[N(v)\cap A]$ and such that $\overline{m}\geqslant c_{0}\cdot q$. If every vertex of $A$ samples a random color $\chi(v)\in C$ and retains it only if $\chi(v)\notin\chi(N(v)\cap A)$, then
node $v$ has at least $\frac{\overline{m}/q}{3\cdot 10^{4}}$ repeated colored in $N(v)$
with probability at least $1-\exp(-\Omega(\overline{m}/q))$.
To implement Lemma 6.1 when $\Delta$ is small, we resort to the shattering technique. To ensure that Proposition 6.9 applies after shattering, we use two disjoint sets of colors $C_{1}=\{1,2,\ldots,\Delta/2\}$ and $C_{2}=\{\Delta/2+1,\ldots,c\}$ for pre-shattering and post-shattering respectively. In post-shattering, some vertices will only have part of their neighbors sample colors. We thus need to ensure that the post-shattering instance retains enough of the original sparsity. This can be achieved by activating nodes randomly as Proposition 6.10 describes.
Proposition 6.10.
Suppose $H$ is a graph of maximum degree $\Delta$ and let $v$ be a vertex such that $H[N(v)]$ contains $\overline{m}$ anti-edges. Sample every vertex into a set $A$ with probability $p\geqslant 8/\Delta$. The number of anti-edges induced by $H[N(v)\cap A]$ is at least $p^{2}\overline{m}/2$ w.p. at least $1-\exp(-\Omega(p\overline{m}/\Delta))$.
It follows immediately that Proposition 6.10 holds in expectation. One obtains concentration through, e.g., Janson’s inequality. The proof of Proposition 6.10 is deferred to Appendix C.
We can now describe the algorithm for Lemma 6.1 run on the set of vertices $S$. It begins by activating nodes by sampling them into a set $R\subseteq S$ w.p. 1/2 , and runs the random experiment of Proposition 6.9 on activated nodes only. When $\Delta\geqslant(\log n)^{50}$, the claims of Lemma 6.1 already hold with high probability. When $\Delta$ is small, the algorithm begins by identifying failures (see below), it then retracts some colors (Algorithm 8), and finally use a deterministic algorithm for fixing the remaining parts of the graph (Algorithm 8).
To identify the aforementioned failures we define the following events. Define the following event for each $v\in S$:
  1. $E_{1}(v)$:
    $v$ has more than $\max\{(11/20)\deg_{S}(v),\Delta/20\}$ neighbors in $R$,
For each $v\in S$ with $\deg_{S}(v)\geqslant\Delta-3\sqrt{\Delta}$ consider the following events
  1. $E_{2}(v)$:
    it has fewer than $3\sqrt{\Delta}$ repeated colors in $F[N(v)\cap R]$, or
  2. $E_{3}(v)$:
    it has fewer than $9\cdot 10^{5}\Delta^{3/2}/8$ anti-edges in $F[N(v)\cap(S\setminus R)]$.
For each clique $A_{i}\in A_{L}\cup A_{H}$ define the following event:
  1. $E_{CC}(i)$:
    CC is violated for clique $A_{i}$
Pre-Shattering:
1 Each vertex of $S$ joins $R$ w.p. $1/2$
2 Each vertex in $R$ tries a random color in $C_{1}=[\Delta/2]$
3
Retractions:
4 Let $\mathcal{A}^{\prime}$ be the set of events $E_{1}(v)$, $E_{2}(v)$, $E_{3}(v)$, or $E_{CC}(i)$ (defined above) that occur.
5Retract the colors of (i.e., uncolor) the vertices in $\operatorname{vbl}(\mathcal{A}^{\prime})$ and remove them from $R$
6
Post-Shattering:
7 Solve the following LLL using the deterministic algorithm of Theorem 3:
Every vertex in $\{v\mid\operatorname{dist}_{F}(v,\operatorname{vbl}(\mathcal{A}^{\prime}))\leqslant 2\}\cap(S\setminus R)$ tries a color in $C_{2}=\{\Delta/2+1,\ldots,\Delta-k\}$ with probability $3/4$.
For each $v\in N^{\leqslant 3}(\operatorname{vbl}(\mathcal{A}^{\prime}))$, we have the bad event
  1. $E_{1}^{\prime}(v)$:
    $v$ has more than $\max\{(4/5)\cdot\deg_{S\setminus R}(v),\Delta/20\}$ colored neighbors in $S\setminus R$.
For every $v\in N^{\leqslant 1}(\operatorname{vbl}(\mathcal{A}^{\prime}))$ with $\deg_{S}(v)\geqslant\Delta-3\sqrt{\Delta}$, we also have the following bad event
  1. $E^{\prime}_{2}(v)$:
    $v$ has fewer than $1.05\sqrt{\Delta}$ colors repeated in $F[N(v)\cap(S\setminus R)]$
For each $A_{i}\in A_{L}\cup A_{H}$ with $\operatorname{vbl}(E_{CC}(i))\cap N^{\leqslant 1}(\operatorname{vbl}(\mathcal{A^{\prime}}))\neq\emptyset$ define the following event:
  1. $E_{CC}(i)$:
    CC is violated for clique $A_{i}$.
Algorithm 8 Implementation of Slack Generation (Lemma 6.1)
Let us begin by proving that events from the pre-shattering are unlikely.
Claim 6.11.
For all $v$ in $S$, each event $E_{i}(v)$, $i\in[3]$ occurs w.p. at most $\exp(-\Omega(\Delta^{1/40}))$. For each $A_{i}\in A_{L}\cup A_{H}$ the event $E_{CC}(i)$ occurs with probability at most $\Delta\cdot\exp(-\Omega(\Delta^{1/40}))$.
Proof.
For the same reason as in Lemma 5.5, ${\mathbb{P}}[E_{CC}(i)]\leqslant\Delta\exp(-\Omega(\Delta^{1/40}))$. Also, since vertices try colors with probability $1/2$, the Chernoff Bound directly implies that ${\mathbb{P}}[E_{1}(v)]\leqslant\exp(-\Omega(\Delta))$.
Let us now consider some fixed $v$ with $\deg_{S}(v)\geqslant\Delta-3\sqrt{\Delta}$ and bound the probability of $E_{2}(v)$ and $E_{3}(v)$. By Lemma 4.1(1), the vertex $v\in S$ has $\overline{m}\geqslant 9\cdot 10^{5}\Delta^{3/2}$ anti-edges in its neighborhood. By Proposition 6.10 on the subgraph of $F$ induced by $S$ and $p=1/2$, there are fewer than $9\cdot 10^{5}\Delta^{3/2}/8$ of anti-edges in $F[N(v)\cap R]$ w.p. at most $\exp(-\Omega(\Delta^{1/2}))$. Suppose that $F[N(v)\cap R]$ has $9\cdot 10^{5}\cdot\Delta^{3/2}/8$ induced anti-edges. By Proposition 6.9 (with $C=C_{1}$, $q\geqslant\Delta/3$ and $\overline{m}=9\cdot 10^{5}\Delta^{3/2}/8$), after vertices of $R$ try a random color from $C_{1}$, the number of pairs of non-adjacent neighbors colored the same in $N(v)\cap R$ is smaller than $\frac{9\cdot 10^{5}\Delta^{3/2}/8}{3\cdot 10^{4}\cdot\Delta}\geqslant 3\sqrt{\Delta}$ w.p. at most $\exp(-\Omega(\Delta^{1/2}))$. Hence, the event $E_{2}(v)$ occurs w.p. at most $\exp(-\Omega(\Delta^{1/2}))$.
To bound the probability of $E_{3}(v)$, observe that each vertex of $S$ joins the complement of $R$ independently w.p. $1/2$. By Proposition 6.10 on $H=F[S]$ where nodes are sampled in $A=S\setminus R$ w.p. $p=1/2$, the graph $F[N(v)\cap(S\setminus R)]$ contains fewer than $9\cdot 10^{5}\cdot\Delta^{3/2}/8$ anti-edges w.p. at most $\exp(-\Omega(\sqrt{\Delta}))$.
Next, we argue about the size of the connected components in the post-shattering instance.
Claim 6.12.
Suppose $\Delta\leqslant(\log n)^{50}$. Let $G_{\operatorname{dep}}$ be the dependency graph of the LLL of Algorithm 8 in Algorithm 8. With high probability, its largest connected component has size at most $\operatorname{\text{{\rm poly}}}(\log n)$.
Proof.
In order to bound the size of the largest connected component of $G_{\operatorname{dep}}$ we use Lemma 3.2; next we argue that the lemma applies. We take the whole graph $F$ as graph $G$ from Lemma 3.2, because every vertex $v$ of $H$ has one variable encoding whether it is contained in $R$ and its color choice in Algorithm 8. The LLLs (formally sets of sets) $\mathcal{A}$ required to use the lemma are given by the collections of $\operatorname{vbl}(E_{1}(v)),\operatorname{vbl}(E_{2}(v)),\operatorname{vbl}(E_{3}(v))$ and $\operatorname{vbl}(E_{CC}(i))$ for the respective $v\in V(H)$ and uncolored cliques $A_{i}\in A_{L}\cup A_{H}$. The LLL for $\mathcal{B}$ is given by $\operatorname{vbl}(E^{\prime}_{1}(v)),\operatorname{vbl}(E^{\prime}_{2}(v))$ and $\operatorname{vbl}(E_{CC}(i))$ for the respective nodes $v\in V(H)$ and uncolored cliques in $A_{i}\in A_{L}\cup A_{H}$. The set $\mathcal{A}^{\prime}$ consists of the events $\operatorname{vbl}(E_{1}(v)),\operatorname{vbl}(E_{2}(v)),\operatorname{vbl}(E_{3}(v))$ and $\operatorname{vbl}(E_{CC}(i))$ that hold after Algorithm 8 of Algorithm 8. Let $\mathcal{B}^{\prime}$ with $c_{\mathcal{B}^{\prime}}=3$ be as in the statement of Lemma 3.2. Let $c_{1}=5,c_{2}=2$, and $c_{4}=3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+2$. We show that the preconditions Steps (S1), (S2), (S4) and (S3) of Lemma 3.2 hold.
Step (S1) holds with a similar reasoning as in 5.3 (and because $F$ includes the cliques) for $c_{1}=5$. Step (S2) holds for $c_{2}=2$ as a variable is only contained in events with variables in its $2$-hop neighborhood. Step (S3) follows directly with the definition of $\mathcal{A}^{\prime}$. Step (S4) requires an upper bound of $1/\Delta(F)^{c_{4}}$ on the probability for each bad event $A\in\mathcal{A}$ to be contained in $\mathcal{A}^{\prime}$ where $c_{1}$ is an arbitrary constant satisfying $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1$. By claim 6.11 this probability is at most $\Delta(G)\cdot\exp(-\Delta(G)^{1/40})$. As $\Delta(G)=\Theta(\Delta(F))$ Step (S4) holds for $c_{4}$ (as chosen above) as long as $\Delta(G)$ is larger than a sufficiently large absolute constant $\Delta_{0}$. The set of events included in the LLL of Algorithm 8 are all included in $\mathcal{B}^{\prime}$ with the $c_{\mathcal{B}^{\prime}}$ defined above; a smaller choice of $c_{\mathcal{B}^{\prime}}$ is possible, but it is more tedious to verify that all events are then included in $\mathcal{B}^{\prime}$. Hence by Lemma 3.2, the connected components of $G_{\operatorname{dep}}$ have size at most $O(\Delta(F)^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}+c_{2}}\log n)=\operatorname{\text{{\rm poly}}}(\log n)$ with high probability.
The technical heart of this proof is the following claim, which argues that Proposition 6.9 still applies in the post-shattering instance. This comes from the introduction of $E_{3}(v)$ to ensure that pre-shattering preserved enough sparsity for the post-shattering step.
Claim 6.13.
In the post-shattering phase, events $E^{\prime}_{1}(v)$ and $E^{\prime}_{2}(v)$ occur w.p. at most $\exp(-\Omega(\Delta^{1/2}))$.
Proof.
The bound on ${\mathbb{P}}[E^{\prime}_{1}(v)]$ follows directly with a Chernoff bound as each node only picks a color with probability $3/4$, that is, in expectation $3/4\deg_{S\setminus R}(v)$ neighbors participate in the coloring process of the post-shattering phase, and the constant factor deviation (lower bounded by $\Delta/20)$ forbidden by $E^{\prime}_{1}(v)$ is exponentially small in $\Omega(\Delta)$.
Let $v\in N^{\leqslant 1}(\operatorname{vbl}(\mathcal{A}^{\prime}))\cap S$ be a vertex with $\deg_{S}(v)\geqslant\Delta-3\sqrt{\Delta}$. We claim that, after retractions, vertex $v$ has at least $9\cdot 10^{5}\cdot\Delta^{3/2}/8$ anti-edges in $F[N(v)\cap N^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))\cap(S\setminus R)]$, where $\mathcal{A}^{\prime}$ is the set of the occurring events previously defined. If $E_{3}(v)$ occurred during pre-shattering, we have that $N(v)\cap R=\emptyset$ because its neighbors are removed from $R$ during the retraction step. Otherwise, then it had $9\cdot 10^{5}\cdot\Delta^{3/2}/8$ anti-edges in $F[N(v)\cap(S\setminus R)]$ before retractions because $E_{3}(v)$ does not occur. Removing vertices from $R$ can only increase the number of anti-edges in $F[N(v)\cap(S\setminus R)]$; hence, it also has $9\cdot 10^{5}\cdot\Delta^{3/2}/8$ anti-edges in $F[N(v)\cap N^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))\cap(S\setminus R)]$ after retractions as $N(v)\cap S\subseteq N^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))\cap S$.
Consider now some vertex $v\in N^{\leqslant 1}(\operatorname{vbl}(\mathcal{A}^{\prime}))\cap S$ and let us upper bound the probability of $E_{2}^{\prime}(v)$. As we just argued, every such vertex $v$ has $9\cdot 10^{5}\Delta^{3/2}/8$ anti-edges in $F[N(v)\cap(S\setminus R)]$ (and all vertices of $N(v)\cap(S\setminus R)$ are trying colors if activated). By Proposition 6.10 (with $p=3/4$), node $v$ has fewer than $(3/4)^{2}/2\cdot 9\cdot 10^{5}\cdot\Delta^{3/2}/8$ w.p. at most $\exp(-\Omega(\Delta^{1/2}))$. So, by Proposition 6.9, the node $v$ has fewer than
$\frac{(3/4)^{2}/2\cdot 9\cdot 10^{5}\Delta^{3/2}/8}{3\cdot 10^{4}\cdot\Delta}\geqslant 1.05\sqrt{\Delta}$
repeated colors in $N(v)$ w.p. at most $\exp(-\Omega(\Delta^{1/2}))$. So the event $E^{\prime}_{2}(v)$ occurs w.p. at most $\exp(-\Omega(\Delta^{1/2}))$.
And we can now conclude with the proof of our main result, which we restate here. See 6.1
Proof.
Both coloring steps (Algorithms 8 and 8) produce a proper coloring with disjoint sets of color, and hence the resulting partial coloring is proper. By Claim 6.11, when $\Delta\geqslant(\log n)^{50}$, the set $\mathcal{A}^{\prime}$ is empty with high probability. Hence, Algorithm 8 ends after $O(1)$ rounds for such values of $\Delta$ and the partial coloring produced verifies the claims of Lemma 6.1. When $\Delta\leqslant(\log n)^{50}$, by Claim 6.12, the connected components of the dependency graph in the post-shattering instance have size at most $\operatorname{\text{{\rm poly}}}(\log n)$. The probability of $E_{CC}(i)$, $E_{1}^{\prime}(v)$, and $E^{\prime}_{2}(v)$ is $\Delta^{-\omega(1)}$ respectively from Lemma 4.4, the Chernoff Bound and Claim 6.13. Since $\operatorname{vbl}(E_{i}(v))\subseteq N^{\leqslant 2}(v)$ and $\operatorname{vbl}(E_{CC}(i))$ consists of every vertex of $S$ within 2 hops from $A_{i}$ in $F$, the dependency degree of the LLL is $\operatorname{\text{{\rm poly}}}(\Delta(F))=\operatorname{\text{{\rm poly}}}(\Delta)$. Hence, they form an LLL that can be solved in $\widetilde{O}(\log^{3}\log n)$ by Theorem 3.
It remains to verify that the claimed properties indeed hold. CC is respected during both pre-shattering and post-shattering so Part (b) holds as well.
Part (a). All the vertices in $V\setminus N^{\leqslant 1}(\operatorname{vbl}(\mathcal{A}^{\prime}))$ have at least $3\sqrt{\Delta}$ colors that appear at least twice in $N(v)\cap S$ because $E_{2}(v)$ does not hold for such vertices and none of their neighbors retract their color. All vertices in $v\in N^{\leqslant 1}(\operatorname{vbl}(\mathcal{A}^{\prime}))$ with $\deg_{S}(v)\geqslant\Delta-3\sqrt{\Delta}$ have $1.05\sqrt{\Delta}$ repeated colors because $E^{\prime}_{2}(v)$ does not hold after the post-shattering step. Overall, Part (a) of Lemma 6.1 holds for all vertices of $S$.
Part (c). Vertices $v\in V\setminus N^{\leqslant 3}(\operatorname{vbl}(\mathcal{A}^{\prime}))$ have at most $11\Delta/20$ colored neighbors because $E_{1}(v)$ does not hold and none of their neighbors gets colored in post-shattering. A vertex $v$ with $E_{1}(v)\in\mathcal{A}^{\prime}$ has at most $4\Delta/5$ colored neighbors after post-shattering because all of its neighbors are uncolored at the beginning of Algorithm 8 and $E_{1}^{\prime}(v)$ does not hold. For every other vertex, i.e., $v\in N^{\leqslant 3}(\operatorname{vbl}(\mathcal{A}^{\prime}))$ but $E_{1}(v)$ did not occur during pre-shattering, the number of colored neighbors is at most
$\deg_{R}(v)+\max\{(4/5)\deg_{S\setminus R}(v),\penalty 10000\ \Delta/20\}$
because $E_{1}^{\prime}(v)$ does not hold and only vertices of $R$ are colored during pre-shattering. If we have that $(4/5)\deg_{S\setminus R}(v)\leqslant\Delta/20$, then since $\deg_{R}(v)\leqslant(11/20)\Delta$, the number colored neighbors is at most $(11/20+1/20)\Delta<19\Delta/20$. Otherwise, using once again that $\deg_{R}(v)\leqslant 11\Delta/20$, the colored degree is at most
$\deg_{R}(v)+(4/5)(\deg_{S}(v)-\deg_{R}(v))$ $\leqslant(4/5)\deg_{S}(v)+(1/5)\deg_{R}(v)$
[…]
7 Coloring Cliques
The goal of this section is to extend the coloring to the cliques in $A_{H}$ and $A_{L}$, respectively. In Algorithm 1, the cliques of $A_{H}$ are colored first, and thus when coloring these nodes we need to ensure that CC is satisfied for all cliques in $A_{L}$. We prove the following:
See 4.8
At a high level, we follow the outline of the approach of [45]. In the first phase, we compute a defective coloring (i.e., a coloring with monochromatic edges) of the cliques to be colored by assigning a random permutation of the colors not used in $\mathrm{All}_{i}$ to the vertices of $A_{i}$. About $\sqrt{\Delta}$ vertices of each clique have a conflict in that they are assigned the same color as an external neighbor (see Lemma 7.2(a)). In the second phase, we perform color swaps: each node with a conflict finds a partner within its own clique with whom it can swap colors to remove the conflict without introducing new ones. This is where the CC property is crucially used.
Our Step 1 is largely similar to the first phase, with the difference of making the conflicts one-way directed in order to simplify the formulation in the shattering framework. We implement the second phase quite differently and do so in two steps (Steps 2 and 3). In our Step 2, we find for each conflicted node a set of candidates for swapping. These sets are designed so that the task of selecting a final swapping partner from the candidate sets is independent across different cliques. In Step 3, we then choose the swap partners locally based on a bipartite matching and apply the swaps in parallel.
7.1 Step 1: Synchronized Color Trial
We compute for each clique $A_{i}$ in $A^{\prime}$ a random permutation of the $|A_{i}|=c-|\mathrm{All}_{i}|$ colors not used by $\mathrm{All}_{i}$ (following [45]) and assign them to the nodes of $A_{i}$. This introduces a limited amount of conflicts, or monochromatic edges. We orient these monochromatic edges for easier shattering formulation: if $\{u,v\}$ is oriented from $u$ to $v$, it means that $v$ should change its color in Step 2.
Definition 7.1.
For a defective coloring $\gamma$ and an orientation of the monochromatic edges of $\gamma$, we define $\mathrm{Unhappy}_{i}$ as the set of vertices of each $A_{i}$ in $A^{\prime}$ with an incoming monochromatic edge Let $\mathrm{Unhappy}=\bigcup_{i}\mathrm{Unhappy}_{i}$ and call a node unhappy if it belongs to $\mathrm{Unhappy}$.
Given the upper bound on external degrees, we expect each clique to have $O(\sqrt{\Delta})$ unhappy nodes. We can achieve that w.h.p. via an LLL formulation.
Lemma 7.2.
There is a $\mathsf{LOCAL}$ algorithm that computes a defective coloring $\gamma$ of $A^{\prime}\subseteq A$ and an orientation of its monochromatic edges such that
  1. (a)
    $|\mathrm{Unhappy}_{i}|\leqslant(10^{8}+1)\sqrt{\Delta}$ for all $A_{i}\in A^{\prime}$, and
  2. (b)
    All cliques of $A^{\prime}$ and all uncolored cliques not in $A^{\prime}$ satisfy the CC property.
The algorithm runs in a single round when $\Delta\geqslant(\log^{50}n)$ and in $\widetilde{O}(\log^{3}\log n)$-rounds otherwise.
The proof is deferred to Section 7.4.
7.2 Step 2: Finding Safe Swaps
We now compute a pool of potential partners for each unhappy vertex. A linear fraction of the nodes in the clique satisfies a minimal requirement. We then cull the pool by probabilistic subsampling. Any candidate that could cause a conflict if some other swap were performed is then eliminated. This eliminates potential inter-clique conflicts. In Step 3, we then select the actual swaps to avoid intra-clique conflicts.
A node $u$ is a swappable candidate for a node $v$ (in the same $A_{i}$) if swapping their colors leaves them both conflict-free when the coloring of the rest of the graph is unchanged. Let $\mathrm{Swappable}_{v}$ be the set of swappable candidates for $v$.
Definition 7.3.
For each $v\in A_{i}\in A^{\prime}$ the set $\mathrm{Swappable}_{v}$ consists of all nodes $u\in A_{i}$ that satisfy
  1. (a)
    $u\not\in\mathrm{Unhappy}_{i}$,
  2. (b)
    $\gamma(u)$ does not appear on an external neighbor of $v$, and
  3. (c)
    $\gamma(v)$ does not appear on an external neighbor of $u$.
We next show that a constant fraction of the nodes of $A_{i}$ are swappable (for any given node $v$). This lemma is the only place where we use the CC property. Its proof is purely deterministic, given the CC property of the clique and the guarantees of Lemma 7.2.
Lemma 7.4.
All $v\in A_{i}\in A^{\prime}$ have $|\mathrm{Swappable}_{v}|\geqslant\Delta/10$.
Proof.
We bound separately from above the number of nodes violating (a), (b) and (c) in Definition 7.3. Nodes violating (a): By Lemma 7.2(a) at most $4\sqrt{\Delta}$ vertices of $A_{i}$ are in $\mathrm{Unhappy}_{i}$.
Nodes violating (b): The external degree of $v$ is bounded by $\sqrt{\Delta}$ (by Lemma 4.1(3,5)), and hence at most $\sqrt{\Delta}$ vertices in $A_{i}$ have a color appearing in the external neighborhood of $v$.
Nodes violating (c): Note that CC holds for the clique $A_{i}$, as it held before Step 1 (recall Lemma 4.8(1)) and was maintained by Step 2 (by Lemma 7.2(b)). Hence, by 4.3 each color (here applied to $\gamma(v)$) appears on an external neighbor (recall, one outside of $A_{i}\cup\mathrm{All}_{i}$) of at most $4\Delta/5$ vertices in $A_{i}$.
Putting everything together,
$|\mathrm{Swappable}_{v}|\geqslant|A_{i}|-4\Delta/5-5\sqrt{\Delta}\geqslant\Delta/10,$
using the lower bound on $|A_{i}|$ from Lemma 4.1(a).
While individual swaps can be done safely (with the sets $\mathrm{Swappable}_{v}$), pairs of swaps could introduce new conflicts. Instead, we seek in this step to truncate the sets to obtain a safe candidate system. Intuitively, a candidate is safe with respect to the candidates of other cliques if it can swap its color obliviously to what happens in other cliques without creating monochromatic edges.
Definition 7.5.
Given a collection $\mathscr{T}=\{T_{v}\subseteq\mathrm{Swappable}_{v}:v\in\mathrm{Unhappy}\}$ of candidate sets for each unhappy node, we call a candidate $u\in\mathrm{Swappable}_{v}$ unsafe for $v$ with regard to $\mathscr{T}$ if one of the following holds:
  1. $i)$
    $v$ has an external neighbor $w$ that has a candidate $w^{\prime}\in T_{w}$ with $\gamma(w^{\prime})=\gamma(u)$;
  2. ii)ii)
    $v$ has an external neighbor $w\in T_{w^{\prime}}$ that is a candidate for a node $w^{\prime}$ with $\gamma(w^{\prime})=\gamma(u)$;
  3. $iii)$
    $u$ has an external neighbor $w\in T_{w^{\prime}}$ that is a candidate for a node $w^{\prime}$ with $\gamma(w^{\prime})=\gamma(v)$.
A candidate that is not unsafe is safe.
We say that $\mathscr{T}$ is a safe candidate system if the sets $T_{v}$ contain only safe candidates for $v$ for all $v\in\mathrm{Unhappy}$.
Definition 7.5 has three instead of four cases, because if $u$ has an external neighbor $w$ that has a candidate $w^{\prime}$ with $\gamma(w^{\prime})=\gamma(v)$, then $w^{\prime}$ is also an unsafe candidate for $w$. Note that whether a candidate is safe or not depends on the candidate sets of nodes in adjacent cliques.
The most involved step of our algorithm for coloring cliques is the construction of a safe candidate system as described in Lemma 7.6. We defer the proof to Section 7.5 to preserve the flow of the paper.
Lemma 7.6 (Subsampling for safe candidates).
There is a $\widetilde{O}(\log^{3}\log n)$ $\mathsf{LOCAL}$ algorithm that, w.h.p., computes a safe candidate system $\mathscr{T}$ such that
  1. 1.
    $|T_{v}|\geqslant\Delta^{17/40}/40$ for each $v\in\mathrm{Unhappy}$,
  2. 2.
    Each $u\in A_{i}\setminus\mathrm{Unhappy}_{i}$ belongs to at most $\Delta^{17/40}/80$ sets $T_{v}$ with $v\in\mathrm{Unhappy}_{i}$, and
  3. 3.
    (Strong version of CC) For each uncolored $A_{j}\not\in A^{\prime}$ and color $x$, at most $2\cdot\Delta^{37/40}$ vertices in $A_{j}$ have a neighbor outside of $A_{j}\cup\mathrm{All}_{j}\cup\operatorname{Big}_{j}^{+}$ that has a candidate of color $x$ or is a candidate for a node of color $x$.
7.3 Step 3: Local Matching
We finally show how to turn a safe candidate system (Lemma 7.6) into a proper coloring, concluding the proof of Lemma 4.8. We first observe that safeness implies that the color selection becomes a fully local problem within each clique.
Observation 7.7.
Suppose an unhappy node $v$ performs a swap of colors with one of its safe candidates $u\in T_{v}$. Afterwards, there are no incoming monochromatic edges to $v$ (nor $u$), independent of any swaps performed in other cliques.
Further, if all unhappy nodes perform swaps with safe candidates and if the swaps are disjoint, then the resulting coloring is proper.
Proof.
For the first claim, consider an external neighbor $w$ of $v$ and let $A_{j}$ be its clique. Since $u$ was swappable for $v$, $\gamma(u)\neq\gamma(w)$, so if $u$ did not perform a swap, there is no conflict. Suppose then that $w$ swapped its color with one of its safe candidates $w^{\prime}$. Then $\gamma(w^{\prime})\neq\gamma(u)$, by Definition 7.5(i), so again there is no conflict. The case of conflicts with $u$ is similar.
For the second claim, observe that by the first claim and since all unhappy nodes perform swaps, all monochromatic edges between cliques are eliminated and no new ones are introduced. The disjointness criteria – that no node participates in more than one swap – ensures that no conflicts occur within the cliques.
The remaining task – deciding on swap partners – then reduces to a local bipartite matching problem. Consider the bipartite graph with vertex bipartition $L=\mathrm{Unhappy}_{i}$ and $R=A_{i}\setminus\mathrm{Unhappy}_{i}$, with an edge between each $v\in L$ and $u\in R$ if and only if $u\in T_{v}$. By Lemma 7.6(1,2), each node in $L$ has at least $\Delta^{17/40}/40$ incident edges, while each node in $R$ has at most $\Delta^{17/40}/80$ incident edges. It follows that for every subset $S\subseteq L$, $|N(S)|\geqslant|S|$. Then, by Hall’s theorem there is a matching $M$ that saturates $L$. We perform exactly the color swaps given by the matching $M$, which yields a proper coloring by 7.7.
To conclude with the proof of Lemma 4.8, it remains to argue that the CC property is maintained for the remaining uncolored cliques. This follows for Step 1 by Lemma 7.2(b) and for Step 3 by Lemma 7.6(3).
7.4 Proof of Lemma 7.2
See 7.2
Input : A collection of uncolored cliques $A^{\prime\prime}\subseteq A^{\prime}$
Output : A coloring $\gamma$ of all the $A_{i}\in A^{\prime\prime}$ and an orientation of its monochromatic edges
1
2Each $A_{i}\in A^{\prime\prime}$ samples a permutation $\pi_{i}$ of the colors not already used on $\mathrm{All}_{i}$
3
4Order the vertices of $A_{i}$ as $v_{1},v_{2},\ldots,v_{|A_{i}|}$ arbitrarily and let $\gamma(v_{j})$ be the color $\pi_{i}(j)$
Orient the resulting monochromatic edges $\{u,v\}$ as follows:
  • if $v\in A_{i}\in A^{\prime}$ and $u$ is not in a clique of $A^{\prime}$, orient the edge towards the vertex $v$, and vice versa;
  • if $v\in A_{i}\in A^{\prime}$ and $u\in A_{j}\in A^{\prime}$, orient the edge from $u$ to $v$ if $i<j$ and in the opposite direction otherwise.
Algorithm 9 Synchronized Color Trial on $A_{i}$
Note that due to Lemma 4.1(b), $|\mathrm{All}_{i}|=c-|A_{i}|$, and hence a color is available for each node in $A_{i}$. In this process, each clique corresponds to exactly one random variable. Recall that a vertex is unhappy (Definition 7.1) if it has at least one incoming monochromatic edge. We consider two bad events for each $A_{i}$: define
  1. $E_{a}(i)$
    : $\mathrm{Unhappy}_{i}$ contains more than $10^{8}\sqrt{\Delta}$ vertices; and
  2. $E_{b}(i)$
    : $A_{i}$ is adjacent to a clique of $A^{\prime}$ or to an uncolored clique not in $A^{\prime}$ for which the CC property was broken.
Clearly, the events $E_{a}(i)$ and $E_{b}(i)$ depend on the random permutation of $A_{i}$ and of every $A_{j}$ adjacent to $A_{i}$. Let us argue that they are rare events, no matter what the coloring outside $A_{i}$ is. We emphasize that Claim 7.8 bounds the probability of both events for pre-shattering and post-shattering because it assumes the coloring of $F-A_{i}$ is adversarial.
Claim 7.8.
Let $A_{i}\in A$ be an uncolored clique, and let the coloring of $F-A_{i}$ be arbitrary. When we run Algorithm 9 with an $A^{\prime\prime}$ that contains $A_{i}$, then events $E_{a}(i)$ and $E_{b}(i)$ occur with probability at most $\exp(-\Omega(\Delta^{1/40}))$.
Proof.
Let $\operatorname{Temp}_{i}\subseteq A_{i}$ consist of the nodes of $A_{i}$ that have a color conflicting with an external neighbor. Clearly, $\mathrm{Unhappy}_{i}\subseteq\operatorname{Temp}_{i}$. The colors are drawn from a set of size $|A_{i}|\geqslant\Delta/2$ and at most $10^{8}\sqrt{\Delta}$ of them conflict with those of external neighbors (Lemma 4.1(3,5)). Hence, each node has conflict w.p. at most $2\cdot 10^{8}/\sqrt{\Delta}$, so the expected size of $\operatorname{Temp}_{i}$ is at most $2\cdot 10^{8}\sqrt{\Delta}$. As Molloy and Reed argue in Lemma 39 of [45] by applying McDiarmid’s inequality (Lemma B.5, with $c=r=1$), $|\operatorname{Temp}_{i}|$ is highly concentrated. Specifically, $E_{a}(i)$ occurs w.p. at most $4\exp(-(\sqrt{\Delta})^{2}/(128\cdot(10^{8}+1)\sqrt{\Delta}))\leqslant\exp(-\Delta^{1/3})$ for $\Delta$ sufficiently large.
Let us now bound ${\mathbb{P}}[E_{b}(i)]$. Fix an arbitrary color $x$ and a clique $A_{j}$. Due to Lemma 4.1 the external degree of $A_{i}$ is bounded by $10^{8}\sqrt{\Delta}$. For a fixed external neighbor $u\in A_{\ell}\in A^{\prime\prime}$ of $A_{j}$, the probability that it receives color $x$ in the random permutation of $A_{j}$ is $1/|A_{j}|\leqslant 2/\Delta$ (by Lemma 4.1(b)). Furthermore, at most one vertex in each clique is assigned color $x$ and the random permutation for different cliques are independent. Hence, we can apply Lemma 4.4 with $Q=10^{8}\sqrt{\Delta}$ as it bounds external degrees from above (by Lemma 4.1(3,5)), and get that $A_{j}$ does not satisfy CC for color $x$ with probability at most $\exp(-\Delta^{1/40})$. By union bound on all colors and cliques adjacent to $A_{i}$, the event $E_{b}(i)$ occurs w.p. at most $\Delta^{3/2+1}\cdot\exp(-\Delta^{1/40})\leqslant\exp(-\Omega(\Delta^{1/40}))$.
To ensure that Lemma 7.2(a,b) hold even when $\Delta$ is small, we employ the shattering framework (Lemma 3.2).
1Run Algorithm 9 on every $A_{i}\in A^{\prime}$
2
3Let $A^{\prime\prime}$ be the set of cliques $A_{i}\in A^{\prime}$ for which $E_{a}(i)$ or $E_{b}(i)$ occurs
4
5Uncolor every vertex in cliques $A_{i}\in A^{\prime\prime}$
6
7Solve the following LLL using Theorem 3:
8
Random Process: Run Algorithm 9 on every $A_{i}\in A^{\prime\prime}$ and orient monochromatic edges between $A^{\prime}\setminus A^{\prime\prime}$ and $A^{\prime\prime}$ toward $A^{\prime\prime}$
Bad Events: $E_{a}(i)$ and $E_{b}(i)$ for all $i\in A^{\prime\prime}$
Algorithm 10 Coloring Algorithm for Lemma 7.2
Proof of Lemma 7.2.
We run Algorithm 10 and show that properties (a) and (b) hold after the post-shattering phase. When $\Delta\geqslant(\log n)^{50}$, it follows directly from Claim 7.8 and the union bound that $A^{\prime\prime}=\emptyset$, and hence the algorithm ends in $O(1)$ rounds and is correct with high probability.
We argue now that the random process in Algorithm 10 of Algorithm 10 along with events $E_{a}(i)$ and $E_{b}(i)$ form an LLL whose dependency graph $G_{\operatorname{dep}}$ has $\operatorname{\text{{\rm poly}}}(\log n)$-sized connected components. The events $E_{a}(i)$ and $E_{b}(i)$ depend on the random choices of at most $10^{9}\Delta^{3/2}$ neighboring almost-cliques, hence the dependency degree of the LLL is at most $10^{18}\Delta^{3}$. By Claim 7.8, every event occurs w.p. $\Delta^{-\omega(1)}$, regardless of the coloring produced by the earlier steps of Algorithm 10. Thus, the set of events described in Algorithm 10 indeed describes an LLL. We next want to apply Lemma 3.2 to bound the size of connected components.
Consider the variable graph $H$ of this random process: it has a vertex (i.e., a variable) $v_{i}$ for each $A_{i}\in A^{\prime}$ and vertices $v_{i}$ and $v_{j}$ are connected iff an edge connects $A_{i}$ and $A_{j}$. The collections $\mathcal{A}$ and $\mathcal{B}$ from Lemma 3.2 are all 1-hop neighborhoods in $H$, i.e., all the $\{v_{i}\}\cup N_{H}(v_{i})$ for $A_{i}\in A^{\prime}$. Define $c_{\mathcal{B}^{\prime}}=0$ so that the set $\mathcal{B}^{\prime}$ in Lemma 3.2 includes the 1-hop neighborhood of every $v_{i}$ with $A_{i}\in A^{\prime\prime}$. This is a superset of $A^{\prime\prime}$ (the set of cliques for which we have an event in post-shattering), so it suffices to bound from above the sizes of the connected components induced by the events corresponding to sets of $\mathcal{B}^{\prime}$. Steps (S1) and (S2) with $c_{1}=2$ and $c_{2}=1$ are direct from the choice of sets $\{v_{i}\}\cup N_{H}(v_{i})$ because it has diameter at most 2 and every $v_{i}$ is included only in the sets of its neighbors. Step (S3) is direct from Algorithm 9. By Claim 7.8, each $A_{i}$ is included in $A^{\prime\prime}$ (i.e., a bad event occurred and $A_{i}$ uncolored its vertices) with probability at most $\exp(-\Omega(\Delta^{1/40}))$. The degree of a $v_{i}$ is at most $10^{9}\Delta^{3/2}$, because every vertex in $A_{i}$ may be adjacent to up to $10^{9}\sqrt{\Delta}$ vertices in other cliques. It implies that for $\Delta$ sufficiently large, we can pick a constant $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1=O(1)$ such that each set of $\mathcal{A}$ is kept in $\mathcal{A}^{\prime}$ with probability at most $\Delta(H)^{-c_{4}}$, i.e., Step (S4). Therefore Lemma 3.2 applies and, w.h.p., we can bound the size of the largest component in the dependency graph of the LLL of Algorithm 10 by some $\operatorname{\text{{\rm poly}}}(\log n)$. Let us conclude the proof by proving that the resulting coloring indeed satisfies (a) and (b). For (b), it simply follows from the fact that after retractions (Algorithm 10), the CC is preserved, and since events $E_{b}(i)$ are avoided in post-shattering, CC is also preserved by this coloring step (with a fresh budget). To see why (a) holds, observe that every $A\notin A^{\prime\prime}$ satisfies (a) after Algorithm 10 (otherwise $E_{a}(i)$ occurs and $A_{i}$ is included in $A^{\prime\prime}$). Importantly, if a monochromatic edge $\{u,v\}$ appears where $v$ was colored by Algorithm 10 and $u$ belongs to a clique of $A^{\prime}$ that did not participate in post-shattering (not in $A^{\prime\prime}$), we orient the edge from $u$ to $v$. So, the coloring of a clique $A_{j}$ that participates in the post-shattering phase can never increase the size of $\mathrm{Unhappy}_{i}$ of an adjacent clique $A_{i}$ that does not participate in the post-shattering phase. Finally, a clique $A_{i}\in A^{\prime\prime}$ colored by Algorithm 10 satisfies (a) at the end of the algorithm, as the event $E_{a}(i)$ would otherwise occur in the post-shattering LLL.
7.5 Proof of Lemma 7.6
See 7.6
Our approach is based on the following random procedure: for each node $v\in\mathrm{Unhappy}$,
  1. 1.
    each node $u\in\mathrm{Swappable}_{v}$ samples itself into a set $S_{v}$ with probability $p=\Delta^{-23/40}$; then
  2. 2.
    all candidates $u\in S_{v}$ that are safe for $v$ w.r.t. $\mathscr{S}=\{S_{v}:v\in\mathrm{Unhappy}\}$ join $T_{v}$,
  3. 3.
    output the safe candidate system $\mathscr{T}=\{T_{v}:v\in\mathrm{Unhappy}\}$.
The sampling is done independently for different nodes $v,v^{\prime}\in\mathrm{Unhappy}$, that is, a node $u$ may be contained in some $S_{v}$ but not in $S_{v^{\prime}}$. We emphasize that, when $\Delta\leqslant\operatorname{\text{{\rm poly}}}(\log n)$, that process does not succeed everywhere; hence, we use (again) the shattering technique. The pre-shattering step will produces large enough sets $T_{v}^{pre}$ of safe candidates for most cliques. The post-shattering step then computes sets $T_{v}$ for the remaining cliques based on the same LLL and Theorem 3.
Before proving Lemma 7.6, we prove the following technical claim — that we will use both in pre- and post-shattering — to bound the number of unsafe candidates. The set of cliques $A^{\prime\prime}\subseteq A^{\prime}$ will be all cliques of $A^{\prime}$ in pre-shattering, and, in post-shattering, all the cliques for which we made retractions. Claim 7.9 shows that a set $S$ of candidates contains only a constant fraction of unsafe candidates with regard to candidates sampled by neighboring cliques.
Note that the random process in Claim 7.9 depends only on the outside randomness, i.e., the set $S$ is fixed deterministically. We also emphasize that the success probability depends on the degree $k$ of the outside vertices into $S$. We abuse notation slightly and write $A^{\prime}\setminus A_{i}$ for $A^{\prime}\setminus\{A_{i}\}$.
Claim 7.9.
Let $A^{\prime\prime}\subseteq A^{\prime}$ be a subset of the cliques we wish to color. Suppose that every node in $Swappable_{w}$ for some $w\in\mathrm{Unhappy}_{i}$ with $A_{i}\in A^{\prime\prime}$ joins the candidate set $S_{w}$ w.p. $p<10^{-10}\Delta^{-1/2}$. Consider a fixed $v\in A_{i}$ and $S\subseteq\mathrm{Swappable}_{v}$, and call $k$ the maximum number of neighbors in $S$ that vertices $w\in A_{j}\in A^{\prime}\setminus A_{i}$ can have, i.e., $k=\max_{w\in A_{j}\in A^{\prime}\setminus A_{i}}|N(w)\cap S|$. Then, $S$ contains fewer than $|S|/20$ unsafe candidates for $v$ w.r.t. $\mathscr{S}=\{S_{w}:w\in\mathrm{Unhappy}\}$ w.p. at least $1-\exp\left(-\Omega\left(|S|/k\right)\right)$.
Proof.
Recall from Definition 7.5 that a node $u\in S$ can be an unsafe candidate for $v$ for three reasons. We first bound the number of candidates that are bad due to $(i)$ or $(ii)$. Then, we analyze the number of unsafe candidates due to $(iii)$.
Counting unsafe candidates due to rule $(i)$ or $(ii)$: For $u\in S$, let $X_{u}$ be the indicator random variable equal to one iff either $(i)$ or $(ii)$ in Definition 7.5 holds w.r.t. the sets of candidates $\{S_{w}\}_{w}$. To have $X_{u}=1$ by rule $(i)$, node $v$ must have an external neighbor $w\in\mathrm{Unhappy}$ with a candidate colored $\gamma(u)$ and sampled in $S_{w}$. To have $X_{u}=1$ by rule $(ii)$, node $v$ must have an external neighbor in some $S_{w}$ where $w$ is colored $\gamma(u)$. For a given external neighbor, at most one of those can occur, each w.p. at most $2p$; thus, $\mathbb{E}[X_{u}]\leqslant 10^{8}\sqrt{\Delta}\cdot p<1/100$ (recall that by Lemma 4.1, node $v$ has at most $10^{8}\sqrt{\Delta}$ external neighbors). The $\{X_{u}\}_{u\in A_{i}}$ are independent because each $X_{u}$ depends only on the randomness of nodes colored $\gamma(u)$, and all of the nodes of $A_{i}$ are colored differently. We apply the Chernoff Bound to the sum of the $X_{u}$ over $u\in S$ and obtain that the probability of having more than $|S|/20$ nodes removed is bounded from above by $\exp(-\Omega(|S|))$.
Counting unsafe candidates due to rule $(iii)$: For $u\in S$, let $Y_{u}$ be the indicator random variable equal to one iff $(iii)$ in Definition 7.5 holds for $u$ w.r.t. sets $\mathscr{S}$. We have that $Y_{u}=1$ only if there exists an external neighbor $w\in A_{j}\in A^{\prime\prime}$ of $u$ that was sampled in the candidate set of the node colored $\gamma(v)$ in $A_{j}$ (if any). Hence, by union bound, $\mathbb{E}[Y_{u}]\leqslant p\cdot 10^{8}\sqrt{\Delta}<1/100$. Note that $Y_{u}$ and $Y_{u^{\prime}}$ can be dependent for $u\neq u^{\prime}\in S$ as $u$ and $u^{\prime}$ may have the same external neighbor $w$ that is a candidate for a node colored $\gamma(v)$. However, the variables $\{Y_{u}:u\in S\}$ form a read-$k$ family for $k=\max_{w\in A_{j}\in A^{\prime\prime}\setminus A_{i}}|N(w)\cap S|$, w.r.t. the independent boolean variables $Z_{w}$ equal to one iff $w$ was sampled in the candidate set of the node colored $\gamma(v)$ in its clique. Each variable $Z_{w}$ influences only the $Y_{u}$ for which the edge $\{u,w\}$ exists, which amounts to at most $k$ variables definition of $k$. Hence, the read-$k$ bound implies that
${\mathbb{P}}\left[\sum_{u\in S}Y_{u}>|S|/20\right]\leqslant{\mathbb{P}}\left[\left|\sum_{u\in S}Y_{u}-\mathbb{E}\left[\sum_{u\in S}Y_{u}\right]\right|>|S|/30\right]\leqslant\exp\left(-\Omega(|S|/k)\right)$
Overall, the number of unsafe candidates for $v$ in $S$ is bounded from above by the sum of $X_{u}+Y_{u}$ over $u\in S$. By union bound, the sum of the $X_{u}$ and the sum of the $Y_{u}$ are both smaller than $|S|/20$ w.p. at least $1-2\exp(-\Omega(|S|/k))$; hence the claim.
We are now ready to stick the analysis above together to prove Lemma 7.6.
Pre-shattering.
For each node $v\in\mathrm{Unhappy}$, each node $u\in\mathrm{Swappable}_{v}$ samples itself into the set $S_{v}$ with probability $p=\Delta^{-23/40}$. Denote by $\mathscr{S}=\{S_{v}:v\in\mathrm{Unhappy}\}$ the randomly generated collection of candidate sets. Let us define a some bad events for every $A_{i}\in A^{\prime}$:
  1. $B_{1}(i)$:
    Some $v\in\mathrm{Unhappy}_{i}$ has fewer than $\Delta^{17/40}/20$ safe candidates w.r.t. $\mathscr{S}$;
  2. $B_{1}^{\prime}(i)$:
    Some $w\notin A_{j}\in A^{\prime}\setminus A_{i}$ has at least $\Delta^{1/10}$ neighbors in some $S_{v}$ where $v\in\mathrm{Unhappy}_{i}$;
  3. $B_{2}(i)$:
    Some $u\in A_{i}\setminus\mathrm{Unhappy}$ belongs to more than $\Delta^{17/40}/80$ sets in $\mathscr{S}$;
  4. $B_{3}(i)$:
    For some uncolored $A_{j}\not\in A^{\prime}$ adjacent to $A_{i}$ and color $x$, more than $\Delta^{37/40}$ vertices in $A_{j}$ have a neighbor outside of $A_{j}\cup\mathrm{All}_{j}\cup\operatorname{Big}_{j}^{+}$ that either has a candidate of color $x$ or is a candidate for a node of color $x$;
  5. $B_{4}(i)$:
    The set $\mathrm{Swappable}_{v}$ for some $v\in\mathrm{Unhappy}_{i}$ contains $\Delta/20$ unsafe candidates w.r.t. $\mathscr{S}$.
It should be clear that those events depend on the colors and random decisions within $O(1)$ distance. Let us now argue that they occur with probability at most $\exp(-\Delta^{1/40})$.
Claim 7.10.
Event $B_{1}(i)$ and $B_{1}^{\prime}(i)$ occur with probability at most $\exp(-\Delta^{1/40})$.
Proof.
We argue about a given $v\in\mathrm{Unhappy}_{i}$. The claim then follows by union bound over all such $v$.
We first argue that the probability on the randomness of $S_{v}$ that $S_{v}$ either (1) contains fewer than $\Delta^{17/40}/20$ nodes, or (2) has $\Delta^{1/10}$ edges to some $w\in A_{j}\in A^{\prime}\setminus A_{i}$ is small. Fix a node $v\in A_{i}$, and recall that $|\mathrm{Swappable}_{v}|\geqslant\Delta/10$ (Lemma 7.4). Hence, the expected size of $S_{v}$ is $p|\mathrm{Swappable}_{v}|=\Delta^{17/40}/10$. By Chernoff, we obtain that $|S_{v}|$ has fewer than $\Delta^{17/40}/19$ w.p. at most $\exp(-\Omega(\Delta^{17/40}))$. Meanwhile, a node $w\notin A_{j}\in A^{\prime}\setminus A_{i}$ has at most $10^{8}\sqrt{\Delta}$ neighbors in $A_{i}$ and thus $o(1)$ expected neighbors in $S_{v}$. By Chernoff, node $w$ has more than $\Delta^{1/10}$ neighbors in $S_{v}$ w.p. at most $\exp(-\Omega(\Delta^{1/10}))$. By union bound on all $v\in\mathrm{Unhappy}_{i}$ and $w$ adjacent to $A_{i}$, we get that ${\mathbb{P}}[B_{1}^{\prime}(i)]\leqslant 10^{16}\Delta^{2}\exp\left(-\Omega(\Delta^{1/10})\right)\leqslant\exp\left(-\Delta^{1/40}\right)$.
To bound the probability of $B_{1}(i)$, it suffices to bound the probability of $B_{1}(i)\cap\overline{B_{1}^{\prime}(i)}\cap\overline{E(i)}$, where $E(i)$ is the event that some $S_{v}$ with $v\in\mathrm{Unhappy}_{i}$ contains fewer than $\Delta^{17/40}/19$ nodes. Consider any given realization of the sets $S_{v}$ for $v\in\mathrm{Unhappy}_{i}$ such that $\overline{B_{1}^{\prime}(i)}\cap\overline{E(i)}$ holds; then the random process is the one described in Claim 7.9 with $A^{\prime\prime}=A^{\prime}\setminus A$ and $S=S_{v}$ for some fixed $v\in\mathrm{Unhappy}_{i}$. Since $S_{v}$ is such that $\overline{B_{1}^{\prime}(v)}\cap\overline{E(i)}$ holds, it verifies the assumption of Claim 7.9 with $k=\Delta^{1/10}$, and thus $S_{v}$ contains at most $|S_{v}|/20$ unsafe candidates. Under $\overline{E(i)}$, it implies that $S_{v}$ contains at least $\Delta^{17/40}/20$ safe candidates. By union bound, the probability (on the randomness of $S_{w}$ for $w\notin\mathrm{Unhappy}\setminus\mathrm{Unhappy}_{i}$) that some $v$ has too few safe candidates is
$|\mathrm{Unhappy}_{i}|\exp\left(-\Omega\left(\frac{\Delta^{17/40}}{\Delta^{1/10}}\right)\right)\leqslant\exp(-\Omega(\Delta^{13/40}))\ .$
Since this holds for any realization of $\{S_{v}:v\in\mathrm{Unhappy}_{i}\}$ where $\overline{B_{1}^{\prime}(i)}\cap\overline{E(i)}$ holds, we get a bound on the probability of $B_{1}(v)$ as ${\mathbb{P}}[B_{1}(i)]\leqslant{\mathbb{P}}[B_{1}(i)\cap\overline{B_{1}^{\prime}(i)}\cap\overline{E(i)}]+{\mathbb{P}}[B_{1}^{\prime}(i)]+{\mathbb{P}}[E(i)]\leqslant\exp(-\Delta^{1/40})$.
Claim 7.11.
Event $B_{2}(i)$ occurs with probability at most $\exp(-\Delta^{1/40})$.
Proof.
In expectation, a vertex is sampled into $\Delta^{-23/40}|\mathrm{Unhappy}_{i}|=o(1)$ sets of $\mathscr{S}$. The claim follows via a Chernoff bound.
Claim 7.12.
Event $B_{3}(i)$ occurs with probability at most $\exp(-\Omega(\Delta^{1/40}))$
Proof.
Fix a clique $A_{j}$ (either in $A^{\prime}$ or uncolored) and a color $x$. For a fixed external neighbor $u\in A_{\ell}\neq A_{j}$, the vertex with color $x$ in $u$’s clique (if any) gets sampled into $u$’s candidate set $S_{u}$ or samples $u$ into its candidate set with probability at most $p\leqslant\Delta^{-23/40}$. On the other hand, by assumption in Lemma 4.8, the external degrees of uncolored cliques not in $A^{\prime}$ is at most $30\Delta^{1/4}$ (recall that in this step, we do not need to guarantee that CC is maintained by cliques of $A^{\prime}$, in particular, cliques of $A_{H}$ do not need to verify Lemma 7.6(3)). And since the sampling is independent, it is easy to verify that (P7.1) holds with $Q=30\Delta^{1/4}$, thus that $B_{3}(i)$ occurs because of clique $A_{j}$ with probability at most $\exp(-\Delta^{1/40})$ by Lemma 4.4. By union bound on the colors and the at most $10^{8}\Delta^{3/2}$ cliques adjacent to $A_{i}$, we obtain the desired probability bound.
Claim 7.13.
Event $B_{4}(i)$ holds w.p. at most $\exp(-\Omega(\Delta^{1/2}))$.
Proof.
Fix some $v\in\mathrm{Unhappy}_{i}$. Recall that every $w\in A_{j}\in A^{\prime}\setminus A_{i}$ has at most $k:=10^{8}\sqrt{\Delta}$ external neighbors, thus at most $k$ neighbors in $\mathrm{Swappable}_{v}$. Hence Claim 7.9 with $S=\mathrm{Swappable}_{v}$ and $k$ as above implies that $\mathrm{Swappable}_{v}$ contains more than $|\mathrm{Swappable}_{v}|/20$ unsafe candidates w.r.t. sets $\{S_{w}:w\in\mathrm{Unhappy}\}$ with probability at most $\exp(-\Omega(|\mathrm{Swappable}_{v}|/k))\leqslant\exp(-\Omega(\Delta^{1/2}))$, where the last inequality uses that $|\mathrm{Swappable}_{v}|\geqslant\Delta/10$ (Lemma 7.4). The claim follows by union bound on all $v\in\mathrm{Unhappy}_{i}$.
Retractions & Candidates $T^{pre}_{v}$.
If $\Delta\geqslant(\log n)^{50}$, w.h.p., none of the bad events occur and we are done. We henceforth assume that $\Delta\leqslant(\log n)^{50}$. To find enough safe candidates even in cliques where bad events occur, we retract some sets of candidates. Let $A^{\prime\prime}\subseteq A^{\prime}$ be the set of cliques $A_{i}\in A^{\prime}$ for which
  1. 1.
    some bad event $B_{j}(i)$ for $j\in[4]$ occurred, or
  2. 2.
    $A_{i}$ is incident to some $A_{i^{\prime}}$ for which a bad event $B_{j}(i^{\prime})$ occurred.
For all the cliques $A_{i}\in A^{\prime}\setminus A^{\prime\prime}$ and $v\in\mathrm{Unhappy}_{i}$, let $T^{pre}_{v}$ be the subset of $S_{v}$ containing all the safe candidates w.r.t. the sets $S_{v}$ sampled during pre-shattering. Recall that removing candidates does not turn safe candidates into unsafe ones; hence the sets $T^{pre}_{v}$ contain only safe candidates w.r.t. $\mathscr{T}^{pre}=\{T_{v}^{pre}:v\in\mathrm{Unhappy}_{i},A_{i}\in A^{\prime}\setminus A^{\prime\prime}\}$. We say that the sets $S_{v}$ for $v\in A_{i}\in A^{\prime\prime}$ were retracted as we henceforth ignore them.
Post-shattering.
To compute candidates for sets of $A^{\prime\prime}$ we consider the LLL induced by the same random sampling process as before except that it runs only in $A^{\prime\prime}$: every vertex $u\in\mathrm{Swappable}_{v}$ for $v\in\mathrm{Unhappy}_{i}$ with $A_{i}\in A^{\prime\prime}$ samples itself into a set $S_{v}^{\prime}$ with probability $p=\Delta^{-23/40}$. We call the sets of candidates $S_{v}^{\prime}$ rather than $S_{v}$ to emphasize that they were sampled during post-shattering. The random collection of candidate sets thereby produced is called $\mathscr{S}^{\prime}$. Consider the following bad events: for each $A_{i}\in A^{\prime\prime}$ or $A_{i}$ adjacent to some $A_{j}\in A^{\prime\prime}$,
  1. $P_{1}(i)$:
    Some $v\in\mathrm{Unhappy}_{i}$ has fewer than $\Delta^{17/40}/40$ safe candidates w.r.t. $\mathscr{S}^{\prime}\cup\mathscr{T}^{pre}$;
and for each $A_{i}\in A^{\prime\prime}$, let
  1. $P_{2}(i)$:
    Some $u\in A_{i}\setminus\mathrm{Unhappy}_{i}$ is a candidate for more than $\Delta^{17/40}/80$ nodes in $\mathrm{Unhappy}_{i}$;
  2. $P_{3}(i)$:
    For some uncolored $A_{j}\not\in A^{\prime}$ adjacent to $A_{i}$ and color $x$, more than $2\cdot\Delta^{37/40}$ vertices in $A_{j}$ have a neighbor outside of $A_{j}\cup\mathrm{All}_{j}\cup\operatorname{Big}_{j}^{+}$ that either has a candidate of color $x$ or is a candidate for a node of color $x$.
We emphasize that the safety of the candidates is w.r.t. $\mathscr{T}^{pre}$ obtained from pre-shattering and $\mathscr{S}^{\prime}$ sampled in post-shattering. Let us first argue that $P_{1}(i)$, $P_{2}(i)$ and $P_{3}(i)$ are rare events even in the presence of the fixed pre-shattering sets $S_{v}$.
Claim 7.14.
For $A_{i}\in A^{\prime\prime}$ or $A_{i}$ adjacent to some $A_{j}\in A^{\prime\prime}$, the event $P_{1}(i)$ occurs w.p. at most $\exp(-\Omega(\Delta^{1/40}))$.
Proof.
Suppose first that $A_{i}\in A^{\prime\prime}$. If $B_{4}(i)$ holds, then $A_{i}$ and all the adjacent cliques belong to $A^{\prime\prime}$. Hence the same argument as in Claim 7.10 (for $B_{1}(i)$) implies that $P_{1}(i)$ holds with probability at most $\exp(-\Omega(\Delta^{1/40}))$. Otherwise, suppose that $\overline{B_{4}(i)}$ holds, hence for all $v\in\mathrm{Unhappy}_{i}$, the sets $\mathrm{Swappable}_{v}$ contain at least $\Delta/20$ safe candidates w.r.t. $\mathscr{T}^{pre}$. We follow the same analysis as for Claim 7.10, but we only have half as many vertices to pick from. For a fixed $v\in\mathrm{Unhappy}_{i}$, w.p. at least $1-\exp(-\Omega(\Delta^{17/40}))$, at least $\Omega(\Delta^{17/40})$ nodes $u\in\mathrm{Swappable}_{v}$ that are safe candidates w.r.t. $\mathscr{T}^{pre}$ and get sampled into $S_{v}^{\prime}$. Also, w.p. at least $1-10^{8}\Delta^{3/2}\exp(-\Omega(\Delta^{1/10}))$, every $w\in A_{j}\in A^{\prime\prime}\setminus A_{i}$ has fewer than $\Delta^{1/10}$ neighbors in $S_{v}^{\prime}$. Now, Claim 7.9 with $S$ the set of safe candidates (w.r.t. $\mathscr{T}^{pre}$) in $S_{v}^{\prime}$ and $k=\Delta^{1/10}$ implies that w.p. $1-\exp(-\Omega(\Delta^{17/40}/k))\geqslant 1-\exp(-\Omega(\Delta^{1/40}))$. Overall, when $A_{i}\in A^{\prime\prime}$, the event $P_{1}(i)$ holds w.p. at most $\exp(-\Omega(\Delta^{1/40}))$.
Suppose now that $A_{i}\notin A^{\prime\prime}$ but is adjacent to some $A_{j}\in A^{\prime\prime}$. Fix some $v\in\mathrm{Unhappy}_{i}$. Since $A_{i}\notin A^{\prime\prime}$, the event $B_{1}^{\prime}(i)$ does not hold, thus every $w$ in some $A_{j}\in A^{\prime\prime}$ has fewer than $\Delta^{1/10}$ neighbors in $T^{pre}_{v}$. By Claim 7.9 with $S=T_{v}^{pre}$, the set $T_{v}^{pre}$ contains at most $|T_{v}^{pre}|/20$ unsafe candidates w.r.t. $\mathscr{S}^{\prime}$ w.p. at least $1-\exp(-\Omega(\Delta^{17/40}/\Delta^{1/10}))=1-\exp(-\Omega(\Delta^{13/40}))$. Since $T_{v}^{pre}$ contained at least $\Delta^{17/40}/20$ nodes (otherwise $B_{1}(i)$ would hold and $A_{i}\in A^{\prime\prime}$), the set $T_{v}^{pre}$ contains at least $(1-1/20)\Delta^{17/40}/20\geqslant\Delta^{17/40}/40$ safe candidates w.r.t. sets $\mathscr{T}^{pre}\cup\mathscr{S}^{\prime}$ with probability at least $1-\exp(-\Omega(\Delta^{13/40}))$. The claim follows by union bound over all $v\in\mathrm{Unhappy}_{i}$.
The bound on the probability of $P_{2}(i)$ and $P_{3}(i)$ follows the same argument as in pre-shattering. Note that we allow the post-shattering phase to have a fresh budget.
Claim 7.15.
Let $A_{i}\in A^{\prime\prime}$. Then $P_{2}(i)$ and $P_{3}(i)$ each occur w.p. at most $\exp(-\Delta^{1/40})$.
Proof.
See Claims 7.11 and 7.12.
As for the synchronized color trial, we can see the post-shattering random process as having one random variable for each clique $A_{i}\in A^{\prime\prime}$. For each event $P_{1}(i)$, $P_{2}(i)$, and $P_{3}(i)$, its set of variables corresponds to the clique $A_{i}$ and each neighboring clique of $A^{\prime\prime}$. Therefore the dependency degree of those events is at most $10^{16}\Delta^{3}$. Since each event occurs with probability at most $\exp(-\Omega(\Delta^{1/40}))$, it indeed defines an LLL that can be solved by Theorem 3.
The shattering argument showing that the connected components of the dependency graph have size at most $\operatorname{\text{{\rm poly}}}(\log n)$ is identical to that of Lemma 7.2, so we do not repeat it here. It follows that computing the sets $\mathscr{S}^{\prime}$ such that none of the events $P_{1}(i)$, $P_{2}(i)$, or $P_{3}(i)$ occur takes $\widetilde{O}(\log^{3}N)=\widetilde{O}(\log^{3}\log n)$ rounds.
The algorithm outputs the safe candidate system $\mathscr{T}=\{T_{v}:v\in\mathrm{Unhappy}\}$ where $T_{v}$ is the set of safe candidates for $v$ in $T_{v}^{pre}$ w.r.t. $\mathscr{T}^{pre}\cup\mathscr{S}^{\prime}$ if $v\in A_{i}\in A^{\prime}\setminus A^{\prime\prime}$ (it succeeded in pre-shattering), and $T_{v}$ is the set of safe candidates for $v$ in $S^{\prime}_{v}$ w.r.t. $\mathscr{T}^{pre}\cup\mathscr{S}^{\prime}$ if $v\in A_{i}\in A^{\prime\prime}$ (it retracted its candidates after pre-shattering). This choice of candidate sets satisfies Lemma 7.6(3) because after retractions in the pre-shattering, none of the $B_{3}(i)$ holds, and after post-shattering, none of the $P_{3}(i)$ hold. Lemma 7.6(1,2) hold after retractions for every clique $A_{i}\notin A^{\prime\prime}$ because $B_{1}(i)$ and $B_{2}(i)$ do not occur and retracting candidates in neighboring cliques cannot make them occur. If $A_{i}$ has neighboring cliques in $A^{\prime\prime}$, the event $P_{1}(i)$ is included in post-shattering so Lemma 7.6(1) continues to hold after selecting sets $\mathscr{S}^{\prime}$. For $A_{i}\in A^{\prime\prime}$, Lemma 7.6(1,2) hold after post-shattering because neither of $P_{1}(i)$ nor $P_{2}(i)$ can occur. $\blacksquare$
8 Ultrafast Coloring High-Degree Graphs
In this section, we show that the coloring can be computed in $O(\log^{*}n)$ rounds when $\Delta=\Omega(\log^{50}n)$, rather than $O(\log\Delta\cdot\operatorname{\text{{\rm poly}}}\log\log n)$ many rounds when using the iterative approach described in Section 5  [^8] We have not tried to optimize the constant exponent ”50”.
Several parts of our algorithm actually run in $O(1)$ rounds when $\Delta$ is large (e.g., when $\Delta\geqslant\log^{50}n$) simply because the probability that each bad event from our LLLs occurs is $\operatorname{\text{{\rm poly}}}(n)\exp(-\Omega(\Delta^{1/40})\leqslant\operatorname{\text{{\rm poly}}}(n)\exp(-\Omega(\log^{5/4}n))\leqslant 1/\operatorname{\text{{\rm poly}}}(n)$. This applies to the coloring of the cliques (see ColorCliques from Section 7), slack generation (see SlackGeneration from Section 6.3), and vertex splitting (Lemma 6.3 from Section 6).
The remaining hurdle is that the algorithm ColorWithMuchSlack uses $O(\log\Delta)$ iterations. In this section, we explain how it can be replaced by an algorithm of [35] for the $(\deg+1)$-list-coloring problem. It runs in $O(\log^{*}n)$ rounds when $\Delta>(\log n)^{3}$. Recall that the input to ColorWithMuchSlack is a $\Pi$-ous subgraph $H$.
See 4.6
If it was not for the CC constraint, one could simply run the $(\deg+1)$-list-coloring algorithm of [35] to extend the coloring to $H$. Given an instance of $(\deg+1)$-list-coloring, it is not possible in general to ensure that the CC constraint will be maintained. Here, we use that, thanks to $\Pi$, the nodes of $H$ have lists significantly larger than their degree.
In more detail, we observe that all the coloring steps of [35] are based on three forms of randomized color trials that we have already considered:
  • Random color trials (RCT), when nodes try a random color from their palette,
  • Multi-color trials (MCT), when nodes try multiple colors from their palette, and
  • Synchronized color trials (SCT), when the nodes of an almost-clique are assigned a permutation of the colors in a palette.
Using $\Pi$ and Lemma 4.4, we can show that each of the $O(\log^{*}n)$ steps of [35] maintain the CC property with high probability.
Proof of Lemma 4.6.
One technical issue needs to be resolved before applying the algorithm of [35]. We have lower bounds on the palettes of nodes but not explicit lower bounds on their degrees. To address this, we construct the graph $H^{\prime}$ from $H$ by adding $|L(v)|-(\deg_{H}(v)+1)$ dummy vertices incident to each $v$ each with two arbitrary colors. Note that every vertex of $H$ has degree at least $U\cdot\Delta^{0.22}$ in $H^{\prime}$ (by $\Pi$(b)).
The first step of [35] computes (deterministically) a vertex partition of $H^{\prime}$ into sets $V_{sp},C_{1},C_{2},\ldots,C_{q}$ with the properties described in [1, Lemma 4.2]. The algorithm then colors $H^{\prime}$ in two steps (see Algorithm 8 in [35]): first color $H^{\prime}[V_{sp}]$ with [35, Algorithm 4], and the color $H^{\prime}[C_{1}\cup\ldots\cup C_{q}]$ with [35, Algorithm 5].
Both algorithms begin with a random color trial — for the sparse vertices, only some vertices try a random color. See [35, Algorithm 3]. Since every vertex of $H$ samples one color out of a list of $U\Delta^{0.22}$ while the external degree of uncolored cliques is upper bounded by $U$ (by $\Pi$), Lemma 4.4 and union bound implies that CC is not maintained for some clique with probability at most $n\Delta\exp(-\Delta^{1/40})<1/\operatorname{\text{{\rm poly}}}(n)$. Note that the dummy nodes do not verify $\Pi$, as their lists are of size two, but they are not adjacent to the cliques of $F$.
To extend the coloring to $H[V_{sp}]$, [35] uses an algorithm called SlackColor consisting of $O(\log^{*}n)$ iterations of of MCT. See [35, Lemma 1] and [35, Algorithm 10] with $s_{\min}=U\Delta^{0.22}\geqslant\Delta^{0.47}$ and $\kappa=1/2$. More precisely, each vertex of $V_{sp}$ (not dummy vertices) picks up to $\log n\leqslant\Delta^{1/50}$ colors u.a.r. from their list. By $\Pi$, all palettes contain at least $U\Delta^{0.22}$ colors. Hence, any given color is picked with probability proportional to at most
$\frac{\log n}{U\Delta^{0.22}}\leqslant\frac{1}{U\Delta^{0.22-1/50}}\leqslant\frac{1}{U\Delta^{1/5}}\ .$
Hence, we can apply Lemma 4.4 as before to reason that each step of MCT maintains CC with probability at least $1-\exp(-\Omega(\sqrt{s_{\min}}))-\Delta\exp(-\Omega(s_{\min}))\geqslant 1-1/\operatorname{\text{{\rm poly}}}(n)$.
The analysis of slack color for nodes of $H[C_{1}\cup\ldots\cup C_{q}]$ is the same as above. We emphasize that we do not need the put-aside set, Step 3 and 7 in [35, Algorithm 5], because $\Pi$ ensures that all nodes of $H$ have more than $\log^{2}n$ colors in their lists. After slack generation, [35, Algorithm 5] runs a synchronized color trial before it calls SlackColor on (most) nodes. See [35, Algorithm 7]. In this algorithm, (most of) the vertices in each $C_{i}$ receive a random color from a selected leader. As such, the probability that any given color is tried by a vertex in some $C_{i}$ is $O(1/|C_{i}|)$. Since the vertices have $\deg(v)=|L(v)|-1\geqslant U\Delta^{0.22}$, where the equality holds because of the dummy vertices and the inequality because of $\Pi$, an almost-clique $C_{i}$ contains at least $\Omega(U\Delta^{0.22})$ vertices. And thus the CC constraint is maintained by the SCT as every vertex tries gets some color $x$ with probability at most $O(1/|C_{i}|)\leqslant O(\frac{1}{U\Delta^{0.22}})\leqslant\frac{1}{U\Delta^{1/5}}$.
Appendix A Graph Decomposition by Molloy-Reed
Based on Molloy and Reed [45], Bamas and Esperet give the following structural decomposition of the graph $F$.
Lemma A.1 (Lemma 4.5 of [4], based on Lemma 12 in [45]).
We can construct from $H$ in $O(1)$ rounds of $\mathsf{LOCAL}$ a graph $F$ of maximum degree at most $10^{9}\Delta$ (such that a $c$-coloring of $H$ can be deduced from any $c$-coloring of $F$ in $O(1)$ rounds) and find a partition of the vertices of $F$ into $S,B,A_{1},\ldots,A_{t}$ such that:
  • (a)
    Every $A_{i}$ is a clique with $c-10^{8}\sqrt{\Delta}\leqslant|A_{i}|\leqslant c$.
  • (b)
    Every vertex of $A_{i}$ has at most $10^{8}\sqrt{\Delta}$ neighbors in $F-A_{i}$.
  • (c)
    There is a set $\mathrm{All}_{i}\subseteq B$ of $c-|A_{i}|$ vertices which are adjacent to all of $A_{i}$. Every other vertex of $F-A_{i}$ is adjacent to at most $\tfrac{3}{4}\Delta+10^{8}\sqrt{\Delta}$ vertices of $A_{i}$.
  • (d)
    Every vertex of $S$ either has fewer than $\Delta-3\sqrt{\Delta}$ neighbors in $S$ or has at least $900\Delta^{3/2}$ non-adjacent pairs of neighbors within $S$.
  • (e)
    Every vertex of $B$ has fewer than $c-\sqrt{\Delta}+9$ neighbors in $F-\bigcup_{j}A_{j}$.
  • (f)
    If a vertex $v\in B$ has at least $c-\Delta^{3/4}$ neighbors in $F-\bigcup_{j}A_{j}$, then there is some $i$ such that: $v$ has at most $c-\sqrt{\Delta}+9$ neighbors in $F-A_{i}$ and every vertex of $A_{i}$ has at most $30\Delta^{1/4}$ neighbors in $F-A_{i}$.
  • (g)
    For every $A_{i}$, every two vertices outside of $A_{i}\cup\mathrm{All}_{i}$ which have at least $2\Delta^{9/10}$ neighbors in $A_{i}$ are joined by an edge of $F$.
Given Lemma A.1, the proof of Lemma 4.1 is relatively straightforward.
See 4.1
Proof.
Lemma A.1 decomposes the graph $F$ into $S$,$B$, a collection of cliques $A=\bigcup_{j}A_{j}$, and mentions a set $\mathrm{All}_{i}\subseteq B$ for each clique $A_{i}$. The more fine-grained decomposition of this lemma follows the algorithm of [45]. Let $B_{H}\subseteq B$ be the vertices with at most $c-\Delta^{3/4}$ neighbors in $F\setminus\bigcup_{j}A_{j}$. The set $B_{L}$ equals $B\setminus B_{H}$. $A_{H}$ contains all cliques $A_{i}\in A$ such that each vertex of $A_{i}$ has at least $30\Delta^{1/4}$ neighbors outside of $A_{i}\cup\mathrm{All}_{i}$. $A_{L}$ equals $A\setminus A_{H}$. We now prove each of the properties separately.
  1. 1.
    $S$: The decomposition of the graph $F$ in Lemma A.1 leaves a crucial property open, namely that the vertices in $S$, the sparse vertices, have a degree of at most $\Delta$ in their induced subgraph. Property Lemma A.1(d) only specifies that a vertex in $S$ has a degree of at most $\Delta-3\sqrt{\Delta}$ or has at least $900\Delta^{3/2}$ non-edges in its neighborhood. But for the upcoming coloring step of coloring the vertices in $S$ it is crucial that their degree is also bounded above by $\Delta$ in any case. However, in their proof of Lemma A.1 in [49] the authors clearly state that this property also applies.
    We next reason why we can claim $9.9\cdot 10^{5}\cdot\sqrt{\Delta}$ non-edges in the neighborhood of vertices in $S$, instead of only $900\sqrt{\Delta}$ as claimed in Lemma A.1. The structural decomposition is defined in Lemma 12 of [45]. It begins with the decomposition from [44] where nodes of $S$ are $d$-sparse for $d=10^{6}\sqrt{\Delta}$, meaning that their neighborhood has at most $\binom{\Delta}{2}-d\Delta$ edges. As they claim in the proof of Lemma 12(d), the only part where they add edges in $F[S]$ that did not exist in $G$ is in Modification 2 (see Section 5.3, page 161). Importantly, this step adds edges only between big-neighbors. In the proof of Lemma 12 (section 5.3, page 163), Molloy and Reed argue that $v\in S$ with more than $\Delta-3\sqrt{\Delta}$ neighbors in $S$ is not in any $Big_{i}$ set.
    Now, the argument is two-fold. If $u\in N(v)\cap S$ belongs to some $Big_{i}$, Molloy and Reed reason that $u$ is incident to at least $\Delta^{9/10}/2$ non-edges in $F[N(v)\cap S]$. Hence, if $v$ has at least $\alpha\Delta^{6/10}$ big neighbors, it also has at least $(\alpha/2)\Delta^{3/2}$ anti-edges in $F[N(v)\cap S]$, where $\alpha=2\cdot 10^{6}$. On the other hand, if it has fewer than $\alpha\Delta^{6/10}$ big neighbors, its induced neighborhood gained at most $(\alpha\Delta^{6/10})^{2}\leqslant\alpha^{2}\Delta^{12/10}=o(\Delta^{3/2})$ edges. As such, it has sparsity at least $d-o(\Delta^{1/2})\geqslant 9.9\cdot 10^{5}$.
  2. 2.
    By definition, vertices in $B_{H}$ have at most $c-\Delta^{3/4}$ neighbors in $F\setminus\bigcup_{j}A_{j}$. Hence, for each $v\in B_{H}$ at most $x(v)=c-\Delta^{3/4}-deg_{B_{H}}(v)$ neighbors are in $S$ and already colored. Thus, the list of $v$ when processing $B_{H}$ is at least of size $|L(v)|\geqslant c-x(v)=deg_{B_{H}}+\Delta^{3/4}$.
  3. 3.
    Follows from Lemma A.1 (b).
  4. 4.
    For each $v\in B_{L}$, Lemma 12(f) guarantees that there is some $A_{i}\in A$ such that $v$ has at most $c-\sqrt{\Delta}+9$ neighbors outside of $A_{i}$. Lemma 12(f) even states that $A_{i}\in A_{L}$. In other words $v$ has at most $x(v)=c-\sqrt{\Delta}+9-deg_{B_{L}}(v)$ nodes outside of $A_{i}\in A_{L}$ and $B_{L}$. As all nodes in $A_{L}$ and $B_{L}$ are uncolored, we obtain $|L(v)|\geqslant c-x(v)\geqslant\deg_{B_{L}}(v)+\sqrt{\Delta}-9$.
  5. 5.
    The upper bound on the external degree follows from the definition of $A_{L}$.
Additionally, we prove the second list of properties:
  1. a)
    This is identical to Lemma A.1 (a).
  2. b)
    This follows from Lemma A.1 (c).
  3. c)
    The set $\operatorname{Big}_{i}^{+}$ forms a clique due to Lemma A.1 (g). The upper bound of the degree into $A_{i}$ from a vertex in $\operatorname{Big}_{i}^{+}$ follows from Lemma A.1 (c).
Appendix B Concentration Inequalities
Theorem 4 (Janson’s inequality).
Let $S\subseteq V$ be a random subset formed by sampling each $v\in V$ independently with probability $p$. Let $\mathcal{A}$ be any collection of subsets of $V$. For each $A\in\mathcal{A}$, let $I_{A}:=\bm{1}(A\subseteq S)$ be an indicator variable for the event that $A$ is contained in $S$. Let $f:=\sum_{A\in\mathcal{A}}I_{A}$ and $\mu:=\operatorname{\mathbb{E}}[f]$. Define
$K:=\frac{1}{2}\sum_{A,B\in\mathcal{A},A\cap B\neq\emptyset}\operatorname{\mathbb{E}}[I_{A}I_{B}]$ (1)
Then, for any $0\leqslant t\leqslant\operatorname{\mathbb{E}}[f]$,
${\mathbb{P}}[f\leqslant\operatorname{\mathbb{E}}[f]-t]\leqslant\exp\left(-\frac{t^{2}}{2\mu+K}\right)$
We use the classic Chernoff Bound on sums of independent random variables. See [20] or [21] for details.
Proposition B.1 (Chernoff bounds).
Let $X_{1},\ldots,X_{n}$ be a family of independent random variables with values in $[0,1]$, and let $X=\sum_{i\in[n]}X_{i}$. Suppose $\mu_{\mathsf{L}}\leqslant\mathbb{E}[X]\leqslant\mu_{\mathsf{H}}$, then
$\text{for all }\delta\geqslant 0,\text{ we have that }$ ${\mathbb{P}}\left[X>(1+\delta)\mu_{\mathsf{H}}\right]$ $\leqslant\exp\left(-\frac{\delta^{2}}{2+\delta}\mu_{\mathsf{H}}\right)\ .$ (2)
$\text{for all }\delta\in(0,1),\text{ we have that }$ ${\mathbb{P}}\left[X<(1-\delta)\mu_{\mathsf{L}}\right]$ $\leqslant\exp\left(-\frac{\delta^{2}}{2}\mu_{\mathsf{L}}\right)\ .$ (3)
Let $Y_{1},\ldots,Y_{n}$ be boolean random variables (with values in $\{0,1\}$). We say they form a read-$k$ family if they can be expressed as a function of independent random variables $X_{1},\ldots,X_{m}$ such that each $X_{j}$ influences at most $k$ variables $Y_{i}$. More formally, there are sets $P_{i}\subseteq[m]$ for each $i\in[n]$ such that: (1) for each $i\in[n]$, the variable $Y_{i}$ is a function of $\{X_{j}:j\in P_{i}\}$ and (2) $|\{i:j\in P_{i}\}|\leqslant k$ for each $j\in[m]$.
Proposition B.2 (read-$k$ bound, [25]).
Let $Y_{1},\ldots,Y_{n}$ be a read-$k$ family of boolean variables, and let $Y$ be their sum. Then, for any $\delta>0$,
${\mathbb{P}}[\left|Y-\mathbb{E}{Y}\right|>\delta n]\leqslant 2\exp\left(-\frac{2\delta^{2}n}{k}\right)\ .$
Lemma B.3 (Azuma-Hoeffding).
Let $Z=Z_{1}+\ldots+Z_{n}$ be the sum of n random variables and $X_{0},\ldots,X_{n}$ be a sequence, where $Z_{i}$ is uniquely determined by $X_{0},\ldots,X_{i}$. Let $\mu_{i}=\mathbb{E}[Z_{i}\nonscript\>|{}\nonscript\>\mathopen{}X_{0},\ldots,X_{i-1}]$, $\mu=\sum_{i}\mu_{i}$, and $a_{i}\leqslant Z_{i}\leqslant a^{\prime}_{i}$. Then
${\mathbb{P}}[{Z\geqslant\mu+t}],{\mathbb{P}}[Z\leqslant\mu-t]\leqslant\exp{\left(-\frac{t^{2}}{2\sum_{i}(a^{\prime}_{i}-a_{i})^{2}}\right)}.$
We use the following Talagrand inequality. A function $f(x_{1},\dots,x_{n})$ is called $c$-Lipschitz iff the value of any single $x_{i}$ affects $f$ by at most $c$. Additionally, $f$ is $r$-certifiable if for every $x=(x_{1},\dots,x_{n})$, (1) there exists a set of indices $J(x)\subseteq[n]$ such that $|J(x)|\leqslant r\cdot f(x)$, and (2) if $x^{\prime}$ agrees with $x$ on the coordinates in $J(x)$, then $f(x^{\prime})\geqslant f(x)$.
Lemma B.4 (Talagrand’s Inequality [45]).
Let $X_{1},\dots,X_{n}$ be independent random variables and $f(X_{1},\dots,X_{n})$ be a $c$-Lipschitz, $r$-certifiable function. For any $b\geqslant 1$:
${\mathbb{P}}[|f-\operatorname{\mathbb{E}}[f]|>b+60c\sqrt{r\operatorname{\mathbb{E}}[f]}]\leqslant 4\exp{\left(-\frac{b^{2}}{8c^{2}r\operatorname{\mathbb{E}}[f]}\right)}$
McDiarmid [43] extended Talagrand’s inequality to the setting where $X$ depends also on permutations.
Lemma B.5 (McDiarmid’s Inequality [45, Section 3.1]).
Let $X$ be a non-negative random variable determined by independent trials $T_{1},\ldots,T_{n}$ and independent permutations $\Pi_{1},\ldots,\Pi_{m}$. Suppose that for every set of possible outcomes of the trials and permutations, we have:
  1. 1.
    changing the outcome of any of the trials affects the outcome by at most $c$,
  2. 2.
    interchanging two elements in any one permutation can affect $X$ by at most $c$, and
  3. 3.
    for each $s\geqslant 0$, if $X\geqslant s$ then there exists a set of at most $rs$ choices whose outcomes certify that $X\geqslant s$.
Then for any $t>0$, we have
[…]
Appendix C Slack Generation
See 6.9
Proof.
Let $N_{A}(v)=N(v)\cap A$ for short. Let $X\subseteq\binom{N_{A}(v)}{2}$ be the set of non-edges, where $|X|=\overline{m}$. Let $\chi(w)$ be the random color chosen by $w$, and let $\varphi(w)$ be the possible permanent color, or $\varphi(w)=\bot$ in case of a conflict.
Let $Z$ be the number of colors $\psi\in[q]$ such that there exists a non-edge $\{u,w\}\in X$ with $\chi(u)=\chi(w)=\psi$, and for all such non-edges, $\varphi(u)=\varphi(w)=\psi$, i.e. all the nodes retain the color (see below for a formal definition with quantifiers). Say that a non-edge $\{u,w\}\in X$ is successful if $\varphi(u)=\varphi(w)\neq\bot$, and no node in $N_{A}(v)\setminus\{u,w\}$ picks the same color. Let $Y_{\{u,w\}}$ be an indicator function for the event that $\{u,w\}$ is successful. We have $\operatorname{\mathbb{E}}[Z]\geqslant\sum_{\{u,w\}\in X}\mathbb{E}[Y_{\{u,w\}}]$, since each non-edge with $Y_{\{u,w\}}=1$ counts towards $Z$. The probability of a non-edge being successful is at least
${\mathbb{P}}[Y_{\{u,w\}}=1]\geqslant\frac{1}{q}\left(\frac{q-1}{q}\right)^{2\Delta-2}\left(\frac{q-1}{q}\right)^{\Delta-2}\geqslant\frac{1}{q}\left(1-\frac{1}{q}\right)^{3\Delta}$
where $u$ and $w$ select the same color w.p. $1/q$; with probability $(\frac{q-1}{q})^{2\Delta-2}$ none of the neighbors of $u$ nor $v$ choose that particular color, and $(\frac{q-1}{q})^{\Delta-2}$ is the probability that no other neighbors of $v$ choose that color. Using that $q\geqslant\Delta/3$ and that $\Delta\geqslant 30$, we can lower bound this probability as
${\mathbb{P}}[Y_{\{u,w\}}=1]\geqslant\frac{1}{q}\exp\left(-\frac{3\Delta}{q-1}\right)\geqslant\frac{e^{-10}}{q}$ (using that $1-x\geqslant\exp\left(-\frac{1}{x-1}\right)$ for all $x>1$)
which gives
$\operatorname{\mathbb{E}}[Z]\geqslant\sum_{\{u,w\}\in X}\mathbb{E}[Y_{\{u,w\}}]\geqslant e^{-10}\cdot\overline{m}/q\ .$
Next, we show that $Z$ is concentrated around its mean. Let $T$ be the number of colors that are randomly chosen in TryColor by both nodes of at least one non-edge in $X$. Let $D$ be the number of colors that are chosen in TryColor by both nodes of at least one non-edge in $X$, but are not retained by at least one of them. Formally,
$T$ :=# $:=\#\text{ colors }\psi\text{ s.t. }$ $\exists\{u,w\}\in X:\chi(u)=\chi(w)=\psi$
$D$ :=# $:=\#\text{ colors }\psi\text{ s.t. }$ $\big(\exists\{u,w\}\in X:\chi(u)=\chi(w)=\psi\big)\;\wedge$
$\big(\exists\{u,w\}\in X:(\chi(u)=\chi(w)=\psi)\wedge(\varphi(u)=\bot\vee\varphi(w)=\bot)\big)$
$Z$ :=# $:=\#\text{ colors }\psi\text{ s.t. }$ $\big(\exists\{u,w\}\in X:\chi(u)=\chi(w)=\psi\big)\;\wedge$
$\big(\forall\{u,w\}\in X:(\chi(u)=\chi(w)=\psi)\implies(\varphi(u)=\varphi(w)=\psi)\big)$
We have Z=TDZ=T-D, since $D$ counts the colors where the implication in the definition of $Z$ fails.
We upper bound $\mathbb{E}[T]$ (which implies the same bound for $D$, as $D\leqslant T$). For a fixed color $c$, the probability that both $u,w$ pick $c$ is at most $1/q^{2}$. By union bound, the probability that $c$ is picked by at least one non-edge is at most $\overline{m}/q^{2}$. There are $q$ colors, so $\mathbb{E}[T]\leqslant q\cdot\overline{m}/q^{2}=\overline{m}/q$.
The functions $T$ and $D$ are $r$-certifiable with $r=2$ and $r=3$, respectively. See the appendix and Lemma B.4 for the definition of an $r$-certifiable function. $T$ and $D$ are both $2$-Lipschitz: whether a node is activated, and which color it picks affects the outcome by at most $c=2$. We apply Lemma B.4 with $b=\mathbb{E}[Z]/10-60c\sqrt{r\cdot\mathbb{E}[T]}$ (which is non-negative for $\overline{m}/q\geqslant c_{0}$ and $c_{0}$ a large enough constant):
${\mathbb{P}}[|T-\mathbb{E}[T]|\geqslant b+60c\sqrt{r\cdot\mathbb{E}[T]}]$ $={\mathbb{P}}[|T-\mathbb{E}[T]|\geqslant\mathbb{E}[Z]/10]$
$\leqslant 4\exp\left(-\frac{\big(\mathbb{E}[Z]/10-60c\sqrt{r\cdot\mathbb{E}[T]}\big)^{2}}{8c^{2}r\mathbb{E}[T]}\right)$
$\leqslant\exp\left(-\Theta(1)\left(\frac{\mathbb{E}[Z]^{2}}{\mathbb{E}[T]}-\frac{\mathbb{E}[Z]}{\sqrt{\operatorname{\mathbb{E}}[T]}}+O(1)\right)\right)$
$\leqslant\exp\left(-\Omega(\overline{m}/q)\right)$
In the last inequality, we used that $\operatorname{\mathbb{E}}[Z]\geqslant e^{-10}\overline{m}/q$ and $\mathbb{E}[T]\leqslant\overline{m}/q$. The same concentration bound applies for $D$, meaning that ${\mathbb{P}}[|D-\mathbb{E}[D]|\geqslant\operatorname{\mathbb{E}}[Z]/10]\leqslant\exp\left(-\Omega(\overline{m}/q)\right)$. By union bound, neither of the events $|T-\mathbb{E}[T]|\geqslant\operatorname{\mathbb{E}}[Z]/10$ and $|D-\mathbb{E}[D]|\geqslant\operatorname{\mathbb{E}}[Z]/10$ occur with probability at least $1-2\exp\left(-\Omega(\overline{m}/q)\right)$. Hence,
$Z=T-D\geqslant\operatorname{\mathbb{E}}[T]-\operatorname{\mathbb{E}}[Z]/10-(\operatorname{\mathbb{E}}[D]+\operatorname{\mathbb{E}}[Z]/10)=(4/5)\cdot\operatorname{\mathbb{E}}[Z]\geqslant\frac{\overline{m}/q}{3\cdot 10^{4}}\ ,$
with probability at least $1-\exp\left(-\Omega(\overline{m}/q)\right)$.
See 6.10
Proof.
Let $\overline{E}$ be the set of non-edges in $H[N(v)]$. For each non-edge $e\in\overline{E}$, define an indicator variable $I_{e}$ for the event that the non-edge $e$ is preserved in $H[N(v)\cap A]$. Call $Z$ the number of anti-edges in $H[N(v)\cap A]$, i.e., we have $\operatorname{\mathbb{E}}[Z]=\sum_{e\in\overline{E}}\operatorname{\mathbb{E}}[I_{e}]=\overline{m}p^{2}$. We can bound the quantity Eq 1 as
$K=\frac{1}{2}\sum_{e,e^{\prime}\in\overline{E},e\cap e^{\prime}\neq\emptyset}\operatorname{\mathbb{E}}[I_{e}I_{e^{\prime}}]\leqslant\frac{1}{2}\overline{m}(2\Delta-2)p^{3}\leqslant\overline{m}\Delta p^{3}\ ,$
where the first inequality uses that endpoints of $e$ have maximum degree $\Delta$. Using Theorem 4 with $t=\operatorname{\mathbb{E}}[Z]/2$, we can compute ${\mathbb{P}}[Z\leqslant p^{2}\overline{m}/2]={\mathbb{P}}[Z\leqslant\operatorname{\mathbb{E}}[f]-t]$, where
${\mathbb{P}}[Z\leqslant\operatorname{\mathbb{E}}[f]-t]\leqslant\exp\left(-\frac{t^{2}}{2\operatorname{\mathbb{E}}[f]+K}\right)$ $\leqslant\exp\left(-\frac{p^{4}\overline{m}^{2}/4}{2p^{2}\overline{m}+p^{3}\overline{m}\Delta}\right)$
$=\exp\left(-\frac{p^{2}\overline{m}}{8+4p\Delta}\right)\leqslant\exp\left(-\frac{p\overline{m}}{5\Delta}\right)$
using that $p\Delta\geqslant 8$ by assumption.

### 01 Apr 2026 v2

@@
 In Section 2.3, we attempt to give a bird’s eye view of the techniques allowing us to solve each instance of the Lovász Local Lemma in sublogarithmic time.
-2.1 Color Coverage (CC) & $\Pi$-ous subgraphs
+2.1 Color Coverage ($C$) & $\Pi$-ous subgraphs
 Due to the impossibility of completing all partial solutions in a non-greedy coloring problem, the hardest part of the problem is coloring the last vertices.
@@
 Concretely, this means that for every clique and every color, only a small number of vertices are forbidden from using that color.
-We formalize this requirement as the color coverage (henceforth CC) property.
-Ensuring that CC holds until coloring the dense parts of the graph is essential to completing the coloring.
-See Section 4.2 for the precise definition of CC.
+We formalize this requirement as the color coverage (henceforth $C$) property.
+Ensuring that $C$ holds until coloring the dense parts of the graph is essential to completing the coloring.
+See Section 4.2 for the precise definition of $C$.
 $\Pi$-ous subgraphs.
-The CC condition can be maintained if all earlier coloring steps are sufficiently random.
-If each dense node has at most $U$ neighbors outside its clique, and each of them picks one random color out of $U\cdot\Delta^{0.22}$ colors, then few of these outside neighbors pick the same color, thereby maintaining the CC condition.
-Thus, the key to ensuring the CC condition in each step of the algorithm is to maintain large lists of available colors $L(v)$ (colors that do not currently appear in the neighborhood of $v$) compared to the number of outside neighbors of the cliques.
+The $C$ condition can be maintained if all earlier coloring steps are sufficiently random.
+If each dense node has at most $U$ neighbors outside its clique, and each of them picks one random color out of $U\cdot\Delta^{0.22}$ colors, then few of these outside neighbors pick the same color, thereby maintaining the $C$ condition.
+Thus, the key to ensuring the $C$ condition in each step of the algorithm is to maintain large lists of available colors $L(v)$ (colors that do not currently appear in the neighborhood of $v$) compared to the number of outside neighbors of the cliques.
 We capture this condition through $\Pi$, which we state in a simplified form as follows:
@@
 Subgraphs satisfying $\Pi$ are called $\Pi$-ous (pronounced pious).
-We can color a $\Pi$-ous subgraph $H$ while maintaining CC by a combination of iterated random color trials (Rct) and multi color trials (Mct).
+We can color a $\Pi$-ous subgraph $H$ while maintaining $C$ by a combination of iterated random color trials (Rct) and multi color trials (Mct).
 Condition (b) of $\Pi$ ensures that the list of available colors for each vertex of $H$ has $U\cdot\Delta^{0.22}$ slack that remains and ensures a sufficiently large list regardless of how the coloring of $H$ proceeds.
@@
 2.2 Overview of the Coloring Algorithm
-Our algorithm centers on identifying $\Pi$-ous subgraphs that can be colored while maintaining CC, thereby ensuring that the dense vertices can be colored last.
+Our algorithm centers on identifying $\Pi$-ous subgraphs that can be colored while maintaining $C$, thereby ensuring that the dense vertices can be colored last.
 We build on the structural decomposition of Molloy and Reed, which nearly [^4] The subgraphs induced by $B_{H}$ and $B_{L}$ are $\Pi$-ous by construction, however the subgraph induced by $S$ is not $\Pi$-ous right away; see Section 6 for more details.
@@
 Each such unhappy vertex $v$ seeks a partner $u$ in $A_{i}$ so that both become happy if they swap their colors.
-CC crucially guarantees a large supply of valid swap candidates.
+$C$ crucially guarantees a large supply of valid swap candidates.
 To execute swaps in all cliques concurrently, we first independently downsample the candidate sets and remove those causing external conflicts, i.e., swaps that cannot be executed safely in parallel with sampled candidate swaps in adjacent cliques.
@@
 As one randomly colors vertices in the pre-shattering phase, one may be forced to uncolor certain vertices.
-For instance, if a color appears too frequently around a clique and CC is violated.
+For instance, if a color appears too frequently around a clique and $C$ is violated.
 This may interfere with the progress measures of our algorithm, which are typically formulated as additional bad events.
@@
 Since we only pick candidates in pre- and post-shattering using separate budgets, this increases the number of neighbors with a given color by a factor of at most two compared to having one single budget for both steps.
-Conveniently, the same analysis bounds the probability of breaking CC for the pre- and post-shattering.
+Conveniently, the same analysis bounds the probability of breaking $C$ for the pre- and post-shattering.
 Removing candidates does not make other candidates bad, however, resampling candidates to repair local failures can affect neighboring cliques, and potentially undo the progress made during pre-shattering.
@@
 It therefore suffices to upper bound the size of the largest connected component of $H[N_{G}^{\leqslant c_{\mathcal{B}^{\prime}}}(\mathcal{A}^{\prime})]$.
-By (S2), (S4), and the union bound over all sets $A$ with $v$ as a vertex, we have that
+By (S2), (S4), and the union bound over all sets AA with $v$ as a vertex, we have that
 ${\mathbb{P}}\left[v\in\bigcup_{A\in\mathcal{A}^{\prime}}A\right]\leqslant\Delta^{c_{2}-c_{4}}\ .$
@@
 ${\mathbb{P}}\left[\operatorname{dist}_{G}(v,\mathcal{A^{\prime}})\leqslant c_{\mathcal{B}^{\prime}}\right]\leqslant\Delta^{c_{\mathcal{B}^{\prime}}+1}\cdot\Delta^{c_{2}-c_{4}}\leqslant\Delta(H)^{(c_{\mathcal{B}^{\prime}}+1+c_{2}-c_{4})/(3c_{1})}\ .$
-By (S3), whether $A\in\mathcal{A}^{\prime}$ depend only on the random variables $X_{u}$ of vertices in $A$, which has diameter one in $H$.
-Whether $v\in N_{G}^{\leqslant c_{\mathcal{B}^{\prime}}}(\mathcal{A}^{\prime})$ depends on whether $A\in\mathcal{A}^{\prime}$ for all the $A$ with a vertex in $N^{\leqslant c_{\mathcal{B}^{\prime}}}_{G}(v)$, hence it depends only on the random values within $c_{\mathcal{B}^{\prime}}+1$ hops in $H$.
+By (S3), whether $A\in\mathcal{A}^{\prime}$ depend only on the random variables $X_{u}$ of vertices in AA, which has diameter one in $H$.
+Whether $v\in N_{G}^{\leqslant c_{\mathcal{B}^{\prime}}}(\mathcal{A}^{\prime})$ depends on whether $A\in\mathcal{A}^{\prime}$ for all the AA with a vertex in $N^{\leqslant c_{\mathcal{B}^{\prime}}}_{G}(v)$, hence it depends only on the random values within $c_{\mathcal{B}^{\prime}}+1$ hops in $H$.
 Let $C_{1}=(c_{4}-c_{\mathcal{B}^{\prime}}-c_{2}-1)/(3c_{1})$, $C_{2}=c_{\mathcal{B}^{\prime}}+1$ and $C_{3}=10$; it is easy to verify that $C_{1}>C_{3}+4C_{2}+2$ for our choice of $c_{1},c_{2},c_{4},c_{\mathcal{B}^{\prime}}$.
@@
 The proof of Lemma 4.1 appears in Appendix A.
-4.2 The Color Coverage (CC) Property
-In this section, we formally introduce the color coverage property, which we abbreviate CC, that was outlined in Section 2.
+4.2 The Color Coverage ($C$) Property
+In this section, we formally introduce the color coverage property, which we abbreviate $C$, that was outlined in Section 2.
 Informally speaking we wish to ensure that the coloring outside of cliques $A_{i}$ looks random enough.
@@
 While $\mathrm{All}_{i}$ contains the nodes that are connected to all nodes in $A_{i}$, the nodes in $\operatorname{Big}_{i}^{+}$ are connected to most nodes (but still not too many) of $A_{i}$.
-Hence, coloring a single node in $\operatorname{Big}_{i}^{+}$ may bring $A_{i}$ close to not satisfying CC.
+Hence, coloring a single node in $\operatorname{Big}_{i}^{+}$ may bring $A_{i}$ close to not satisfying $C$.
 But due to Lemma 4.1 the nodes in $\operatorname{Big}_{i}^{+}$ form a clique for $A_{i}$; hence, no two nodes in $\operatorname{Big}_{i}^{+}$ can receive the same color.
-Property CC.
+Property $C$.
 Consider some arbitrary step $j$ of a coloring algorithm.
 Let $\text{CC}_{j}(i,x)$ be the number of vertices in $A_{i}$ that have a neighbor $v\not\in A_{i}\cup\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ that gets colored with $x$ in step $j$.
-We say that CC is within budget for a single coloring step if $\text{CC}_{j}(i,x)<\Delta^{37/40}$ holds for every uncolored clique $A_{i}$ and for every color $x$.
-We say that CC holds for a clique $A_{i}$ at a point in time if $C$ has been within budget for every coloring step applied so far.
+We say that $C$ is within budget for a single coloring step if $\text{CC}_{j}(i,x)<\Delta^{37/40}$ holds for every uncolored clique $A_{i}$ and for every color $x$.
+We say that $C$ holds for a clique $A_{i}$ at a point in time if $C$ has been within budget for every coloring step applied so far.
 Essentially, the $\Delta^{37/40}$ budget ensures that across all steps, each color remains sufficiently under-used around every clique, enabling valid swaps later.
@@
 Observation 4.3.
-If we maintain CC throughout the coloring process for a clique $A_{i}$, then for all colors $x$ and at any point in the coloring process, at most $4\Delta/5$ vertices of $A_{i}$ have a neighbor outside of $A_{i}\cup\mathrm{All}_{i}$ with color $x$.
+If we maintain $C$ throughout the coloring process for a clique $A_{i}$, then for all colors $x$ and at any point in the coloring process, at most $4\Delta/5$ vertices of $A_{i}$ have a neighbor outside of $A_{i}\cup\mathrm{All}_{i}$ with color $x$.
 Proof.
@@
 That means that we color vertices in a pre-shattering and a post-shattering phase, both of which count as a separate coloring step in the context of this observation.
-In other words, the pre-shattering and post-shattering phases in these LLLs obtain a separate CC budget which simplifies the analysis significantly.
+In other words, the pre-shattering and post-shattering phases in these LLLs obtain a separate $C$ budget which simplifies the analysis significantly.
 , depending on the approach.
@@
 The probability that at least $\Delta^{37/40}$ sets contain at least one marked vertex is at most $\exp(-\Delta^{1/40})$.
-The high level idea of using Lemma 4.4 to show that CC is maintained for cliques in each step of our coloring procedure is as follows.
+The high level idea of using Lemma 4.4 to show that $C$ is maintained for cliques in each step of our coloring procedure is as follows.
 Most of our coloring steps are based on some variant of random color trials, in which each node picks one (or multiple) candidate colors and then permanently retains one of its candidate colors if none of its neighbors chose it as a candidate.
-To prove that such a process maintains CC, fix a color $x\in[c]$, a clique $A_{i}$, and define a set $N_{v}$ for each vertex $v$ of the clique containing its external neighbors, except those in $\operatorname{Big}_{i}^{+}$.
+To prove that such a process maintains $C$, fix a color $x\in[c]$, a clique $A_{i}$, and define a set $N_{v}$ for each vertex $v$ of the clique containing its external neighbors, except those in $\operatorname{Big}_{i}^{+}$.
 Let $Q$ be an upper bound on the sizes of the sets.
@@
 Hence, if the list of each vertex is of size $Q\cdot\Delta^{1/5}$ the properties of Lemma 4.4 are satisfied.
-We obtain that CC is maintained for color $x$ and clique $A_{i}$ with probability $\exp(-\Delta^{1/40})$.
+We obtain that $C$ is maintained for color $x$ and clique $A_{i}$ with probability $\exp(-\Delta^{1/40})$.
 Note that in this process the events whether color $x$ appears in the neighborhood of two vertices $v\neq v^{\prime}$ may not be independent as $v$ and $v^{\prime}$ may have a common external neighbor.
@@
 Property (b) ensures that every vertex in $H$ has slack $U\cdot\Delta^{0.22}$, and importantly this slack remains as other vertices of the subgraph $H$ are colored.
-This allows us to color $H$ with $O(\log\Delta)$ iterated random color trial while maintaining Property CC.
+This allows us to color $H$ with $O(\log\Delta)$ iterated random color trial while maintaining Property $C$.
 In Section 5 we prove the following lemma for coloring $\Pi$-ous subgraphs.
@@
 Let $H$ be a $\Pi$-ous subgraph of $F-\bigcup_{i}A_{i}$.
-There is a $\mathsf{LOCAL}$ algorithm that, w.h.p., extends the coloring to all the vertices of $H$ in $\widetilde{O}(\log\Delta\cdot\log^{3}\log n)$ rounds while maintaining Property CC.
+There is a $\mathsf{LOCAL}$ algorithm that, w.h.p., extends the coloring to all the vertices of $H$ in $\widetilde{O}(\log\Delta\cdot\log^{3}\log n)$ rounds while maintaining Property $C$.
 Note that when $\Delta=n^{\varepsilon}$, Lemma 4.5 does not guarantee a sublogarithmic runtime.
-For $\Delta\geqslant(\log n)^{50}$, we adapt the state-of-the-art $(\deg+1)$-list-coloring algorithm of [35] for coloring a $\Pi$-ous subgraph while maintaining CC; see Section 8 for details.
+For $\Delta\geqslant(\log n)^{50}$, we adapt the state-of-the-art $(\deg+1)$-list-coloring algorithm of [35] for coloring a $\Pi$-ous subgraph while maintaining $C$; see Section 8 for details.
 Using Lemma 4.5 for low-degree graphs incurs an additional $O(\log\Delta)=O(\log\log n)$ factor in the runtime, but greatly simplifies the analysis.
@@
 Let $H$ be a $\Pi$-ous subgraph of $F-\bigcup_{i}A_{i}$ and assume $\Delta\geqslant(\log n)^{50}$.
-There is a $\mathsf{LOCAL}$ algorithm that list-colors all vertices of $H$ in $O(\log^{*}n)$ rounds while maintaining Property CC.
+There is a $\mathsf{LOCAL}$ algorithm that list-colors all vertices of $H$ in $O(\log^{*}n)$ rounds while maintaining Property $C$.
 We cannot use Lemma 4.5 directly to color the vertices of $S$ because they do not have enough initial slack to satisfy $\Pi$.
 As in [45], we begin by solving an LLL to provide them with $\Omega(\sqrt{\Delta})$ slack which is still not sufficient for $\Pi$.
-We then color the vertices of $S$ with Lemma 4.5 (or Lemma 4.6) in two batches to ensure that $\Pi$ holds, thereby that CC is maintained.
+We then color the vertices of $S$ with Lemma 4.5 (or Lemma 4.6) in two batches to ensure that $\Pi$ holds, thereby that $C$ is maintained.
 Details can be found in Section 6.
@@
 Let $c\geqslant\Delta-k_{\Delta}+1$ and $S$ be the set of vertices in $F$ as in Lemma 4.1-(1).
-There is a $\widetilde{O}(\log^{4}\log n)$-round distributed algorithm that, w.h.p., $c$-colors the vertices of $S$ while maintaining CC.
+There is a $\widetilde{O}(\log^{4}\log n)$-round distributed algorithm that, w.h.p., $c$-colors the vertices of $S$ while maintaining $C$.
 When $\Delta\geqslant(\log n)^{50}$, it runs in $O(\log^{*}n)$ rounds.
-As explained above, to color the cliques, we require that they satisfy CC.
-Since the cliques from $A_{L}$ are colored later, we must be careful about maintaining CC for those as we color $A_{H}$.
+As explained above, to color the cliques, we require that they satisfy $C$.
+Since the cliques from $A_{L}$ are colored later, we must be careful about maintaining $C$ for those as we color $A_{H}$.
 We prove Lemma 4.8 in Section 7.
@@
   1. 1.
-    all $A_{i}\in A^{\prime}$ satisfy CC, and
+    all $A_{i}\in A^{\prime}$ satisfy $C$, and
   2. 2.
     every uncolored vertex in some $A_{i}\notin A^{\prime}$ has at most $30\Delta^{1/4}$ external neighbors.
-Then, there is a $\widetilde{O}(\log^{3}\log n)$-round $\mathsf{LOCAL}$ algorithm that, w.h.p., $c$-colors all the cliques in $A^{\prime}$ while maintaining CC for all the uncolored cliques outside $A^{\prime}$.
+Then, there is a $\widetilde{O}(\log^{3}\log n)$-round $\mathsf{LOCAL}$ algorithm that, w.h.p., $c$-colors all the cliques in $A^{\prime}$ while maintaining $C$ for all the uncolored cliques outside $A^{\prime}$.
 For $\Delta\geqslant(\log n)^{50}$, it runs in $O(1)$ rounds.
@@
 The collection of cliques in $A_{H}$ is colored via Lemma 4.8.
-We can apply the lemma as their CC was maintained by earlier coloring steps (Lemmas 4.7, 4.5 and 4.6).
-We emphasize that, by Lemma 4.8, CC of cliques in $A_{L}$ still holds after coloring cliques in $A_{H}$.
+We can apply the lemma as their $C$ was maintained by earlier coloring steps (Lemmas 4.7, 4.5 and 4.6).
+We emphasize that, by Lemma 4.8, $C$ of cliques in $A_{L}$ still holds after coloring cliques in $A_{H}$.
 Coloring $B_{L}$.
@@
 Coloring $A_{L}$.
-The cliques in $A_{L}$ are also colored via Lemma 4.8, which can be done as CC holds for each of these cliques.
+The cliques in $A_{L}$ are also colored via Lemma 4.8, which can be done as $C$ holds for each of these cliques.
 Note that Lemma 4.8-(2) vacuously holds at this step because all the uncolored cliques are in $A^{\prime}=A_{L}$.
@@
 5 Coloring With Much Slack
-The goal of this section is to prove Lemma 4.5, that is, we aim to color an induced subgraph $H$ of $F-\bigcup_{i}A_{i}$ that is $\Pi$-ous while at the same time maintaining CC for all uncolored cliques.
+The goal of this section is to prove Lemma 4.5, that is, we aim to color an induced subgraph $H$ of $F-\bigcup_{i}A_{i}$ that is $\Pi$-ous while at the same time maintaining $C$ for all uncolored cliques.
 Recall that $H$ is $\Pi$-ous if each node has sufficient slack in $H$ compared to a given upper bound $U\geqslant\Delta^{1/4}$ on the external degree of uncolored cliques.
@@
   2. 2.
-    CC is maintained for all uncolored cliques.
+    $C$ is maintained for all uncolored cliques.
 Algorithm 3 runs in $\widetilde{O}(\log^{3}\log n)$ rounds.
@@
 Let $H$ be an uncolored induced subgraph in which every vertex has $\Delta^{9/20}$ slack and each uncolored vertex in each $A_{i}$ has at most $\Delta^{1/10}$ neighbors in $H$.
-There is a $\widetilde{O}(\log^{3}\log n)$-round $\mathsf{LOCAL}$ algorithm that, w.h.p., colors $H$ while maintaining CC for every uncolored clique $A_{i}$.
+There is a $\widetilde{O}(\log^{3}\log n)$-round $\mathsf{LOCAL}$ algorithm that, w.h.p., colors $H$ while maintaining $C$ for every uncolored clique $A_{i}$.
 When $\Delta\geqslant(\log n)^{50}$, the algorithm runs in $O(1)$ rounds.
@@
 The application of Lemma 5.2 to $H^{\prime}$ completes the coloring of $H$.
-The CC is preserved for all uncolored cliques as it is guaranteed by the outcome of Lemmas 5.1 and 5.2.
+The $C$ is preserved for all uncolored cliques as it is guaranteed by the outcome of Lemmas 5.1 and 5.2.
 The runtime consists of the $O(\log\Delta)$ iterations from Lemma 5.1 that take $\widetilde{O}(\log^{3}\log n)$ rounds each and $\widetilde{O}(\log^{3}\log n)$ rounds for the one application of Lemma 5.2.
 This completes the proof.
-5.1 Iterated Random Color Trial with CC (Lemma 5.1)
+5.1 Iterated Random Color Trial with $C$ (Lemma 5.1)
 The goal of this section is to prove Lemma 5.1.
-We present a $\operatorname{\text{{\rm poly}}}(\log\log n)$ algorithm that partially colors the graph $H$ while maintaining CC and additionally reduces the uncolored degree of vertices in $H$ and $\bigcup_{i}A_{i}$ below $\Delta^{1/10}$.
+We present a $\operatorname{\text{{\rm poly}}}(\log\log n)$ algorithm that partially colors the graph $H$ while maintaining $C$ and additionally reduces the uncolored degree of vertices in $H$ and $\bigcup_{i}A_{i}$ below $\Delta^{1/10}$.
 The analysis focuses on a single iteration of Random Color Trial (Rct).
-We show that in each iteration of the Rct procedure the uncolored degree of a vertex drops by a factor of (11/180)(1-1/180) compared to the previous iteration while maintaining CC.
+We show that in each iteration of the Rct procedure the uncolored degree of a vertex drops by a factor of (11/180)(1-1/180) compared to the previous iteration while maintaining $C$.
 We set up an LLL that, if solved correctly, yields a partial coloring guaranteeing this drop in the uncolored degree.
@@
 and for each uncolored clique $A_{i}$, we have
-$E_{CC}(i)$: CC is not maintained for some color w.r.t. $A_{i}$.
+$E_{CC}(i)$: $C$ is not maintained for some color w.r.t. $A_{i}$.
 Observation 5.3.
@@
 […]
-Lemma 5.5.
+Lemma $5.$
 Let $H^{\prime}$ be an induced uncolored subgraph of a $\Pi$-ous subgraph $H$ of $F-\bigcup_{i}A_{i}$.
@@
 By Lemma 4.1-(a), we have $|A_{i}|<\Delta$.
-We ignore vertices of $\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ for their contribution is not accounted for in CC.
+We ignore vertices of $\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ for their contribution is not accounted for in $C$.
 Hence, no relevant vertex of $H$ belongs to more than $2\Delta^{9/10}$ such subsets.
@@
 As candidate color choices are independent for different vertices, the probability that $\ell$ vertices pick color $x$ is at most $1/(Q\cdot\Delta^{1/5})^{\ell}$.
-Therefore, by Lemma 4.4, CC fails for clique $A_{i}$ and color $x$ w.p.
+Therefore, by Lemma 4.4, $C$ fails for clique $A_{i}$ and color $x$ w.p.
 at most $\exp(-\Delta^{1/40})$.
@@
 Remark 5.6.
-As detailed in CC, there is a fresh budget of $CC(i,x)<\Delta^{\frac{37}{40}}$ for each color $x$, each clique $A_{i}$, and also for each coloring step of the algorithm.
+As detailed in $C$, there is a fresh budget of $CC(i,x)<\Delta^{\frac{37}{40}}$ for each color $x$, each clique $A_{i}$, and also for each coloring step of the algorithm.
 In particular, in Algorithms 3 and 3 of Algorithm 3 the events $E_{CC}(i)$ get a separate budget of $\Delta^{\frac{37}{40}}$ each.
@@
 Step (S1) holds with $c_{1}=5$ due to 5.3 (and because $F$ includes the cliques): a vertex $v$ holds variables used only by events $E(u)$ with $u\in N_{H}^{\leqslant 2}(v)$ (which has diameter at most 4) or events $E_{CC}(i)$ with $A_{i}$ with $\operatorname{dist}_{G}(v,A_{i})\leqslant 2$ (which has diameter at most 5 in $F$).
-Step (S2) requires a bound $\Delta(F)^{c_{2}}$ for some constant $c_{2}$ on the number of events in which each variable appears and it holds with $c_{2}=2$ as each variable is only used by events with variables in its $2$-hop neighborhood.
+Step (S2) requires a bound $\Delta(F)^{c_{2}}$ for some constant $c_{2}$ on the number of events in which each variable appears and it holds with $c_{2}=2$ as each variable is only used by events with variables in its 22-hop neighborhood.
 Step (S3) requires that for each event $A\in\mathcal{A}$ we can determine $1(A\in\mathcal{A}^{\prime})$ by only evaluating the variables in $\operatorname{vbl}(A)$ which immediately holds by the definition of $\mathcal{A}^{\prime}$.
@@
 As $\Delta(G)=\Theta(\Delta(F))$ Step (S4) holds for $c_{4}$ (as chosen above) as long as $\Delta(G)$ is larger than a sufficiently large absolute constant $\Delta_{0}$.
-The set of events included in the LLL of Algorithm 3 are all included in $\mathcal{B}^{\prime}$ (with $c_{\mathcal{B}^{\prime}}$ defined above) as vertices with distance at most $2$ to $\operatorname{vbl}(\mathcal{A}^{\prime})$ participate in the post-shattering random process and any event $E(v)$ or $E_{CC}(i)$ adjacent to such a vertex participates in the post-shattering as well; note that the event $E_{CC}(i)$ for a clique $A_{i}$ that has the closest vertex $w$ in $H^{\prime}$ in distance $2$ does not need to participate despite $w\in\operatorname{vbl}(E_{CC}(i))$ as the event is avoided regardless the color choice of $w$.
+The set of events included in the LLL of Algorithm 3 are all included in $\mathcal{B}^{\prime}$ (with $c_{\mathcal{B}^{\prime}}$ defined above) as vertices with distance at most 22 to $\operatorname{vbl}(\mathcal{A}^{\prime})$ participate in the post-shattering random process and any event $E(v)$ or $E_{CC}(i)$ adjacent to such a vertex participates in the post-shattering as well; note that the event $E_{CC}(i)$ for a clique $A_{i}$ that has the closest vertex $w$ in $H^{\prime}$ in distance 22 does not need to participate despite $w\in\operatorname{vbl}(E_{CC}(i))$ as the event is avoided regardless the color choice of $w$.
 Hence by Lemma 3.2, the connected components of $G_{\operatorname{dep}}$ have size at most $O(\Delta(F)^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}+c_{2}}\log n)=\operatorname{\text{{\rm poly}}}(\log n)$ with high probability.
@@
 Since $H^{\prime}$ contains all the uncolored neighbors of $v$, the uncolored degree of $v$ at the end of the algorithm is upper bounded by the uncolored degree of $v$ in $H^{\prime}$.
-After retracting colors in $\operatorname{vbl}(\mathcal{A}^{\prime})$, none of the $E_{CC}(i)$ holds anymore and thus CC is maintained by the pre-shattering phase.
+After retracting colors in $\operatorname{vbl}(\mathcal{A}^{\prime})$, none of the $E_{CC}(i)$ holds anymore and thus $C$ is maintained by the pre-shattering phase.
 During post-shattering, only vertices of $N_{H}^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))$ get colored.
-Hence, CC continues to hold for the $A_{i}$ that are not adjacent to some vertex in $N_{H}^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))$.
-If $A_{i}$ is adjacent to such a vertex, then CC is maintained by the post-shattering coloring (with a fresh budget) because $E_{CC}(i)$ would otherwise hold.
+Hence, $C$ continues to hold for the $A_{i}$ that are not adjacent to some vertex in $N_{H}^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))$.
+If $A_{i}$ is adjacent to such a vertex, then $C$ is maintained by the post-shattering coloring (with a fresh budget) because $E_{CC}(i)$ would otherwise hold.
 5.2 Finishing Off The Coloring via MCT (Lemma 5.2)
@@
 $E^{\prime}(v)$: $v$ does not get colored,
-$E_{CC}(i)$: CC is not maintained for some color w.r.t. $A_{i}$.
+$E_{CC}(i)$: $C$ is not maintained for some color w.r.t. $A_{i}$.
 Observation 5.9.
@@
 No vertex lies in more than $2\Delta^{9/10}$ sets, because in that case it would belong to some $\operatorname{Big}_{i}^{+}$.
-Now we can apply Lemma 4.4 and get that for a fixed clique and color the probability that CC fails is at most $\exp(-\Delta^{1/40})$.
+Now we can apply Lemma 4.4 and get that for a fixed clique and color the probability that $C$ fails is at most $\exp(-\Delta^{1/40})$.
 With a union bound over all colors, $E_{CC}(i)$ holds w.p.
@@
 If $\Delta\geqslant(\log n)^{50}$, then by Lemmas 5.10 and 5.11, none of the bad events occur after running Algorithm 4 in pre-shattering, w.h.p..
-Hence, $H$ is colored and CC is maintained w.h.p..
+Hence, $H$ is colored and $C$ is maintained w.h.p..
 We henceforth assume that $\Delta\leqslant(\log n)^{50}$.
@@
 At the end, all nodes of $H$ are colored as all events $E^{\prime}(v)$ are either avoided already after the pre-shattering phase or after the post-shattering phase; note that any uncolored node after the pre-shattering phase is contained in $\operatorname{vbl}(\mathcal{A}^{\prime})$.
-After retracting colors in $\operatorname{vbl}(\mathcal{A}^{\prime})$, none of the $E_{CC}(i)$ holds anymore and thus also CC is maintained by the pre-shattering phase.
+After retracting colors in $\operatorname{vbl}(\mathcal{A}^{\prime})$, none of the $E_{CC}(i)$ holds anymore and thus also $C$ is maintained by the pre-shattering phase.
 During post-shattering, only vertices of $\operatorname{vbl}(\mathcal{A}^{\prime})$ get colored.
-Hence, CC continues to hold for the $A_{i}$ that are not adjacent to some vertex in $\operatorname{vbl}(\mathcal{A}^{\prime})$.
-If $A_{i}$ is adjacent to such a vertex, then CC is maintained by the post-shattering coloring (with a fresh budget) as it is implied by avoiding $E_{CC}(i)$.
+Hence, $C$ continues to hold for the $A_{i}$ that are not adjacent to some vertex in $\operatorname{vbl}(\mathcal{A}^{\prime})$.
+If $A_{i}$ is adjacent to such a vertex, then $C$ is maintained by the post-shattering coloring (with a fresh budget) as it is implied by avoiding $E_{CC}(i)$.
 6 Coloring the Sparse Nodes
-In this section, we show how to color the sparse nodes $S$ while maintaining $CC$. Recall that this is the first step of Algorithm 1.
+In this section, we show how to color the sparse nodes $S$ while maintaining CCCC. Recall that this is the first step of Algorithm 1.
 See 4.7
@@
   2. (b)
-    CC is maintained; and
+    $C$ is maintained; and
   3. (c)
@@
 For sufficiently large $\Delta$, it holds that $\Delta^{3/4}/40>U_{1}\cdot\Delta^{0.22}$, thereby proving Part (b) of $\Pi$.
-As for Part (a), recall that a vertex in $A$ has at most $U_{1}=10^{8}\sqrt{\Delta}$ neighbors in $S$ (by Lemma 4.1(3)).
+As for Part (a), recall that a vertex in AA has at most $U_{1}=10^{8}\sqrt{\Delta}$ neighbors in $S$ (by Lemma 4.1(3)).
 Proof of 2.
@@
 Pre-Shattering
-2 Each vertex of $S^{\prime}$ joins $S_{2}$ independently w.p. $p$.
+2 Each vertex of $S^{\prime}$ joins $S_{2}$ independently w.p. pp.
 3 For each $v$ with $\deg_{S^{\prime}}(v)\geqslant\alpha p^{-1}\log\Delta$, consider the event
@@
 The probability that a bad event $E_{d}(v)$ occurs during pre-shattering is low by a direct application of the Chernoff Bound.
-Claim 6.6.
+Claim $6.$
 For all $v$, the event $E_{d}(v)$ occurs w.p. at most $2\exp(-p\deg_{S^{\prime}}(v)/12)$.
@@
 We say that a color is repeated if it is used at least two times.
-We use the following variant of the result in which a set $A$ of active nodes try colors from a custom set of $q$ colors, which we prove in Appendix C.
+We use the following variant of the result in which a set AA of active nodes try colors from a custom set of $q$ colors, which we prove in Appendix C.
 Proposition 6.9.
@@
 Let $\Delta$ and $q$ be positive integers with $q\geqslant\Delta/3$.
-Let $A\subseteq V(G)$ be a subset of nodes such that $F[A]$ has maximum degree at most $\Delta$, and CC a set of $q$ colors.
+Let $A\subseteq V(G)$ be a subset of nodes such that $F[A]$ has maximum degree at most $\Delta$, and $C$ a set of $q$ colors.
 Consider a node $v\in V$ with at least $\overline{m}$ non-edges in $F[N(v)\cap A]$ and such that $\overline{m}\geqslant c_{0}\cdot q$.
-If every vertex of $A$ samples a random color $\chi(v)\in C$ and retains it only if $\chi(v)\notin\chi(N(v)\cap A)$, then
+If every vertex of AA samples a random color $\chi(v)\in C$ and retains it only if $\chi(v)\notin\chi(N(v)\cap A)$, then
 node $v$ has at least $\frac{\overline{m}/q}{3\cdot 10^{4}}$ repeated colored in $N(v)$
@@
 Suppose $H$ is a graph of maximum degree $\Delta$ and let $v$ be a vertex such that $H[N(v)]$ contains $\overline{m}$ anti-edges.
-Sample every vertex into a set $A$ with probability $p\geqslant 8/\Delta$.
+Sample every vertex into a set AA with probability $p\geqslant 8/\Delta$.
 The number of anti-edges induced by $H[N(v)\cap A]$ is at least $p^{2}\overline{m}/2$ w.p.
@@
   1. $E_{CC}(i)$:
-    CC is violated for clique $A_{i}$
+    $C$ is violated for clique $A_{i}$
 Pre-Shattering:
@@
   1. $E_{CC}(i)$:
-    CC is violated for clique $A_{i}$.
+    $C$ is violated for clique $A_{i}$.
 Algorithm 8 Implementation of Slack Generation (Lemma 6.1)
@@
 Step (S1) holds with a similar reasoning as in 5.3 (and because $F$ includes the cliques) for $c_{1}=5$.
-Step (S2) holds for $c_{2}=2$ as a variable is only contained in events with variables in its $2$-hop neighborhood.
+Step (S2) holds for $c_{2}=2$ as a variable is only contained in events with variables in its 22-hop neighborhood.
 Step (S3) follows directly with the definition of $\mathcal{A}^{\prime}$.
@@
 Hence, they form an LLL that can be solved in $\widetilde{O}(\log^{3}\log n)$ by Theorem 3.
-It remains to verify that the claimed properties indeed hold. CC is respected during both pre-shattering and post-shattering so Part (b) holds as well.
+It remains to verify that the claimed properties indeed hold. $C$ is respected during both pre-shattering and post-shattering so Part (b) holds as well.
 Part (a).
@@
 The goal of this section is to extend the coloring to the cliques in $A_{H}$ and $A_{L}$, respectively.
-In Algorithm 1, the cliques of $A_{H}$ are colored first, and thus when coloring these nodes we need to ensure that CC is satisfied for all cliques in $A_{L}$.
+In Algorithm 1, the cliques of $A_{H}$ are colored first, and thus when coloring these nodes we need to ensure that $C$ is satisfied for all cliques in $A_{L}$.
 We prove the following:
@@
 In the second phase, we perform color swaps: each node with a conflict finds a partner within its own clique with whom it can swap colors to remove the conflict without introducing new ones.
-This is where the CC property is crucially used.
+This is where the $C$ property is crucially used.
 Our Step 1 is largely similar to the first phase, with the difference of making the conflicts one-way directed in order to simplify the formulation in the shattering framework.
@@
   2. (b)
-    All cliques of $A^{\prime}$ and all uncolored cliques not in $A^{\prime}$ satisfy the CC property.
+    All cliques of $A^{\prime}$ and all uncolored cliques not in $A^{\prime}$ satisfy the $C$ property.
 The algorithm runs in a single round when $\Delta\geqslant(\log^{50}n)$ and in $\widetilde{O}(\log^{3}\log n)$-rounds otherwise.
@@
 We next show that a constant fraction of the nodes of $A_{i}$ are swappable (for any given node $v$).
-This lemma is the only place where we use the CC property.
-Its proof is purely deterministic, given the CC property of the clique and the guarantees of Lemma 7.2.
+This lemma is the only place where we use the $C$ property.
+Its proof is purely deterministic, given the $C$ property of the clique and the guarantees of Lemma 7.2.
 Lemma 7.4.
@@
 Nodes violating (b): The external degree of $v$ is bounded by $\sqrt{\Delta}$ (by Lemma 4.1(3,5)), and hence at most $\sqrt{\Delta}$ vertices in $A_{i}$ have a color appearing in the external neighborhood of $v$.
-Nodes violating (c): Note that CC holds for the clique $A_{i}$, as it held before Step 1 (recall Lemma 4.8(1)) and was maintained by Step 2 (by Lemma 7.2(b)).
+Nodes violating (c): Note that $C$ holds for the clique $A_{i}$, as it held before Step 1 (recall Lemma 4.8(1)) and was maintained by Step 2 (by Lemma 7.2(b)).
 Hence, by 4.3 each color (here applied to $\gamma(v)$) appears on an external neighbor (recall, one outside of $A_{i}\cup\mathrm{All}_{i}$) of at most $4\Delta/5$ vertices in $A_{i}$.
@@
   3. 3.
-    (Strong version of CC) For each uncolored $A_{j}\not\in A^{\prime}$ and color $x$, at most $2\cdot\Delta^{37/40}$ vertices in $A_{j}$ have a neighbor outside of $A_{j}\cup\mathrm{All}_{j}\cup\operatorname{Big}_{j}^{+}$ that has a candidate of color $x$ or is a candidate for a node of color $x$.
+    (Strong version of $C$) For each uncolored $A_{j}\not\in A^{\prime}$ and color $x$, at most $2\cdot\Delta^{37/40}$ vertices in $A_{j}$ have a neighbor outside of $A_{j}\cup\mathrm{All}_{j}\cup\operatorname{Big}_{j}^{+}$ that has a candidate of color $x$ or is a candidate for a node of color $x$.
 7.3 Step 3: Local Matching
@@
 We first observe that safeness implies that the color selection becomes a fully local problem within each clique.
-Observation 7.7.
+Observation $7.$
 Suppose an unhappy node $v$ performs a swap of colors with one of its safe candidates $u\in T_{v}$.
@@
 Then, by Hall’s theorem there is a matching $M$ that saturates $L$.
-We perform exactly the color swaps given by the matching $M$, which yields a proper coloring by 7.7.
-To conclude with the proof of Lemma 4.8, it remains to argue that the CC property is maintained for the remaining uncolored cliques.
+We perform exactly the color swaps given by the matching $M$, which yields a proper coloring by $7.$
+To conclude with the proof of Lemma 4.8, it remains to argue that the $C$ property is maintained for the remaining uncolored cliques.
 This follows for Step 1 by Lemma 7.2(b) and for Step 3 by Lemma 7.6(3).
@@
   2. $E_{b}(i)$
-    : $A_{i}$ is adjacent to a clique of $A^{\prime}$ or to an uncolored clique not in $A^{\prime}$ for which the CC property was broken.
+    : $A_{i}$ is adjacent to a clique of $A^{\prime}$ or to an uncolored clique not in $A^{\prime}$ for which the $C$ property was broken.
 Clearly, the events $E_{a}(i)$ and $E_{b}(i)$ depend on the random permutation of $A_{i}$ and of every $A_{j}$ adjacent to $A_{i}$.
@@
 Furthermore, at most one vertex in each clique is assigned color $x$ and the random permutation for different cliques are independent.
-Hence, we can apply Lemma 4.4 with $Q=10^{8}\sqrt{\Delta}$ as it bounds external degrees from above (by Lemma 4.1(3,5)), and get that $A_{j}$ does not satisfy CC for color $x$ with probability at most $\exp(-\Delta^{1/40})$.
+Hence, we can apply Lemma 4.4 with $Q=10^{8}\sqrt{\Delta}$ as it bounds external degrees from above (by Lemma 4.1(3,5)), and get that $A_{j}$ does not satisfy $C$ for color $x$ with probability at most $\exp(-\Delta^{1/40})$.
 By union bound on all colors and cliques adjacent to $A_{i}$, the event $E_{b}(i)$ occurs w.p.
@@
 Let us conclude the proof by proving that the resulting coloring indeed satisfies (a) and (b).
-For (b), it simply follows from the fact that after retractions (Algorithm 10), the CC is preserved, and since events $E_{b}(i)$ are avoided in post-shattering, CC is also preserved by this coloring step (with a fresh budget).
+For (b), it simply follows from the fact that after retractions (Algorithm 10), the $C$ is preserved, and since events $E_{b}(i)$ are avoided in post-shattering, $C$ is also preserved by this coloring step (with a fresh budget).
 To see why (a) holds, observe that every $A\notin A^{\prime\prime}$ satisfies (a) after Algorithm 10 (otherwise $E_{a}(i)$ occurs and $A_{i}$ is included in $A^{\prime\prime}$).
@@
 For a fixed external neighbor $u\in A_{\ell}\neq A_{j}$, the vertex with color $x$ in $u$’s clique (if any) gets sampled into $u$’s candidate set $S_{u}$ or samples $u$ into its candidate set with probability at most $p\leqslant\Delta^{-23/40}$.
-On the other hand, by assumption in Lemma 4.8, the external degrees of uncolored cliques not in $A^{\prime}$ is at most $30\Delta^{1/4}$ (recall that in this step, we do not need to guarantee that CC is maintained by cliques of $A^{\prime}$, in particular, cliques of $A_{H}$ do not need to verify Lemma 7.6(3)).
-And since the sampling is independent, it is easy to verify that (P7.1) holds with $Q=30\Delta^{1/4}$, thus that $B_{3}(i)$ occurs because of clique $A_{j}$ with probability at most $\exp(-\Delta^{1/40})$ by Lemma 4.4.
-By union bound on the colors and the at most $10^{8}\Delta^{3/2}$ cliques adjacent to $A_{i}$, we obtain the desired probability bound.
+On the other hand, by assumption in Lemma 4.8, the external degrees of uncolored cliques not in $A^{\prime}$ is at most $30\Delta^{1/4}$ (recall that in this step, we do not need to guarantee that $C$ is maintained by cliques of $A^{\prime}$, in particular, cliques of $A_{H}$ do not need to verify Lemma 7.6(3)).
+And since the sampling is independent, it is easy to verify that (P7.1) holds with $Q=30\Delta^{1/4}$, thus that $B_{3}(i)$ occurs because of clique $A_{j}$ with probability at most $\exp(-\Delta^{1/40})$ by Lemma $4.$ By union bound on the colors and the at most $10^{8}\Delta^{3/2}$ cliques adjacent to $A_{i}$, we obtain the desired probability bound.
 Claim 7.13.
@@
 See 4.6
-If it was not for the CC constraint, one could simply run the $(\deg+1)$-list-coloring algorithm of [35] to extend the coloring to $H$.
-Given an instance of $(\deg+1)$-list-coloring, it is not possible in general to ensure that the CC constraint will be maintained.
+If it was not for the $C$ constraint, one could simply run the $(\deg+1)$-list-coloring algorithm of [35] to extend the coloring to $H$.
+Given an instance of $(\deg+1)$-list-coloring, it is not possible in general to ensure that the $C$ constraint will be maintained.
 Here, we use that, thanks to $\Pi$, the nodes of $H$ have lists significantly larger than their degree.
@@
   • Synchronized color trials (SCT), when the nodes of an almost-clique are assigned a permutation of the colors in a palette.
-Using $\Pi$ and Lemma 4.4, we can show that each of the $O(\log^{*}n)$ steps of [35] maintain the CC property with high probability.
+Using $\Pi$ and Lemma 4.4, we can show that each of the $O(\log^{*}n)$ steps of [35] maintain the $C$ property with high probability.
 Proof of Lemma 4.6.
@@
 See [35, Algorithm 3].
-Since every vertex of $H$ samples one color out of a list of $U\Delta^{0.22}$ while the external degree of uncolored cliques is upper bounded by $U$ (by $\Pi$), Lemma 4.4 and union bound implies that CC is not maintained for some clique with probability at most $n\Delta\exp(-\Delta^{1/40})<1/\operatorname{\text{{\rm poly}}}(n)$.
+Since every vertex of $H$ samples one color out of a list of $U\Delta^{0.22}$ while the external degree of uncolored cliques is upper bounded by $U$ (by $\Pi$), Lemma 4.4 and union bound implies that $C$ is not maintained for some clique with probability at most $n\Delta\exp(-\Delta^{1/40})<1/\operatorname{\text{{\rm poly}}}(n)$.
 Note that the dummy nodes do not verify $\Pi$, as their lists are of size two, but they are not adjacent to the cliques of $F$.
@@
 $\frac{\log n}{U\Delta^{0.22}}\leqslant\frac{1}{U\Delta^{0.22-1/50}}\leqslant\frac{1}{U\Delta^{1/5}}\ .$
-Hence, we can apply Lemma 4.4 as before to reason that each step of MCT maintains CC with probability at least $1-\exp(-\Omega(\sqrt{s_{\min}}))-\Delta\exp(-\Omega(s_{\min}))\geqslant 1-1/\operatorname{\text{{\rm poly}}}(n)$.
+Hence, we can apply Lemma 4.4 as before to reason that each step of MCT maintains $C$ with probability at least $1-\exp(-\Omega(\sqrt{s_{\min}}))-\Delta\exp(-\Omega(s_{\min}))\geqslant 1-1/\operatorname{\text{{\rm poly}}}(n)$.
 The analysis of slack color for nodes of $H[C_{1}\cup\ldots\cup C_{q}]$ is the same as above.
@@
 Since the vertices have $\deg(v)=|L(v)|-1\geqslant U\Delta^{0.22}$, where the equality holds because of the dummy vertices and the inequality because of $\Pi$, an almost-clique $C_{i}$ contains at least $\Omega(U\Delta^{0.22})$ vertices.
-And thus the CC constraint is maintained by the SCT as every vertex tries gets some color $x$ with probability at most $O(1/|C_{i}|)\leqslant O(\frac{1}{U\Delta^{0.22}})\leqslant\frac{1}{U\Delta^{1/5}}$.
+And thus the $C$ constraint is maintained by the SCT as every vertex tries gets some color $x$ with probability at most $O(1/|C_{i}|)\leqslant O(\frac{1}{U\Delta^{0.22}})\leqslant\frac{1}{U\Delta^{1/5}}$.
 Appendix A Graph Decomposition by Molloy-Reed
@@
 Theorem 4 (Janson’s inequality).
-Let $S\subseteq V$ be a random subset formed by sampling each $v\in V$ independently with probability $p$.
+Let $S\subseteq V$ be a random subset formed by sampling each $v\in V$ independently with probability pp.
 Let $\mathcal{A}$ be any collection of subsets of $V$.
-For each $A\in\mathcal{A}$, let $I_{A}:=\bm{1}(A\subseteq S)$ be an indicator variable for the event that $A$ is contained in $S$.
+For each $A\in\mathcal{A}$, let $I_{A}:=\bm{1}(A\subseteq S)$ be an indicator variable for the event that AA is contained in $S$.
 Let $f:=\sum_{A\in\mathcal{A}}I_{A}$ and $\mu:=\operatorname{\mathbb{E}}[f]$.
@@
   3. 3.
-    for each $s\geqslant 0$, if $X\geqslant s$ then there exists a set of at most $rs$ choices whose outcomes certify that $X\geqslant s$.
+    for each $s\geqslant 0$, if $X\geqslant s$ then there exists a set of at most rsrs choices whose outcomes certify that $X\geqslant s$.
 Then for any $t>0$, we have
@@
 See the appendix and Lemma B.4 for the definition of an $r$-certifiable function.
-$T$ and $D$ are both $2$-Lipschitz: whether a node is activated, and which color it picks affects the outcome by at most $c=2$.
+$T$ and $D$ are both 22-Lipschitz: whether a node is activated, and which color it picks affects the outcome by at most $c=2$.
 We apply Lemma B.4 with $b=\mathbb{E}[Z]/10-60c\sqrt{r\cdot\mathbb{E}[T]}$ (which is non-negative for $\overline{m}/q\geqslant c_{0}$ and $c_{0}$ a large enough constant):